### v1

Abstract
Due to their importance in various emerging applications, efficient algorithms for solving minimax problems have recently received increasing attention. However, many existing algorithms require prior knowledge of the problem parameters in order to achieve optimal iteration complexity. In this paper, we propose a completely parameter-free alternating gradient projection (PF-AGP) algorithm to solve the smooth nonconvex-(strongly) concave minimax problems using a backtracking strategy, which does not require prior knowledge of parameters such as the Lipschtiz constant $L$ or the strongly concave constant $\mu$. The PF-AGP algorithm utilizes a parameter-free gradient projection step to alternately update the outer and inner variables in each iteration. We show that the total number of gradient calls of the PF-AGP algorithm to obtain an $\varepsilon$-stationary point for nonconvex-strongly concave minimax problems is upper bounded by $\mathcal{O}\left(L\kappa^{3}\varepsilon^{-2}\right)$ where $\kappa$ is the condition number, while the total number of gradient calls to obtain an $\varepsilon$-stationary point for nonconvex-concave minimax problems is upper bounded by $\mathcal{O}\left(L^{4}\varepsilon^{-4}\right)$. As far as we know, this is the first completely parameter-free algorithm for solving nonconvex-strongly concave minimax problems, and it is also the completely parameter-free algorithm which achieves the best iteration complexity in single loop method for solving nonconvex-concave minimax problems. Numerical results validate the efficiency of the proposed PF-AGP algorithm.
Keywords: minimax optimization problem; backtracking strategies; parameter-free alternating gradient projection algorithm; iteration complexity; machine learning
1 Introduction
We consider the following minimax optimization problem:
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}f(x,y),$ (P)
where $\mathcal{X}\subseteq\mathbb{R}^{n}$ and $\mathcal{Y}\subseteq\mathbb{R}^{m}$ are nonempty closed and bounded convex sets, and $f:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ is a smooth function. The minimax problem is a critical class of optimization problems, known for its extensive research history and broad applications in areas such as machine learning, signal processing, and various other fields in recent years. Many practical problems can be formulated as in (P), including machine learning, such as generative adversarial networks (GANs) (Arjovsky et al., 2017; Goodfellow et al., 2014; Sanjabi et al., 2018), reinforcement learning (Dai et al., 2018), power control and transceiver design in signal processing (Lu et al., 2020), distributed nonconvex optimization (Giannakis et al., 2017; Giordano et al., 2018; Mateos et al., 2010), robust optimization (Ben and Tal, 2009; Gao and Kleywegt, 2023), statistical learning (Giordano et al., 2018; Shafieezadeh Abadeh et al., 2015) and several other applications.
There are many research results on convex-concave minimax problems, i.e., $f(x,y)$ is convex with respect to $x$ and concave with respect to $y$. Nemirovski (Nemirovski, 2004) introduced the mirror-prox method, and Nestrov (Nesterov, 2007) proposed a dual extrapolation algorithm, both of which achieve an optimal iteration complexity of $\mathcal{O}(\varepsilon^{-1})$ for solving convex-concave minimax problems. For more details, see (Vandenberghe and Boyd, 2004; Chen et al., 2014, 2017; Dang and Lan, 2014; Monteiro and Svaiter, 2010; Nedić and Ozdaglar, 2009; Yang et al., 2020) and references therein.
Recently, various algorithms have been proposed for nonconvex-(strongly) concave minimax problems. For nonconvex-strongly concave minimax problems, several recent works (Jin et al., 2020; Lin et al., 2020a; Rafique et al., 2022; Lu et al., 2020) have proposed various algorithms that achieve a gradient complexity of $\tilde{\mathcal{O}}(\kappa_{y}^{2}\varepsilon^{-2})$ for stationary points of $\Phi(\cdot)=\max_{y\in\mathcal{Y}}f(\cdot,y)$ (when $\mathcal{X}=\mathbb{R}^{n}$), or for stationary points of $f$, where $\kappa_{y}$ is the condition number for $f(x,\cdot)$. (Lin et al., 2020b) proposed an accelerated algorithm that improves the gradient complexity bound to $\tilde{\mathcal{O}}(\sqrt{\kappa_{y}}\varepsilon^{-2})$. Furthermore, Zhang et al. (Zhang et al., 2021) proposed a generic acceleration framework which can improve the iteration complexity to $\mathcal{O}\left(\sqrt{\kappa_{y}}\varepsilon^{-2}\right)$.
For general nonconvex-concave minimax problems, there are two types of algorithms, i.e., multi-loop algorithms and single-loop algorithms. Various multi-loop algorithms have been proposed in (Kong and Monteiro, 2021; Lin et al., 2020b; Nouiehed et al., 2019; Ostrovskii et al., 2021; Rafique et al., 2022; Thekumparampil et al., 2019). The best known iteration complexity of multi-loop algorithms for solving nonconvex-concave minimax problems is $\tilde{\mathcal{O}}(\varepsilon^{-2.5})$, which was achieved by (Lin et al., 2020b). For solving nonconvex-linear minimax problems, Pan et al. (Pan et al., 2021) proposed a new alternating gradient projection algorithm with the iteration complexity of $\mathcal{O}(\varepsilon^{-3})$. Various single-loop algorithms have also been proposed, e.g., the gradient descent-ascent (GDA) algorithm (Lin et al., 2020a), the hybrid block successive approximation (HiBSA) algorithm (Lu et al., 2020), the unified single-loop alternating gradient projection (AGP) algorithm (Xu et al., 2023), and the smoothed GDA algorithm (Zhang et al., 2020). Both the AGP algorithm and the smoothed GDA algorithm achieve the best known iteration complexity, i.e., $\mathcal{O}(\varepsilon^{-4})$, among single-loop algorithms for solving nonconvex-concave minimax problems.
It should be noted that for most of the existing results mentioned above, achieving optimal complexity of the algorithm requires the assumption of knowing precise information about some parameters of the problem, such as the Lipschitz constant, or the strongly concave constant $\mu$, etc. Accurately estimating these parameters is often challenging, and conservative estimates can significantly impact algorithm performance (Lan et al., 2023). Therefore, designing parameter-free algorithms with complexity guarantees without relying on inputs of these parameters has attracted considerable attention recently.
For nonconvex-strongly concave minimax problems, there are two main parameter-free algorithms: one based on AdaGrad stepsizes and the other based on backtracking. In the framework of AdaGrad stepsizes, Yang et al. (Yang et al., 2022) proposed NeAda, which can achieve convergence rates of $\tilde{\mathcal{O}}(\varepsilon^{-2})$ and $\tilde{\mathcal{O}}(\varepsilon^{-4})$ in deterministic and stochastic settings, respectively. Li et al. (Li et al., 2022) proposed TiAda, achieving convergence rates of ${\mathcal{O}}(\varepsilon^{-2})$ and ${\mathcal{O}}(\varepsilon^{-4})$ in deterministic and stochastic settings, respectively. In the framework of backtracking, Xu et al. (Xu et al., 2024) proposed a stochastic gradient descent ascent (GDA) method with backtracking (SGDA-B), achieving an upper bound on the total number of gradient calls of $\mathcal{O}(L\kappa^{2}\log(\kappa)\varepsilon^{-2})$ with $\mathcal{O}(L\kappa^{2}\varepsilon^{-2})$ gradient calls required per backtracking step. Meanwhile, Zhang et al. (Zhang et al., 2024) proposed a proximal alternating gradient descent ascent method (AGDA+), achieving an upper bound on the total number of gradient calls of $\mathcal{O}(L\kappa^{4}\varepsilon^{-2})$ with $\mathcal{O}(1)$ gradient calls required per backtracking step. Both methods do not require knowledge of the specific value of Lipschtiz constant $L$, but still require information about the strongly concave constant $\mu$ in the function. In the framework of backtracking, Xu et al. (Xu et al., 2024) proposed a stochastic GDA method with backtracking (SGDA-B) for nonconvex-concave minimax problems, achieving an upper bound on the total number of gradient calls of $\tilde{\mathcal{O}}(L^{3}\varepsilon^{-4})$ with $\mathcal{O}(L^{3}\varepsilon^{-4})$ gradient calls required per backtracking step.
1.1 Contributions
In this paper, based on the framework of backtracking, we propose two completely parameter-free alternating gradient projection algorithms, i.e., the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm, for solving nonconvex-(strongly) concave minimax problems respectively, which does not require prior knowledge of parameters such as the Lipschtiz constant $L$ or the strongly concave constant $\mu$.
Moreover, we show that the total number of gradient calls of the PF-AGP-NSC algorithm to obtain an $\varepsilon$-stationary point for nonconvex-strongly concave minimax problems is upper bounded by $\mathcal{O}\left(L\kappa^{3}\varepsilon^{-2}\right)$ where $\kappa$ is the condition number, while the total number of gradient calls of the PF-AGP-NC algorithm to obtain an $\varepsilon$-stationary point for nonconvex-concave minimax problems is upper bounded by $\mathcal{O}\left(L^{4}\varepsilon^{-4}\right)$. As far as we know, the PF-AGP-NSC algorithm is the first completely parameter-free algorithm for solving nonconvex-strongly concave minimax problems, and the PF-AGP-NC algorithm is the first completely parameter-free algorithm which achieves the best iteration complexity in single loop algorithms for solving nonconvex-concave minimax problems.
1.2 Organization
In Sect. 2, we propose a parameter-free alternating gradient projection (PF-AGP) algorithm for nonconvex-(strongly) concave minimax problems, and we then analyze the corresponding gradient complexity for two different settings. We report some numerical results in Sect. 3 and make some concluding remarks in the last section.
[…]
$f(y)\geq f(x)+\langle\nabla f(x),y-x\rangle+\frac{\theta}{2}\|y-x\|^{2}.$ (1)
If $-f$ satisfies (1), $f(\cdot)$ is called $\theta$-strongly concave. Denote the sign function as $\mbox{sgn}(x)$, i.e.,
$\mbox{sgn}(x)=\begin{cases}1&\text{if }x>0,\\ -1&\text{if }x\leq 0.\end{cases}$ (2)
2 Two completely parameter-free alternating gradient projection algorithms
In this section, we propose two completely parameter-free alternating gradient projection (PF-AGP) algorithms for solving the nonconvex-strongly concave and the nonconvex-concave minimax problems (P) respectively. Both algorithms do not require prior knowledge of Lipschitz constants, and the algorithm under nonconvex-strong concave settings also does not require prior knowledge of strong concave constants. The two proposed algorithms are based on the AGP algorithm (Xu et al., 2023), which uses the gradient of a regularized version of the original function, i.e.,
$f_{k}(x,y)=f(x,y)-\frac{c_{k}}{2}\|y_{k}\|^{2},$
where $c_{k}\geq 0$ is a regularization parameters. At the $k$-th iteration, AGP algorithm consists of the following two gradient projection steps for updating both $x$ and $y$:
$x_{k+1}$ $=\mathcal{P}_{\mathcal{X}}\left(x_{k}-\tfrac{1}{\beta_{k}}\nabla_{x}f(x_{k},y_{k})\right),$
$y_{k+1}$ $=\mathcal{P}_{\mathcal{Y}}\left(y_{k}+\tfrac{1}{\gamma_{k}}\nabla_{y}{f}(x_{k},y_{k})-\frac{1}{\gamma_{k}}c_{k}y_{k}\right),$
where $\mathcal{P}_{\mathcal{X}}$ and $\mathcal{P}_{\mathcal{Y}}$ is the projection operator onto $\mathcal{X}$ and $\mathcal{Y}$, respectively, and $\beta_{k}>0$, $\gamma_{k}>0$ are stepsize parameters.
Although the AGP algorithm achieves the optimal iteration complexity among single loop algorithms for nonconvex-(strongly) concave minimax problems, it requires the knowledge of the Lipschtiz constant $L$ and the strongly concave coefficient $\mu$, or at least knowledge of the Lipschtiz constant $L$ to calculate the stepsizes $\beta_{k}$ and $\gamma_{k}$ under the nonconvex-strongly concave setting and the nonconvex-concave setting, respectively, which limits its applicability. To eliminate the need for the prior knowledge of the Lipschtiz constant $L$ or the strongly concave coefficient $\mu$, we propose two parameter-free AGP algorithms by employing backtracking strategies for solving the nonconvex-strongly concave and the nonconvex-concave minimax problems, respectively. Similar backtracking strategies have been used in (Lan et al., 2023; Liu and Luo, 2022; Nesterov and Polyak, 2006; Nesterov, 2015).
2.1 Nonconvex-strongly Concave Setting
We first propose a completely parameter-free AGP algorithm for solving nonconvex-strongly concave minimax problems. More specifically, in order to estimate the Lipschitz constants for $\nabla_{x}f(\cdot,y)$, $\nabla_{y}f(\cdot,y)$, $\nabla_{y}f(x,\cdot)$ denoted as $L_{11}$, $L_{12}$ and $L_{22}$, respectively, and the strongly concave constant $\mu$, at each iteration of the proposed algorithm, we aim to find a tuple $(l_{11}^{k,i},l_{12}^{k,i},l_{22}^{k,i},\mu_{k,i})$ by backtracking such that the following conditions for $(x_{k,i},y_{k,i})$ are satisfied:
$C_{1}^{k,i}$ $=f(x_{k,i},y_{k})-f(x_{k},y_{k})-\langle\nabla_{x}f(x_{k},y_{k}),x_{k,i}-x_{k}\rangle-\frac{l^{k,i}_{11}}{2}\|x_{k,i}-x_{k}\|^{2}\leq 0,$ (C1)
$C_{2}^{k,i}$ $=\|\nabla_{y}f(x_{k,i},y_{k})-\nabla_{y}f(x_{k},y_{k})\|-l_{12}^{k,i}\|x_{k,i}-x_{k}\|\leq 0,$ (C2)
$C_{3}^{k,i}$ $=\|\nabla_{y}f(x_{k,i},y_{k,i})-\nabla_{y}f(x_{k,i},y_{k})\|-l_{22}^{k,i}\|y_{k,i}-y_{k}\|\leq 0,$ (C3)
$C_{4}^{k,i}$ $=\langle\nabla_{y}f(x_{k,i},y_{k,i})-\nabla_{y}f(x_{k,i},y_{k}),y_{k,i}-y_{k}\rangle+\mu_{k,i}\|y_{k,i}-y_{k}\|^{2}\leq 0.$ (C4)
Otherwise, we enlarge $l_{11}^{k,i},l_{12}^{k,i},l_{22}^{k,i}$ by a factor of $2$ or reduce $\mu_{k,i}$ by half according to the sign of $C_{1}^{k,i}$, $C_{2}^{k,i}$, $C_{3}^{k,i}$ and $C_{4}^{k,i}$, respectively. We then update $\beta_{k}$ and $\gamma_{k}$ by the new estimates, i.e., $l_{11}^{k,i},l_{12}^{k,i},l_{22}^{k,i},\mu_{k,i}$. The proposed algorithm for solving nonconvex-strongly concave problems, denoted as PF-AGP-NSC, is formally presented in Algorithm 1.
Algorithm 1 A parameter-free alternating gradient projection (PF-AGP-NSC) algorithm for nonconvex-strongly concave minimax problems
  Step 1: Input $x_{1},y_{1}$, $\beta_{1}$, $\gamma_{1}$, $l_{11}^{1}$, $l_{12}^{1}$, $l_{22}^{1}$, $\mu_{1}$; Set $k=1$.
  Step 2: Updata $x_{k}$ and $y_{k}$:
   (a): Set $i=1$, $l_{11}^{k,i}=l_{11}^{k}$, $l_{12}^{k,i}=l_{12}^{k}$, $l_{22}^{k,i}=l_{22}^{k}$, $\mu_{k,i}=\mu_{k}$, $\beta_{k,i}=\beta_{k}$, $\gamma_{k,i}=\gamma_{k}$.
   (b): Update $x_{{k,i}}$ and $y_{{k,i}}$:
$x_{{k,i}}$ $=\mathcal{P}_{\mathcal{X}}\left(x_{k}-\frac{1}{\beta_{k,i}}\nabla_{x}f(x_{k},y_{k})\right),$ (3)
$y_{{k,i}}$ $=\mathcal{P}_{\mathcal{Y}}\left(y_{k}+\frac{1}{\gamma_{k,i}}\nabla_{y}f(x_{{k,i}},y_{k})\right).$ (4)
   (c): Compute $C_{1}^{k,i}$, $C_{2}^{k,i}$, $C_{3}^{k,i}$, $C_{4}^{k,i}$ as in (C1), (C2), (C3), (C4), respectively;
   (d): Update $l^{k,i}_{11}$, $l^{k,i}_{12}$, $l^{k,i}_{22}$, $\mu_{k,i}$:
$l^{k,i+1}_{11}$ $=\frac{\mbox{sgn}(C_{1}^{k,i})+3}{2}l^{k,i}_{11},\quad l^{k,i+1}_{12}=\frac{\mbox{sgn}(C_{2}^{k,i})+3}{2}l^{k,i}_{12},$ (5)
$l^{k,i+1}_{22}$ $=\frac{\mbox{sgn}(C_{3}^{k,i})+3}{2}l^{k,i}_{22},\quad\mu_{k,i+1}=\frac{2}{\mbox{sgn}(C_{4}^{k,i})+3}\mu_{k,i}.$ (6)
   (e): If $C_{1}^{k,i}\leq 0$, $C_{2}^{k,i}\leq 0$, $C_{3}^{k,i}\leq 0$ and $C_{4}^{k,i}\leq 0$, then
     $x_{k+1}=x_{k,i}$, $y_{k+1}=y_{k,i}$, $l_{11}^{k}=l_{11}^{k,i+1}$, $l_{12}^{k}=l_{12}^{k,i+1}$, $l_{22}^{k}=l_{22}^{k,i+1}$, $\mu_{k}=\mu_{k,i+1}$,
     $\beta_{k}=\beta_{k,i}$, $\gamma_{k}=\gamma_{k,i}$, go to Step 3;
     Otherwise, $i=i+1$,
$\beta_{k,i}={l_{11}^{k,i}}+\frac{l_{12}^{k,i}\mu_{k-1,i}}{8(l_{22}^{k-1,i})^{2}}+\frac{32(l_{12}^{k,i})^{2}(l_{22}^{k-1,i})^{2}}{\mu_{k,i}\mu_{k-1,i}^{2}},\quad\gamma_{k,i}=\frac{8(l_{22}^{k,i})^{2}}{\mu_{k,i}},$
     go to Step 2(b).
  Step 3: If some stationary condition is satisfied, stop; Otherwise, set $k=k+1$, go to Step 2.
It is worth noting that the parameter-free NeAda (Yang et al., 2022) and TiAda (Li et al., 2022) algorithms with adaptive step size are only used to solve the minimax problems when $x$ is unconstrained, while the proposed PF-AGP-NSC algorithm can solve the constrained case. Compared with the SGDA-B (Xu et al., 2023) algorithm and AGDA+ (Zhang et al., 2024) algorithm using the backtracking framework under the nonconvex-strongly concave setting, AGDA+ and SGDA-B still require knowledge of the strongly concave constant $\ mu$, while the proposed PF-AGP-NSC algorithm is a completely parameter-free algorithm that does not require any prior knowledge of the parameters.
2.1.1 Complexity analysis
In this subsection, we analyze the iteration complexity of Algorithm 1 for solving nonconvex-strongly concave minimax optimization problems, i.e., $f(x,y)$ is nonconvex with respect to $x$ for any fixed $y\in\mathcal{Y}$, and $\mu$-strongly concave with respect to $y$ for any given $x\in\mathcal{X}$.
Before we prove the iteration complexity of the PF-AGP-NSC algorithm for solving (P), we need to make the following assumption about the smoothness of $f(x,y)$.
Assumption 1
$f(x,y)$ has Lipschitz continuous gradients, i.e., there exist positive scalars $L_{11},L_{12},L_{22}$ such that for any $x,\tilde{x}\in\mathcal{X},y,\tilde{y}\in\mathcal{Y}$,
$\|\nabla_{x}f(x,y)-\nabla_{x}f(\tilde{x},y)\|$ $\leq L_{11}\|x-\tilde{x}\|,$
$\|\nabla_{y}f(x,y)-\nabla_{y}f(\tilde{x},y)\|$ $\leq L_{12}\|x-\tilde{x}\|,$
$\|\nabla_{y}f(x,\tilde{y})-\nabla_{y}f(x,y)\|$ $\leq L_{22}\|\tilde{y}-y\|.$
We denote ${L}=\max\{L_{11},L_{12},L_{22}\}$. To analyze the convergence of Algorithm 1, we define the stationarity gap as the termination criterion as follows.
Definition 1
At each iteration of Algorithm 1, the stationarity gap for problem (P) with respect to $f(x,y)$ is defined as:
$\nabla G(x_{k},y_{k})=\begin{bmatrix}\beta_{k}\left(x_{k}-\mathcal{P}_{\mathcal{X}}\left(x_{k}-\frac{1}{\beta_{k}}\nabla_{x}f(x_{k},y_{k})\right)\right)\\ \gamma_{k}\left(y_{k}-\mathcal{P}_{\mathcal{Y}}\left(y_{k}+\frac{1}{\gamma_{k}}\nabla_{y}f(x_{k},y_{k})\right)\right)\end{bmatrix}.$ (7)
We denote $\nabla{G}_{k}=\nabla{G}(x_{k},y_{k})$, $(\nabla G_{k})_{x}=\beta_{k}(x_{k}-\operatorname{\mathcal{P}_{\mathcal{X}}}(x_{k}-\frac{1}{\beta_{k}}\nabla_{x}f(x_{k},y_{k})))$, and $(\nabla G_{k})_{y}=\gamma_{k}(y_{k}-\operatorname{\mathcal{P}_{\mathcal{Y}}}(y_{k}+\frac{1}{\gamma_{k}}\nabla_{y}f(x_{k},y_{k})))$.
Definition 2
At each iteration of Algorithm 1, the stationarity gap for problem (P) with respect to $f_{k}(x,y)$ is defined as:
$\nabla\tilde{G}(x_{k},y_{k})=\begin{bmatrix}\beta_{k}\left(x_{k}-P_{x}\left(x_{k}-\frac{1}{\beta_{k}}\nabla_{x}f_{k}(x_{k},y_{k})\right)\right)\\ \gamma_{k}\left(y_{k}-P_{y}\left(y_{k}+\frac{1}{\gamma_{k}}\nabla_{y}f_{k}(x_{k},y_{k})\right)\right)\end{bmatrix}.$
Under Assumption 1, we first establish the following lemma, which provides bounds on the changes in the function value when $x_{k}$ is updated at each iteration of Algorithm 1. This proof is similar to that of Lemma 2.1 in (Xu et al., 2023), for the sake of completeness, we give its proof.
Lemma 3
Suppose that Assumption 1 holds. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 1, then we have
$f(x_{k+1},y_{k})-f(x_{k},y_{k})\leq-\left(\beta_{k}-\frac{l^{k}_{11}}{2}\right)\|x_{k+1}-x_{k}\|^{2}.$ (8)
Proof  By the optimality condition for (3), we have
$\langle\nabla_{x}f\left(x_{k},y_{k}\right)+\beta_{k}\left(x_{k+1}-x_{k}\right),x_{k}-x_{k+1}\rangle\geq 0.$ (9)
The backtracking strategy (C1) implies that
$f\left(x_{k+1},y_{k}\right)-f\left(x_{k},y_{k}\right)\leq\langle\nabla_{x}f\left(x_{k},y_{k}\right),x_{k+1}-x_{k}\rangle+\tfrac{l^{k}_{11}}{2}\|x_{k+1}-x_{k}\|^{2}.$ (10)
By adding (9) and (10), we obtain
$f\left(x_{k+1},y_{k}\right)-f\left(x_{k},y_{k}\right)$ $\leq-\left(\beta_{k}-\tfrac{l^{k}_{11}}{2}\right)\|x_{k+1}-x_{k}\|^{2},$
which completes the proof.
Next, we provide an estimate of upper bound of the difference between $f(x_{k+1},y_{k+1})$ and $f(x_{k},y_{k})$. First, we need to make the following assumption on the parameter $\gamma_{k}$.
Assumption 2
$\{\gamma_{k}\}$ is a nonnegative monotonically increasing sequence.
Lemma 4
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 1, then we have
$f(x_{k+1},y_{k+1})-f(x_{k},y_{k})$
$\quad\leq-\left(\beta_{k}-\frac{l_{11}^{k}}{2}-\frac{l_{12}^{k}}{2\gamma_{k-1}}\right)\|x_{k+1}-x_{k}\|^{2}+{\gamma_{k-1}}\|y_{k+1}-y_{k}\|^{2}$
$\quad\quad-\left(\mu_{k-1}-\frac{\gamma_{k-1}}{2}-\frac{(l_{22}^{k-1})^{2}}{2\gamma_{k-1}}\right)\|y_{k}-y_{k-1}\|^{2}.$ (11)
Proof  The optimality condition for $y_{k}$ in (4) implies that
$\langle\nabla_{y}f(x_{k},y_{k-1})-\gamma_{k-1}(y_{k}-y_{k-1}),y_{k+1}-y_{k}\rangle\leq 0.$ (12)
By the concavity of $f(x,y)$ with respect to $y$, and combining (12), we have
$f(x_{k+1},y_{k+1})-f(x_{k+1},y_{k})$
$\leq$ $\langle\nabla_{y}f(x_{k+1},y_{k}),y_{k+1}-y_{k}\rangle$
$\leq$ $\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle+\gamma_{k-1}\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle.$ (13)
Denoting $v_{k+1}=(y_{k+1}-y_{k})-(y_{k}-y_{k-1})$, we can write the first inner product term in the r.h.s. of (13) as
$\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle$
$=$ $\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle+\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),v_{k+1}\rangle$
$+\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle.$ (14)
Next, we estimate the three terms on the right-hand side of (14) respectively. By the backtracking strategy (C2), (C3) and the Cauchy-Schwarz inequality, we have
$\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle\leq\frac{l_{12}^{k}}{2\gamma_{k}}\|x_{k+1}-x_{k}\|^{2}+\frac{\gamma_{k}}{2}\|y_{k+1}-y_{k}\|^{2}$ (15)
and
$\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),v_{k+1}\rangle\leq\frac{(l_{22}^{k-1})^{2}}{2\gamma_{k-1}}\|y_{k}-y_{k-1}\|^{2}+\frac{\gamma_{k-1}}{2}\|v_{k+1}\|^{2}.$ (16)
By the backtracking strategy (C4) of $\mu_{k-1}$, we obtain
$\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle\leq-\mu_{k-1}\|y_{k}-y_{k-1}\|^{2}.$ (17)
Moreover, it can be easily checked that
$\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle=\frac{1}{2}\|y_{k}-y_{k-1}\|^{2}+\frac{1}{2}\|y_{k+1}-y_{k}\|^{2}-\frac{1}{2}\|v_{k+1}\|^{2}.$ (18)
Plugging (14)-(18) into (13) and rearranging the terms, we conclude that
$f(x_{k+1},y_{k+1})-f(x_{k+1},y_{k})$ $\leq\frac{l_{12}^{k}}{2\gamma_{k}}\|x_{k+1}-x_{k}\|^{2}-\left(\mu_{k-1}-\frac{\gamma_{k-1}}{2}-\frac{(l_{22}^{k-1})^{2}}{2\gamma_{k-1}}\right)\|y_{k}-y_{k-1}\|^{2}$
$\quad+\left(\frac{\gamma_{k}}{2}+\frac{\gamma_{k-1}}{2}\right)\|y_{k+1}-y_{k}\|^{2}.$ (19)
By Lemma 8, we have
$f(x_{k+1},y_{k})-f(x_{k},y_{k})\leq-(\beta_{k}-\frac{l_{11}^{k}}{2})\|x_{k+1}-x_{k}\|^{2}.$ (20)
The proof is then completed by combining (19) with (20).
We now establish an important recursion for Algorithm 1.
Lemma 5
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 1, and denoting
$f_{k+1}$ $=f(x_{k+1},y_{k+1}),\quad S_{k+1}=\frac{2\gamma^{2}_{k}}{\mu_{k}}\|y_{k+1}-y_{k}\|^{2},$
$F_{k+1}$ $=f_{k+1}+S_{k+1}-\left(\mu_{k}+\frac{7\gamma_{k}}{2}-\frac{(l^{k}_{22})^{2}}{2\gamma_{k}}-\frac{2(l^{k}_{22})^{2}}{\mu_{k}}\right)\|y_{k+1}-y_{k}\|^{2}-\frac{2\gamma_{k}^{2}}{\mu_{k}}\mathcal{D}^{2}_{y},$ (21)
then we have
$F_{k+1}-F_{k}\leq$ $-\left(\frac{2\mu_{k}\gamma_{k}-(l_{22}^{k})^{2}}{2\gamma_{k}}+\frac{\mu_{k}\gamma_{k}-4(l_{22}^{k})^{2}}{2\mu_{k}}\right)\|y_{k+1}-y_{k}\|^{2}$
$\quad-\left(\beta_{k}-\frac{l_{11}^{k}}{2}-\frac{l_{12}^{k}}{2\gamma_{k-1}}-\frac{2(l_{12}^{k})^{2}\gamma_{k-1}}{\mu_{k-1}\mu_{k}}\right)\|x_{k+1}-x_{k}\|^{2},$ (22)
where $\mathcal{D}_{y}=\max\{\|y_{1}-y_{2}\|\mid y_{1},y_{2}\in\mathcal{Y}\}$.
Proof  By the optimality condition for $y_{k}$ in (4) we obtain
$\langle\nabla_{y}f(x_{k+1},y_{k})-\gamma_{k}(y_{k+1}-y_{k}),y_{k}-y_{k+1}\rangle\leq 0.$ (23)
Combining (23) and (12), we have
$\langle\nabla_{y}f(x_{k},y_{k-1})-\nabla_{y}f(x_{k+1},y_{k})+\gamma_{k}(y_{k+1}-y_{k})-\gamma_{k-1}(y_{k}-y_{k-1}),y_{k+1}-y_{k}\rangle\leq 0,$ (24)
which implies that
$\gamma_{k}\|y_{k+1}-y_{k}\|^{2}-\gamma_{k-1}\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle\leq\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle.$ (25)
Denoting $v_{k+1}=(y_{k+1}-y_{k})-(y_{k}-y_{k-1})$, and using the identity $\frac{1}{2}\langle v_{k+1},y_{k+1}-y_{k}\rangle=\frac{1}{2}\|y_{k+1}-y_{k}\|^{2}+\frac{1}{2}\|v_{k+1}\|^{2}-\frac{1}{2}\|y_{k}-y_{k-1}\|^{2}$, we conclude that
$\gamma_{k}\|y_{k+1}-y_{k}\|^{2}-\gamma_{k-1}\left(\frac{1}{2}\|y_{k+1}-y_{k}\|^{2}+\frac{1}{2}\|y_{k}-y_{k-1}\|^{2}-\frac{1}{2}\|v_{k+1}\|^{2}\right)$
$\quad\leq\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle.$ (26)
By rearranging the terms of the above inequality, we then have
$\left(\gamma_{k}-\frac{\gamma_{k-1}}{2}\right)\|y_{k+1}-y_{k}\|^{2}+\frac{\gamma_{k-1}}{2}\|v_{k+1}\|^{2}-\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}$
$\leq$ $\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle$
$=$ $\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle+\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),v_{k+1}\rangle$
$+\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle.$ (27)
By backtracking strategy (C2) and the Cauchy-Schwarz inequality, we can easily get that
$\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle\leq\frac{(l_{12}^{k})^{2}}{2\mu_{k}}\|x_{k+1}-x_{k}\|^{2}+\frac{\mu_{k}}{2}\|y_{k+1}-y_{k}\|^{2}.$ (28)
Combining (15), (16), (27) and (28), we have
$\frac{\gamma_{k}}{2}\|y_{k+1}-y_{k}\|^{2}-\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}$ $\leq\frac{(l_{12}^{k})^{2}}{2\mu_{k}}\|x_{k+1}-x_{k}\|^{2}+\left(\frac{\mu_{k}}{2}+\frac{\gamma_{k-1}}{2}-\frac{\gamma_{k}}{2}\right)\|y_{k+1}-y_{k}\|^{2}$
$\quad-\left(\mu_{k-1}-\frac{(l_{22}^{k-1})^{2}}{2\gamma_{k-1}}\right)\|y_{k}-y_{k-1}\|^{2}.$ (29)
By multiplying $\frac{4\gamma_{k-1}}{\mu_{k-1}}$ on both sides of (29), using the definition of $S_{k+1}$ and Assumption 2, we obtain
$\begin{split}S_{k+1}-S_{k}&\leq\frac{2(l_{12}^{k})^{2}\gamma_{k-1}}{\mu_{k-1}\mu_{k}}\|x_{k+1}-x_{k}\|^{2}+2\gamma_{k-1}\|y_{k+1}-y_{k}\|^{2}\\ &\quad-\left({4\gamma_{k-1}}-\frac{2(l_{22}^{k-1})^{2}}{\mu_{k-1}}\right)\|y_{k}-y_{k-1}\|^{2}+\left(\frac{2\gamma_{k}^{2}}{\mu_{k}}-\frac{2\gamma^{2}_{k-1}}{\mu_{k-1}}\right)\mathcal{D}_{y}^{2}.\end{split}$
The proof is then completed by Lemma 4 and the definition.
We are now ready to establish the iteration complexity for the PF-AGP-NSC algorithm. In particular, letting $\nabla G(x_{k},y_{k})$ be defined as in Definition 1 and $\varepsilon>0$ be a given target accuracy, we provide a bound on $T(\varepsilon)$, the first iteration index to achieve an $\varepsilon$-stationary point, i.e., $\|\nabla G(x_{k},y_{k})\|\leq\varepsilon$, which is equivalent to
$T(\varepsilon)=\min\{k\mid\|\nabla G(x_{k},y_{k})\|\leq\varepsilon\}.$ (30)
Theorem 6
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 1, $\beta_{k}={l_{11}^{k}}+\frac{l_{12}^{k}\mu_{k-1}}{8(l_{22}^{k-1})^{2}}+\frac{32(l_{12}^{k})^{2}(l_{22}^{k-1})^{2}}{\mu_{k}\mu_{k-1}^{2}}$, $\gamma_{k}=\frac{8(l_{22}^{k})^{2}}{\mu_{k}}$, then it holds that
${T}(\varepsilon)\leq\frac{F_{2}-\underline{F}}{d_{1}\varepsilon^{2}}+1,$ (31)
where $F_{2}$ is defined in (21), $\underline{F}=\underline{f}-\left(\frac{224L_{22}^{2}}{\mu}+\mu_{1}-\frac{2(l_{22}^{1})^{2}}{\mu_{1}}-\frac{\left(l_{22}^{1}\mu\right)^{2}}{128L_{22}^{2}}+\frac{2^{13}L_{22}^{4}}{\mu^{3}}\right)\mathcal{D}^{2}_{y}$, with $\underline{f}=\min_{(x,y)\in X\times Y}f(x,y)$, $\mathcal{D}_{y}=\max\{\|y_{1}-y_{2}\|\mid y_{1},y_{2}\in\mathcal{Y}\}$, $d_{1}=\min\{\frac{\beta_{\text{min}}}{2\beta_{\text{max}}^{2}+4L^{2}_{12}},\frac{15\mu^{3}+2^{7}l_{22}^{1}\mu}{2^{15}L_{22}^{4}}\}$, $\beta_{\text{min}}={l^{1}_{11}}+\frac{l^{1}_{12}\mu}{64(L_{22})^{2}}+\frac{32(l_{12}^{1})^{2}(l^{1}_{22})^{2}}{\mu_{1}^{3}}$ and $\beta_{\text{max}}=2L_{11}+\frac{\mu_{1}L_{12}}{4(l^{1}_{22})^{2}}+\frac{4096(L_{12})^{2}(L_{22})^{2}}{\mu^{3}}$.
Proof  It follows immediately from Definition 1 and (3) that
$\|(\nabla G_{k})_{x}\|=\beta_{k}\|x_{k+1}-x_{k}\|.$ (32)
On the other hand, by the triangle inequality, we obtain that
$\|(\nabla G_{k})_{y}\|$ $\leq\gamma_{k}\left\|P_{y}\left(y_{k}+\frac{1}{\gamma_{k}}\nabla_{y}f(x_{k+1},y_{k})\right)-P_{y}\left(y_{k}+\frac{1}{\gamma_{k}}\nabla_{y}f(x_{k},y_{k})\right)\right\|$
$\quad+\gamma_{k}\|y_{k+1}-y_{k}\|$
$\leq L_{12}\|x_{k+1}-x_{k}\|+\gamma_{k}\|y_{k+1}-y_{k}\|.$ (33)
By combining (32) and (33), and using the Cauchy-Schwarz inequality, we obtain
$\|\nabla G_{k}\|^{2}$ $\leq(\beta_{k}^{2}+2L_{12}^{2})\|x_{k+1}-x_{k}\|^{2}+2\gamma_{k}^{2}\|y_{k+1}-y_{k}\|^{2}.$ (34)
Let $d^{(k)}_{1}={\min\left\{\frac{\frac{2\mu_{k}\gamma_{k}-\left(l^{k}_{22}\right)^{2}}{2\gamma_{k}}+\frac{\mu_{k}\gamma_{k}-4\left(l_{22}^{k}\right)^{2}}{2\mu_{k}}}{2\gamma_{k}^{2}},\frac{\beta_{k}-\frac{l^{k}_{11}}{2}-\frac{l^{k}_{12}}{2\gamma_{k-1}}-\frac{2(l^{k}_{12})^{2}\gamma_{k-1}}{\mu_{k-1}\mu_{k}}}{\beta_{k}^{2}+2L_{12}^{2}}\right\}}>0$. Multiplying both sides of (34) by $d^{(k)}_{1}$, and using (22) in Lemma 5, we have
$d^{(k)}_{1}\|\nabla G_{k}\|^{2}$ $\leq F_{k}-F_{k+1}.$ (35)
By using the definitions of $d_{1}$, $\beta_{k}$ and $\gamma_{k}$, we have $d_{1}<d^{(k)}_{1}$, then we get
$d_{1}\|\nabla G_{k}\|^{2}\leq F_{k}-F_{k+1}.$ (36)
Summing up the above inequalities from $k=2$ to $k=T(\varepsilon)$, we obtain
$\sum_{k=2}^{T(\varepsilon)}d_{1}\|\nabla G_{k}\|^{2}$ $\leq F_{2}-F_{T(\varepsilon)+1}.$ (37)
Note that by the definition of $F_{k+1}$ in Lemma 5, we have
$F_{T(\varepsilon)+1}$ $=f_{T(\varepsilon)+1}+S_{T(\varepsilon)+1}-\frac{2\gamma_{T(\varepsilon)}^{2}}{\mu_{T(\varepsilon)}}\mathcal{D}^{2}_{y}$
$\quad-\left(\frac{7}{2}\gamma_{T(\varepsilon)+1}+\mu_{T(\varepsilon)+1}-\frac{2(l_{22}^{T(\varepsilon)+1})^{2}}{\mu_{T(\varepsilon)+1}}-\frac{(l_{22}^{T(\varepsilon)+1})^{2}}{2\gamma_{T(\varepsilon)+1}}\right)||y_{T(\varepsilon)+1}-y_{T(\varepsilon)}||^{2}$
$\geq\underline{f}-\left(\frac{224L_{22}^{2}}{\mu}+\mu_{1}-\frac{2(l_{22}^{1})^{2}}{\mu_{1}}-\frac{\left(l_{22}^{1}\mu\right)^{2}}{128L_{22}^{2}}+\frac{2^{13}L_{22}^{4}}{\mu^{3}}\right)\mathcal{D}^{2}_{y}$
$=\underline{F},$ (38)
where the inequality follows from the definitions of $\underline{f}$ and $\mathcal{D}_{y}$, and the fact that $S_{k}\geq 0$ ($\forall k\geq 1$). We then conclude from (37) that
$\sum_{k=2}^{T(\varepsilon)}d_{1}\|\nabla G_{k}\|^{2}$ $\leq{F_{2}}-F_{T(\varepsilon)+1}\leq{F}_{2}-\underline{F},$ (39)
which, in view of the definition of $T(\varepsilon)$, implies that
$d_{1}(T(\varepsilon)-1)\varepsilon^{2}\leq F_{2}-\underline{F},$ (40)
or equivalently,
$T(\varepsilon)\leq\frac{F_{2}-\underline{F}}{d_{1}\varepsilon^{2}}+1.$ (41)
We complete the proof.
Remark 7
At the $k$-th iteration of Algorithm 1, the number of backtracking step is upper bounded by $\log_{2}\left(\frac{2L_{11}}{l_{11}^{1}}\right)+\log_{2}\left(\frac{2L_{12}}{l_{12}^{1}}\right)+\log_{2}\left(\frac{2L_{22}}{l_{22}^{1}}\right)+\log_{2}\left(\frac{2\mu_{1}}{\mu}\right)=\mathcal{O}(1)$. By Theorem 6, the total number of gradient calls of Algorithm 1 to obtain an $\varepsilon$-stationary point that satisfies $\|\nabla G(x_{k},y_{k})\|\leq\varepsilon$ is upper bounded by $\mathcal{O}(L\kappa^{3}\varepsilon^{-2})$ under the nonconvex-strongly concave setting, where $\kappa$ is the condition number.
2.2 Nonconvex-concave Setting
In this subsection, we propose a completely parameter-free AGP algorithm for solving nonconvex-concave minimax problems. Similar to the idea of the PF-AGP-NSC Algorithm, we estimate the Lipschitz constants $L_{11}$, $L_{12}$ and $L_{22}$ iteratively, however we do not need to estimate the strongly concave constant $\mu$. At each iteration of the proposed algorithm, we aim to find a tuple $(l^{k,i}_{11},l^{k,i}_{12},l^{k,i}_{22})$ by backtracking such that conditions (C1), (C2), (C3) and the following condition (C5) are satisfied at $(x_{k,i},y_{k,i})$:
$C_{5}^{k,i}$ $=\langle\nabla_{y}f_{k,i}(x_{k,i},y_{k,i})-\nabla_{y}f_{k,i}(x_{k,i},y_{k}),y_{k,i}-y_{k}\rangle+\frac{c^{2}_{k,i}+c_{k,i}{l}^{k,i}_{22}}{{l}^{k,i}_{22}+2c_{k,i}}\|y_{k,i}-y_{k}\|^{2}$
$\qquad+\frac{1}{{l}^{k,i}_{22}+2c_{k,i}}\|\nabla_{y}f_{k,i}(x_{k,i},y_{k,i})-\nabla_{y}f_{k,i}(x_{k,i},y_{k})\|^{2}\leq 0.$ (C5)
The proposed algorithm, denoted as PF-AGP-NC, is formally presented in Algorithm 2.
Please note that the existing SGDA-B algorithm (Xu et al., 2023) for solving nonconvex-concave minimax problems requires the realization of the diameter of the feasible set $\mathcal{Y}$ in order to construct the key auxiliary problem, while the PF-AGP-NC algorithm we propose is a completely parameter-free algorithm that does not require any prior parameters.
Algorithm 2 A parameter-free alternating gradient projection (PF-AGP-NC) algorithm for nonconvex-concave minimax problems
  Step 1: Input $x_{1},y_{1}$, $\beta_{1}$, $\gamma_{1}$, $c_{1}$, $l_{11}^{1}$, $l_{12}^{1}$, $l_{22}^{1}$; Set $k=1$.
  Step 2: Updata $x_{k}$ and $y_{k}$:
   (a): Set $i=1$, $l_{11}^{k,i}=l_{11}^{k}$, $l_{12}^{k,i}=l_{12}^{k}$, $l_{22}^{k,i}=l_{22}^{k}$, $\beta_{k,i}=\beta_{k}$, $\gamma_{k,i}=\gamma_{k}$, $c_{k,i}=c_{k}$.
   (b): Update $x_{{k,i}}$ and $y_{{k,i}}$:
$x_{{k,i}}$ $=\mathcal{P}_{\mathcal{X}}\left(x_{k}-\frac{1}{\beta_{k,i}}\nabla_{x}f(x_{k},y_{k})\right),$ (42)
$y_{{k,i}}$ $=\mathcal{P}_{\mathcal{Y}}\left(y_{k}+\frac{1}{\gamma_{k,i}}\nabla_{y}f(x_{{k,i}},y_{k})-\frac{1}{\gamma_{k,i}}{c_{k,i}}y_{k}\right).$ (43)
   (c): Compute $C_{1}^{k,i}$, $C_{2}^{k,i}$, $C_{3}^{k,i}$, $C_{5}^{k,i}$ as in (C1), (C2), (C3), (C5);
   (d): Update $l^{k,i}_{11}$, $l^{k,i}_{12}$, $l^{k,i}_{22}$:
$l^{k,i+1}_{11}$ $=\frac{\mbox{sgn}(C_{1}^{k,i})+3}{2}l^{k,i}_{11},\quad l^{k,i+1}_{12}=\frac{\mbox{sgn}(C_{2}^{k,i})+3}{2}l^{k,i}_{12},$ (44)
$l^{k,i+1}_{22}$ $=\frac{\max\{\mbox{sgn}(C_{3}^{k,i}),\mbox{sgn}(C_{5}^{k,i})\}+3}{2}l^{k,i}_{22}.$ (45)
   (e): If $C_{1}^{k,i}\leq 0$, $C_{2}^{k,i}\leq 0$, $C_{3}^{k,i}\leq 0$ and $C_{5}^{k,i}\leq 0$, then
     $x_{k+1}=x_{k,i}$, $y_{k+1}=y_{k,i}$, $l_{11}^{k}=l_{11}^{k,i+1}$, $l_{12}^{k}=l_{12}^{k,i+1}$, $l_{22}^{k}=l_{22}^{k,i+1}$, $\beta_{k}=\beta_{k,i}$, $\gamma_{k}=\gamma_{k,i}$,
     $c_{k}=c_{k,i}$, go to Step 3;
     Otherwise, $i=i+1$,
$\beta_{k,i}=\frac{l_{12}^{k,i}}{20l^{k-1,i}_{22}}+\frac{1280(l_{12}^{k,i})^{2}\sqrt{k}}{19^{2}l_{12}^{k-1,i}},\quad\gamma_{k,i}=20l^{k,i}_{22},\quad c_{k,i}=\frac{19l^{k,i}_{22}}{k^{1/4}},$ (46)
     go to Step 2(b).
  Step 3: If some stationary condition is satisfied, stop; Otherwise, set $k=k+1$, go to Step 2.
2.2.1 Complexity Analysis
In this subsection, we analyze the iteration complexity of Algorithm 2 for solving (P) under the general nonconvex-concave setting. The backtracking strategy (C5) is a key inequality that we will use to establish some important recursions for the PF-AGP-NC algorithm under the nonconvex-concave setting in the following two lemmas. This is also one of the main differences between proofs in the nonconvex- strongly concave setting and proofs in the nonconvex-concave setting.
Lemma 8
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 2. If $\forall k$, $\beta_{k}>l^{k}_{11}$ and $\gamma_{k}\geq\frac{{l}_{22}^{k}+2c_{k}}{2}$, then
$f({x_{k+1}},y_{k+1})-f({x_{k}},y_{k})$
$\leq$ $-\left(\frac{{\beta}_{k}}{2}-\frac{\left({l_{12}^{k}}\right)^{2}}{2\gamma_{k-1}}\right)\|x_{k+1}-x_{k}\|^{2}+{\gamma_{k-1}}\|y_{k+1}-y_{k}\|^{2}+\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}$
$\quad+\frac{c_{k-1}}{2}\left(\|y_{k+1}\|^{2}-\|y_{k}\|^{2}\right)+\frac{c_{k}}{2}\|y_{k+1}-y_{k}\|^{2}.$ (47)
Proof  The optimality condition for $y_{k}$ in (43) implies that
$\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\gamma_{k-1}(y_{k}-y_{k-1}),y_{k}-y_{k+1}\rangle\leq 0.$ (48)
The concavity of $f_{k}(x_{k+1},y)$ with respect to $y$ together with (48) then imply that
$f_{k}(x_{k+1},y_{k+1})-f_{k}(x_{k+1},y_{k})$ $\leq\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k}),y_{k+1}-y_{k}\rangle$
$\quad+\langle\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle$
$=\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k}),y_{k+1}-y_{k}\rangle$
$\quad+\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),v_{k+1}\rangle$
$\quad+\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle$
$\quad+\gamma_{k-1}\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle,$ (49)
where $v_{k+1}=y_{k+1}-y_{k}-(y_{k}-y_{k-1})$. We now provide upper bounds for the inner product terms on the right hand side of (49). Firstly, by the definition of $f_{k}(x_{k+1},y_{k})$ and $f_{k-1}(x_{k},y_{k})$, the backtracking strategy (C2) and the Cauchy-Schwarz inequality, we have
$\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k}),y_{k+1}-y_{k}\rangle$
$=\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle-(c_{k}-c_{k-1})\langle y_{k},y_{k+1}-y_{k}\rangle$
$\leq\frac{(l_{12}^{k})^{2}}{2\gamma_{k-1}}\|x_{k+1}-x_{k}\|^{2}+\frac{\gamma_{k-1}}{2}\|y_{k+1}-y_{k}\|^{2}-\frac{c_{k}-c_{k-1}}{2}(\|y_{k+1}\|^{2}-\|y_{k}\|^{2})$
$\quad+\frac{c_{k}-c_{k-1}}{2}\|y_{k+1}-y_{k}\|^{2}$
$\leq\frac{(l_{12}^{k})^{2}}{2\gamma_{k-1}}\|x_{k+1}-x_{k}\|^{2}+\frac{\gamma_{k-1}}{2}\|y_{k+1}-y_{k}\|^{2}-\frac{c_{k}-c_{k-1}}{2}(\|y_{k+1}\|^{2}-\|y_{k}\|^{2})$
$\quad+\frac{c_{k}}{2}\|y_{k+1}-y_{k}\|^{2}.$ (50)
Secondly, by the Cauchy-Schwarz inequality, we get
$\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),v_{k+1}\rangle$
$\leq$ $\frac{1}{2\gamma_{k-1}}\|\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1})\|^{2}+\frac{\gamma_{k-1}}{2}\|v_{k+1}\|^{2}.$ (51)
Thirdly, by (C4) we have that
$\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle$
$\leq$ $-\frac{1}{{l}_{22}^{k-1}+2c_{k-1}}\|\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1})\|^{2}$
$\quad-\frac{c_{k-1}{l}_{22}^{k-1}+c^{2}_{k-1}}{{l}_{22}^{k-1}+2c_{k-1}}\|y_{k}-y_{k-1}\|^{2}$
$\leq$ $-\frac{1}{{l}_{22}^{k-1}+2c_{k-1}}\|\nabla_{y}f_{k}(x_{k},y_{k})-\nabla_{y}f_{k}(x_{k},y_{k-1})\|^{2}.$ (52)
We also have the following equation that
$\begin{split}\gamma_{k-1}\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle=\frac{\gamma_{k-1}}{2}\|y_{k+1}-y_{k}\|^{2}+\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}-\frac{\gamma_{k-1}}{2}\|v_{k+1}\|^{2}.\end{split}$ (53)
By $\gamma_{k}\geq\frac{{l}_{22}^{k}+2c_{k}}{2}$, plugging (50)-(53) into (49), and using the definition of $f_{k}(x_{k+1},y_{k+1})$ and $f_{k}(x_{k+1},y_{k})$, we obtain
$f(x_{k+1},y_{k+1})-f(x_{k+1},y_{k})$
$\leq$ $\frac{(l_{12}^{k})^{2}}{2\gamma_{k-1}}\|x_{k+1}-x_{k}\|^{2}+\gamma_{k-1}\|y_{k+1}-y_{k}\|^{2}+\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}$
$\quad+\frac{c_{k-1}}{2}\left(\|y_{k+1}\|^{2}-\|y_{k}\|^{2}\right)+\frac{c_{k}}{2}\|y_{k+1}-y_{k}\|^{2}.$ (54)
By $\beta_{k}>l^{k}_{11}$, we have
$f(x_{k+1},y_{k})-f(x_{k},y_{k})\leq-\frac{\beta_{k}}{2}\|x_{k+1}-x_{k}\|^{2}.$ (55)
By adding (54) and (55), we complete the proof.
Lemma 9
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 2. Denote
$S_{k+1}$ $=\frac{8\gamma_{k}^{2}}{c_{k}}\|y_{k+1}-y_{k}\|^{2}+{8}{\gamma_{k}}\left(\frac{c_{k+1}}{c_{k}}-1\right)\|y_{k+1}\|^{2},$ (56)
$F_{k+1}$ $={f(x_{k+1},y_{k+1})}+S_{k+1}-\frac{15\gamma_{k}}{2}\|y_{k+1}-y_{k}\|^{2}-\frac{c_{k}}{2}\|y_{k+1}\|^{2}.$ (57)
If
$\beta_{k}>l^{k}_{11},\quad\frac{\gamma_{k+1}}{c_{k+1}}-\frac{\gamma_{k}}{c_{k}}\leq\frac{1}{5},\quad\gamma_{k}\geq\frac{{l}^{k}_{22}+2c_{k}}{2},$ (58)
then $\forall k\geq 2$,
$F_{k+1}-F_{k}=$ $-\left(\frac{9\gamma_{k}}{10}-\frac{c_{k}}{2}\right)\|y_{k+1}-y_{k}\|^{2}$
$-\left(+\frac{\beta_{k}}{2}-\frac{(l^{k}_{12})^{2}}{2\gamma_{k-1}}-\frac{16(l^{k}_{12})^{2}\gamma_{k-1}}{c_{k-1}^{2}}\right)\|x_{k+1}-x_{k}\|^{2}$
$+8\left({\gamma_{k}}\frac{c_{k+1}}{c_{k}}-{\gamma_{k-1}}\frac{c_{k}}{c_{k-1}}\right)\|y_{k+1}\|^{2}+\frac{c_{k-1}-c_{k}}{2}\|y_{k+1}\|^{2}.$ (59)
Proof  The optimality condition for $y_{k}$ in (43) implies that
$\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\gamma_{k}(y_{k+1}-y_{k}),y_{k}-y_{k+1}\rangle\leq 0,$ (60)
By (48) and (60), we have
$\langle\nabla_{y}f_{k-1}(x_{k},y_{k-1})-\nabla_{y}f_{k}(x_{k+1},y_{k})+$ $\gamma_{k}(y_{k+1}-y_{k})-\gamma_{k-1}(y_{k}-y_{k-1}),y_{k+1}-y_{k}\rangle\leq 0,$
and
$\gamma_{k}\|y_{k+1}-y_{k}\|^{2}-\gamma_{k-1}\langle y_{k+1}-y_{k},y_{k}-y_{k-1}\rangle$
$\leq$ $\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle,$ (61)
which can be rewritten as
$\left(\gamma_{k}-\frac{\gamma_{k-1}}{2}\right)\|y_{k+1}-y_{k}\|^{2}+\frac{\gamma_{k-1}}{2}\|v_{k+1}\|^{2}-\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}$
$\leq$ $\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle$
$=$ $\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k}),y_{k+1}-y_{k}\rangle+\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),v_{k+1}\rangle$
$+\langle\nabla_{y}f_{k}(x_{k-1},y_{k-1})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle.$ (62)
Using an argument similar to the proof of (50)-(53), by (C5) and the Cauchy-Schwarz inequality, we conclude from the above inequality that
$\left(\gamma_{k}-\frac{\gamma_{k-1}}{2}\right)\|y_{k+1}-y_{k}\|^{2}+\frac{\gamma_{k-1}}{2}\|v_{k+1}\|^{2}-\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}$
$\leq$ $\;\frac{(l_{12}^{k})^{2}}{2a_{k}}\|x_{k+1}-x_{k}\|^{2}+\frac{a_{k}}{2}\|y_{k+1}-y_{k}\|^{2}-\frac{c_{k}-c_{k-1}}{2}(\|y_{k+1}\|^{2}-\|y_{k}\|^{2})$
$+\frac{c_{k}-c_{k-1}}{2}\|y_{k+1}-y_{k}\|^{2}+\frac{1}{2\gamma_{k-1}}\|\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1})\|^{2}$
$+\frac{\gamma_{k-1}}{2}\|v_{k+1}\|-\frac{1}{{l}_{22}^{k-1}+2c_{k-1}}\|\nabla_{y}f_{k}(x_{k},y_{k})-\nabla_{y}f_{k}(x_{k},y_{k-1})\|^{2}$
$-\frac{c_{k-1}{l}_{22}^{k-1}+c^{2}_{k-1}}{{l}_{22}^{k-1}+2c_{k-1}}\|y_{k}-y_{k-1}\|^{2},$ (63)
for any $a_{k}>0$. Obviously, we have $-\frac{c_{k-1}{l}^{k-1}_{22}+c^{2}_{k-1}}{2c_{k-1}+{l}^{k-1}_{22}}\leq-\frac{c_{k-1}{l}^{k-1}_{22}+c^{2}_{k-1}}{2({l}^{k-1}_{22}+c_{k-1})}=-\frac{c_{k-1}}{2}.$ Combining $\gamma_{k}\geq\frac{{l}^{k}_{22}+2c_{k}}{2}$ and rearranging the terms in (63), we obtain
$\frac{\gamma_{k}}{2}\|y_{k+1}-y_{k}\|^{2}+\frac{c_{k}-c_{k-1}}{2}\|y_{k+1}\|^{2}$
$\quad\leq\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}+\left(\frac{\alpha_{k}}{2}+\frac{\gamma_{k-1}}{2}-\frac{\gamma_{k}}{2}\right)\|y_{k+1}-y_{k}\|^{2}+\frac{c_{k}-c_{k-1}}{2}\|y_{k}\|^{2}$
$\quad\quad+\frac{(l^{k}_{12})^{2}}{2a_{k}}\|x_{k+1}-x_{k}\|^{2}-\frac{c_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}.$ (64)
By multiplying $\frac{16\gamma_{k-1}}{c_{k-1}}$ on both sides of the above inequality, setting $a_{k}=\frac{c_{k}}{2}$, and using the definition of $S_{k+1}$, (58) and $\{\gamma_{k}\}$ is nodecreasing, we then obtain
$S_{k+1}-S_{k}$ $\leq 8\left({\gamma_{k}}\frac{c_{k+1}}{c_{k}}-{\gamma_{k-1}}\frac{c_{k}}{c_{k-1}}\right)\|y_{k+1}\|^{2}+\frac{16(l^{k}_{12})^{2}\gamma_{k-1}}{c_{k-1}^{2}}\|x_{k+1}-x_{k}\|^{2}$
$\quad+\frac{28\gamma_{k}}{5}\|y_{k+1}-y_{k}\|^{2}-{8}{\gamma_{k-1}}\|y_{k}-y_{k-1}\|^{2}.$ (65)
Combining (47) and (65), and using the definition of $F_{k+1}$, we complete the proof.
We are now ready to establish the iteration complexity for the PF-AGP-NC algorithm to achieve an $\varepsilon$-stationary point under the general nonconvex–concave setting.
Theorem 10
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 2. If $\beta_{k}=\frac{l_{12}^{k}}{20l^{k-1}_{22}}+\frac{1280(l_{12}^{k})^{2}\sqrt{k}}{19^{2}l_{12}^{k-1}}$, $\gamma_{k}=20l^{k}_{22}$ , $c_{k}=\frac{19l^{k}_{22}}{k^{1/4}}$, then for any given $\varepsilon>0$, $k\geq 2$, we have
$T(\varepsilon)\leq\max\left(\left(\frac{5120\left(L_{12}\right)^{2}d_{2}d_{3}}{19^{2}l_{22}^{1}\varepsilon^{2}}+1\right)^{2},\frac{76^{4}L_{22}^{4}\hat{\mathcal{D}}_{y}^{4}}{\varepsilon^{4}}\right),$
where $d_{2}=F_{2}-\underline{F}+339L_{22}\hat{\mathcal{D}}_{y}^{2}$, $d_{3}=\max\{\bar{d}_{1},\frac{20\cdot 19^{2}L_{22}^{2}}{17(l_{12}^{1})^{2}}\}$, $\hat{\mathcal{D}}_{y}=\max\{\|y\|\,|y\in Y\}$ with $\underline{F}={\underline{f}}-639L_{22}\hat{\mathcal{D}}_{y}^{2}$, $\underline{f}=\min_{(x,y)\in X\times Y}f(x,y)$, $\bar{d}_{1}=32+\frac{1+1600L^{2}_{22}}{128({l}_{12}^{1})^{2}}$.
Proof  We can easily see that the relations in (58) are satisfied by the settings of $c_{k}$, $\gamma_{k}$ and $\beta_{k}$. Denote $\alpha_{k}=\frac{16(l_{12}^{k})^{2}\gamma_{k-1}}{c_{k-1}^{2}}$. Then it follows from the selection of $\beta_{k}$ and $\alpha_{k}$ that
$\frac{\beta_{k}}{2}-\frac{(l^{k}_{12})^{2}}{2\gamma_{k-1}}-\frac{16(l^{k}_{12})^{2}\gamma_{k-1}}{c_{k-1}^{2}}=\alpha_{k}.$ (66)
This observation, in view of Lemma 8, $\frac{c_{k}}{2}\leq\frac{19\gamma_{k}}{40k^{1/4}}\leq\frac{19\gamma_{k}}{40}$ implies that
$\begin{split}&{\alpha_{k}}\left\|x_{k+1}-x_{k}\right\|^{2}+\frac{17\gamma_{k}}{40}\left\|y_{k+1}-y_{k}\right\|^{2}\\ \leq&F_{k}-F_{k+1}+8\left({\gamma_{k}}\frac{c_{k+1}}{c_{k}}-{\gamma_{k-1}}\frac{c_{k}}{c_{k-1}}\right)\|y_{k+1}\|^{2}+\frac{c_{k-1}-c_{k}}{2}\|y_{k+1}\|^{2}.\end{split}$ (67)
We can easily check from the definition of $f_{k}(x,y)$ that
$\|\nabla G_{k}\|-\|\tilde{\nabla}G_{k}\|\leq c_{k}\|y_{k}\|.$ (68)
By replacing $f$ with $f_{k}$, similar to (32) and (33), we immediately obtain that
(~$\|(\tilde{\nabla}G_{k})_{x}\|=\beta_{k}\|x_{k+1}-x_{k}\|,$ (69)
and
(~$\|(\tilde{\nabla}G_{k})_{y}\|\leq\gamma_{k}\|y_{k+1}-y_{k}\|+l^{k}_{12}\|x_{k+1}-x_{k}\|.$ (70)
Combining (69) and (70), and using the Cauchy-Schwarz inequality, we have
~$\|\tilde{\nabla}G_{k}\|^{2}\leq(\beta_{k}^{2}+2(l^{k}_{12})^{2})\|x_{k+1}-x_{k}\|^{2}+2{{\gamma_{k}}^{2}}\|y_{k+1}-y_{k}\|^{2}.$ (71)
Since both $\alpha_{k}$ and $\beta_{k}$ are in the same order when $k$ becomes large enough, it then follows from the definition of $\bar{d}_{1}$ that $\forall k\geq 1$,
$\frac{(\beta_{k})^{2}+2({l^{k}_{12}})^{2}}{\alpha_{k}^{2}}$ $=\frac{\left(\frac{{l^{k}_{12}}}{\gamma_{k-1}}+\frac{64({l}_{12}^{k})^{2}\gamma_{k-1}}{c_{k-1}^{2}}\right)^{2}+2({l}^{k}_{12})^{2}}{\alpha_{k}^{2}}$
$\leq 32+\frac{1+1600L^{2}_{22}}{128({l}_{12}^{1})^{2}}=\bar{d}_{1}.$ (72)
Combining the previous two inequalities in (71) and (72), we obtain
$\|\nabla{\tilde{G}_{k}}\|^{2}\leq{\bar{d}_{1}(\alpha_{k})^{2}}\|x_{k+1}-x_{k}\|^{2}+2\gamma_{k}^{2}\|y_{k+1}-y_{k}\|^{2}.$ (73)
Denote $d_{k}^{(2)}=\frac{1}{\max\{\bar{d}_{1}\alpha_{k},{\frac{17}{80}\gamma_{k}}\}}$, $\tilde{c}_{k}=\frac{19}{20k^{1/4}}$, by multiplying $d^{(2)}_{k}$ on the both sides of (73), and using (67), we have
$d_{k}^{(2)}\|\nabla\tilde{G}_{k}\|^{2}$
$\leq$ $F_{k}-F_{k+1}+8\left({\gamma_{k}}\frac{c_{k+1}}{c_{k}}-{\gamma_{k-1}}\frac{c_{k}}{c_{k-1}}\right)\|y_{k+1}\|^{2}+\frac{c_{k-1}-c_{k}}{2}\|y_{k+1}\|^{2}$
$\leq$ $F_{k}-F_{k+1}+8\left({\gamma_{k}}\frac{c_{k+1}}{c_{k}}-{\gamma_{k-1}}\frac{c_{k}}{c_{k-1}}\right)\|y_{k+1}\|^{2}+40L_{22}\frac{\tilde{c}_{k-1}-\tilde{c}_{k}}{2}\|y_{k+1}\|^{2}.$ (74)
Denoting
$\tilde{T}(\varepsilon)=\min\left\{k\middle|\left\|\nabla\tilde{G}(x_{k},y_{k})\right\|\leq\frac{\varepsilon}{2},k\geq 2\right\},$ (75)
summing both sides of (74) from $k=2$ to $k=\tilde{T}(\varepsilon)$, we then obtain
$\sum_{k=2}^{\tilde{T}{(\varepsilon)}}d^{(2)}_{k}\left\|\nabla\tilde{G}_{k}\right\|^{2}$ $\leq F_{2}-F_{\tilde{T}(\varepsilon)}+8\left(\frac{\gamma_{\tilde{T}(\varepsilon)-1}c_{\tilde{T}(\varepsilon)}}{c_{\tilde{T}(\varepsilon)-1}}-\frac{\gamma_{1}c_{2}}{c_{1}}\right)\hat{\mathcal{D}}_{y}^{2}+40L_{22}\left(\frac{\tilde{c}_{1}}{2}-\frac{\tilde{c}_{\tilde{T}(\varepsilon)}}{2}\right)\hat{\mathcal{D}}^{2}_{y}$
$\leq F_{2}-F_{\tilde{T}(\varepsilon)}+339L_{22}\hat{\mathcal{D}}_{y}^{2}.$ (76)
Note that by the definition of $F_{k+1}$ in Lemma 9, we have
$\begin{split}F_{\tilde{T}(\varepsilon)}&\geq{\underline{f}}-639L_{22}\hat{\mathcal{D}}_{y}^{2}=\underline{F},\end{split}$ (77)
where $\underline{f}=\min_{(x,y)\in X\times Y}f(x,y)$. We then conclude from (76) that
$\sum_{k=2}^{\infty}d^{(2)}_{k}\left\|\nabla\tilde{G}_{k}\right\|^{2}\leq F_{2}-\underline{F}+339L_{22}\hat{\mathcal{D}}_{y}^{2}=d_{2}.$ (78)
We can see from the selection of $d_{3}$ that $d_{3}\geq\max\{\bar{d}_{1},\frac{80\gamma_{k}}{17\alpha_{k}}\}$, which implies $d^{(2)}_{k}\geq\frac{1}{d_{3}\alpha_{k}}$, by multiplying $d_{3}$ on the both sides of (78), and combining the definition of $d_{2}$, we have $\sum_{k=2}^{\infty}\frac{1}{\alpha_{k}}\left\|\nabla\tilde{G}_{k}\right\|^{2}\leq d_{2}d_{3}$, which, by the definition of $\tilde{T}(\varepsilon)$, implies that
$\frac{\varepsilon^{2}}{4}\leq\frac{d_{2}d_{3}}{\sum_{k=2}^{\tilde{T}{(\varepsilon)}}\frac{1}{\alpha_{k}}}\leq\frac{d_{2}d_{3}}{\sum_{k=2}^{\tilde{T}{(\varepsilon)}}\frac{1}{\bar{\alpha}}}.$ (79)
By using the fact $\sum_{k=2}^{\tilde{T}{(\varepsilon)}}\frac{1}{\sqrt{k}}\geq\sqrt{\tilde{T}{(\varepsilon)}}-1$ and (79), we conclude $\frac{\varepsilon^{2}}{4}\leq\frac{1280\cdot\left(L_{12}\right)^{2}d_{2}d_{3}}{19^{2}l_{22}^{1}\left(\sqrt{\tilde{T}(\varepsilon)}-1\right)}$ or equivalently,
$\tilde{T}(\varepsilon)\leq\left(\frac{5120\left(L_{12}\right)^{2}d_{2}d_{3}}{19^{2}l_{22}^{1}\varepsilon^{2}}+1\right)^{2}.$ (80)
On the other hand, if $k\geq\frac{76^{4}L_{22}^{4}\hat{\mathcal{D}}_{y}^{4}}{\varepsilon^{4}}$, then $c_{k}=\frac{19\gamma_{k}}{20k^{1/4}}\leq\frac{\varepsilon}{2\hat{{\mathcal{D}_{y}}}}$. This inequality together with the definition of $\hat{\mathcal{D}}_{y}$, then imply that $c_{k}\left\|y_{k}\right\|\leq\frac{\varepsilon}{2}$. Therefore, there exists a
$T(\varepsilon)$ $\leq\max\left(\tilde{T}{(\varepsilon)},\frac{76^{4}L_{22}^{4}\hat{\mathcal{D}}_{y}^{4}}{\varepsilon^{4}}\right)$
$\leq\max\left(\left(\frac{5120\left(L_{12}\right)^{2}d_{2}d_{3}}{19^{2}l_{22}^{1}\varepsilon^{2}}+1\right)^{2},\frac{76^{4}L_{22}^{4}\hat{\mathcal{D}}_{y}^{4}}{\varepsilon^{4}}\right),$ (81)
such that $\left\|\nabla G_{k}\right\|\leq\left\|\nabla\tilde{G}_{k}\right\|+c_{k}\left\|y_{k}\right\|\leq\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon$.
Remark 11
At the $k$-th iteration of Algorithm 2, the number of backtracking step is upper bounded by $\log_{2}\left(\frac{2L_{11}}{l_{11}^{1}}\right)+\log_{2}\left(\frac{2L_{12}}{l_{12}^{1}}\right)+\log_{2}\left(\frac{2L_{22}}{l_{22}^{1}}\right)=\mathcal{O}(1)$. According to Theorem 10, the total number of gradient calls of Algorithm 2 to obtain an $\varepsilon$-stationary point that satisfies $\|\nabla G(x_{k},y_{k})\|\leq\varepsilon$ is $\mathcal{O}(L^{4}\varepsilon^{-4})$ under the nonconvex-concave setting.
3 Numerical results
In this section, we compare the numerical performance of the proposed PF-AGP-NC algorithm with the AGP algorithm.
3.1 Dirac-GAN problem
The Dirac-GAN problem (Mescheder et al., 2018) can be formulated as the following nonconvex–concave minimax problem :
$\min_{x}\max_{y}L(x,y)=-\log(1+\exp(-xy))+\log 2,$ (82)
where $(0,0)$ is the unique stationary point.
Experimental setup. We set the initial point of the experiment at $(x,y)=(1,1)$. The termination condition is defined as the norm of the gradient $\|\nabla f(x_{i},y_{i})\|$ being less than or equal to $10^{-5}$, where $i$ represents the number of gradient calls.
For the AGP algorithm, let $\alpha_{x}$ and $\beta_{y}$ be the stepsize of $x$ and $y$, respectively. We set $\alpha_{x}=\frac{0.8}{\sqrt{k}}$, $\beta_{y}=0.3$ and $c_{k}=\frac{0.5}{k^{1/4}}$. For the PF-AGP-NC algorithm, we do not need to preset specific parameters.
In the experiments, we counted the number of gradient calls of the PF-AGP-NC algorithm. The results of alternating gradient projection in the $k$ th iteration are represented by nodes in the graph and the number of backtracking step computations is represented by the connecting lines.
Results Figure 1 shows the convergence curves of the AGP and PF-AGP-NC algorithms. It can be seen that although the PF-AGP-NC algorithm involves a double-layer structure, requiring the determination of points in the outer layer through computations in the inner layer, even when considering the number of iterations in the inner layer, the overall convergence speed is still faster than that of the AGP algorithm.
Figure 1: number of gradient calls $i$ vs $\log(||\nabla f(x_{i},y_{i})||)$
4 Conclusions and discussion
In this paper, based on the framework of backtracking, we propose two completely parameter-free alternating gradient projection algorithms, i.e., the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm, for solving nonconvex-(strongly) concave minimax problems respectively, which does not require prior knowledge of parameters such as the Lipschtiz constant $L$ or the strongly concave constant $\mu$. Moreover, we show that the total number of gradient calls of the PF-AGP-NSC algorithm to obtain an $\varepsilon$-stationary point for nonconvex-strongly concave minimax problems is upper bounded by $\mathcal{O}\left(L\kappa^{3}\varepsilon^{-2}\right)$ where $\kappa$ is the condition number, while the total number of gradient calls of the PF-AGP-NC algorithm to obtain an $\varepsilon$-stationary point for nonconvex-concave minimax problems is upper bounded by $\mathcal{O}\left(L^{4}\varepsilon^{-4}\right)$. As far as we know, the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm are the first completely parameter-free algorithms for solving nonconvex-strongly concave minimax problems and nonconvex-concave minimax problems respectively. Numerical results demonstrate the efficiency of the two proposed algorithms.

### v2

@@
 However, many existing algorithms require prior knowledge of the problem parameters in order to achieve optimal iteration complexity.
-In this paper, we propose a completely parameter-free alternating gradient projection (PF-AGP) algorithm to solve the smooth nonconvex-(strongly) concave minimax problems using a backtracking strategy, which does not require prior knowledge of parameters such as the Lipschtiz constant $L$ or the strongly concave constant $\mu$.
-The PF-AGP algorithm utilizes a parameter-free gradient projection step to alternately update the outer and inner variables in each iteration.
-We show that the total number of gradient calls of the PF-AGP algorithm to obtain an $\varepsilon$-stationary point for nonconvex-strongly concave minimax problems is upper bounded by $\mathcal{O}\left(L\kappa^{3}\varepsilon^{-2}\right)$ where $\kappa$ is the condition number, while the total number of gradient calls to obtain an $\varepsilon$-stationary point for nonconvex-concave minimax problems is upper bounded by $\mathcal{O}\left(L^{4}\varepsilon^{-4}\right)$.
-As far as we know, this is the first completely parameter-free algorithm for solving nonconvex-strongly concave minimax problems, and it is also the completely parameter-free algorithm which achieves the best iteration complexity in single loop method for solving nonconvex-concave minimax problems.
+In this paper, we propose two completely parameter-free alternating gradient projection algorithms, i.e., the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm, to solve the smooth nonconvex-strongly concave and nonconvex-concave minimax problems respectively using a backtracking strategy, which does not require prior knowledge of parameters such as the Lipschtiz constant $L$ or the strongly concave constant $\mu$.
+Moreover, we show that the total number of gradient calls of the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm to obtain an $\varepsilon$-stationary point is upper bounded by $\mathcal{O}\left(L\kappa^{3}\varepsilon^{-2}\right)$ and $\mathcal{O}\left(L^{4}\varepsilon^{-4}\right)$ respectively, where $\kappa$ is the condition number.
+As far as we know, the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm are the first completely parameter-free algorithms for solving nonconvex-strongly concave minimax problems and nonconvex-concave minimax problems respectively.
 Numerical results validate the efficiency of the proposed PF-AGP algorithm.
@@
 where $\mathcal{X}\subseteq\mathbb{R}^{n}$ and $\mathcal{Y}\subseteq\mathbb{R}^{m}$ are nonempty closed and bounded convex sets, and $f:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ is a smooth function.
-The minimax problem is a critical class of optimization problems, known for its extensive research history and broad applications in areas such as machine learning, signal processing, and various other fields in recent years.
-Many practical problems can be formulated as in (P), including machine learning, such as generative adversarial networks (GANs) (Arjovsky et al., 2017; Goodfellow et al., 2014; Sanjabi et al., 2018), reinforcement learning (Dai et al., 2018), power control and transceiver design in signal processing (Lu et al., 2020), distributed nonconvex optimization (Giannakis et al., 2017; Giordano et al., 2018; Mateos et al., 2010), robust optimization (Ben and Tal, 2009; Gao and Kleywegt, 2023), statistical learning (Giordano et al., 2018; Shafieezadeh Abadeh et al., 2015) and several other applications.
+Minimax problem is a critical class of optimization problems, known for its extensive research history and broad applications in areas such as machine learning, signal processing, and various other fields in recent years.
+Many practical problems can be formulated as in (P), including generative adversarial networks (GANs) (Arjovsky et al., 2017; Goodfellow et al., 2014; Sanjabi et al., 2018), reinforcement learning (Dai et al., 2018), power control and transceiver design in signal processing (Lu et al., 2020), distributed nonconvex optimization (Giannakis et al., 2017; Giordano et al., 2018; Mateos et al., 2010), robust optimization (Ben and Tal, 2009; Gao and Kleywegt, 2023), statistical learning (Giordano et al., 2018; Shafieezadeh Abadeh et al., 2015) and several other applications.
 There are many research results on convex-concave minimax problems, i.e., $f(x,y)$ is convex with respect to $x$ and concave with respect to $y$.
@@
 In Sect.
-2, we propose a parameter-free alternating gradient projection (PF-AGP) algorithm for nonconvex-(strongly) concave minimax problems, and we then analyze the corresponding gradient complexity for two different settings.
+2, we propose two parameter-free alternating gradient projection (PF-AGP) algorithms for nonconvex-strongly concave and nonconvex-concave minimax problems, and we then analyze their corresponding gradient complexity under two different settings.
 We report some numerical results in Sect.
-3 and make some concluding remarks in the last section.
+3 and make some conclusions in the last section.
 […]
@@
 It is worth noting that the parameter-free NeAda (Yang et al., 2022) and TiAda (Li et al., 2022) algorithms with adaptive step size are only used to solve the minimax problems when $x$ is unconstrained, while the proposed PF-AGP-NSC algorithm can solve the constrained case.
-Compared with the SGDA-B (Xu et al., 2023) algorithm and AGDA+ (Zhang et al., 2024) algorithm using the backtracking framework under the nonconvex-strongly concave setting, AGDA+ and SGDA-B still require knowledge of the strongly concave constant $\ mu$, while the proposed PF-AGP-NSC algorithm is a completely parameter-free algorithm that does not require any prior knowledge of the parameters.
+Compared with the SGDA-B (Xu et al., 2023) algorithm and AGDA+ (Zhang et al., 2024) algorithm using the backtracking framework under the nonconvex-strongly concave setting, AGDA+ and SGDA-B still require knowledge of the strongly concave constant $\mu$, while the proposed PF-AGP-NSC algorithm is a completely parameter-free algorithm that does not require any prior knowledge of the parameters.
 2.1.1 Complexity analysis
-In this subsection, we analyze the iteration complexity of Algorithm 1 for solving nonconvex-strongly concave minimax optimization problems, i.e., $f(x,y)$ is nonconvex with respect to $x$ for any fixed $y\in\mathcal{Y}$, and $\mu$-strongly concave with respect to $y$ for any given $x\in\mathcal{X}$.
-Before we prove the iteration complexity of the PF-AGP-NSC algorithm for solving (P), we need to make the following assumption about the smoothness of $f(x,y)$.
+In this subsection, we analyze the iteration complexity of Algorithm 1 for solving nonconvex-strongly concave minimax optimization problems (P), i.e., $f(x,y)$ is nonconvex with respect to $x$ for any fixed $y\in\mathcal{Y}$, and $\mu$-strongly concave with respect to $y$ for any given $x\in\mathcal{X}$.
+We first need to make the following assumption about the smoothness of $f(x,y)$.
 Assumption 1
@@
 3 Numerical results
-In this section, we compare the numerical performance of the proposed PF-AGP-NC algorithm with the AGP algorithm.
+In this section, we report some numerical results to show the efficiency of the proposed PF-AGP-NSC and PF-AGP-NC algorithms.
+All the numerical tests are implemented in Python 3.12 and run with an NVIDIA RTX 4090D GPU (24GB).
+3.1 Synthetic Minimax Problem
+Synthetic minimax problem can be formulated as the following nonconvex-strongly-concave minimax problem (Luo et al., 2022):
+$\min_{\mathbf{x}\in\mathbb{R}^{3}}\max_{\mathbf{y}\in\mathbb{R}^{2}}f(\mathbf{x},\mathbf{y})=w(x_{3})-\frac{y_{1}^{2}}{40}+x_{1}y_{1}-\frac{5y_{2}^{2}}{2}+x_{2}y_{2},$ (82)
+where $\mathbf{x}=[x_{1},x_{2},x_{3}]^{\top}$, $\mathbf{y}=[y_{1},y_{2}]^{\top}$ and
+$w(x)=\begin{cases}\sqrt{\epsilon}\left(x+(\lambda+1)\sqrt{\epsilon}\right)^{2}-\frac{1}{3}\left(x+(\lambda+1)\sqrt{\epsilon}\right)^{3}-\frac{1}{3}(3\lambda+1)\epsilon^{3/2},&x\leq-\lambda\sqrt{\epsilon};\\ \epsilon x+\frac{\epsilon^{3/2}}{3},&-\lambda\sqrt{\epsilon}<x\leq-\sqrt{\epsilon};\\ -\sqrt{\epsilon}x^{2}-\frac{x^{3}}{3},&-\sqrt{\epsilon}<x\leq 0;\\ -\sqrt{\epsilon}x^{2}+\frac{x^{3}}{3},&0<x\leq\sqrt{\epsilon};\\ -\epsilon x+\frac{\epsilon^{3/2}}{3},&\sqrt{\epsilon}<x\leq\lambda\sqrt{\epsilon};\\ \sqrt{\epsilon}\left(x-(\lambda+1)\sqrt{\epsilon}\right)^{2}+\frac{1}{3}\left(x-(\lambda+1)\sqrt{\epsilon}\right)^{3}-\frac{1}{3}(3\lambda+1)\epsilon^{3/2},&\lambda\sqrt{\epsilon}\leq x.\end{cases}$
+It is easy to verify that this problem has a strict saddle point at $(\mathbf{x}^{*},\mathbf{y}^{*})=([0,0,0]^{\top},[0,0]^{\top})$.
+Experimental setup.
+We compare the proposed PF-AGP-NSC algorithm with the SGDA-B algorithm, the TiAda algorithm and the AGP algorithm for solving (82) with $\epsilon=0.01$, $\lambda=5$.
+The initial point is chosen as $(x_{0},y_{0})=([0,0,2]^{\top},[0,0]^{\top})$ for all four tested algorithms.
+For the PF-AGP-NSC algorithm, we set $l_{11}^{1}=0.01$ , $l_{12}^{1}=0.01$, $l_{22}^{1}=0.01$ and $\mu_{1}=1$.
+For the SGDA-B algorithm, we set $\mu=1$, $\gamma=0.9$, $N=1$, $p=0.5$ and $\tilde{L}=\frac{10}{9}$.
+For the TiAda algorithm, we set $\alpha=0.6$, $\beta=0.4$, $\eta_{x}=0.8$, $\eta_{y}=0.4$, $v^{x}_{0}=v^{y}_{0}=1$.
+For the AGP algorithm, we set $\alpha_{x}=0.14$, $\beta_{y}=1.1$.
+The algorithms terminate when $\|\nabla f(x_{k},y_{k})\|\leq 10^{-5}$ satisfied.
+Figure 1: Numerical Performance of four tested algorithms for solving the Synthetic Minimax problem
+Results.
+Figure 1 shows the stationarity gap of all four test algorithms, which demonstrates that the proposed PF-AGP-NSC algorithm exhibits significantly better convergence performance than that of the SGDA-B algorithm, TiAda algorithm and AGP algorithm.
 3.2 Dirac-GAN problem
-The Dirac-GAN problem (Mescheder et al., 2018) can be formulated as the following nonconvex–concave minimax problem :
+The Dirac-GAN problem can be formulated as the following nonconvex-concave minimax problem (Mescheder et al., 2018):
 $\min_{x}\max_{y}L(x,y)=-\log(1+\exp(-xy))+\log 2,$ (83)
@@
 Experimental setup.
-We set the initial point of the experiment at $(x,y)=(1,1)$.
-The termination condition is defined as the norm of the gradient $\|\nabla f(x_{i},y_{i})\|$ being less than or equal to $10^{-5}$, where $i$ represents the number of gradient calls.
-For the AGP algorithm, let $\alpha_{x}$ and $\beta_{y}$ be the stepsize of $x$ and $y$, respectively.
-We set $\alpha_{x}=\frac{0.8}{\sqrt{k}}$, $\beta_{y}=0.3$ and $c_{k}=\frac{0.5}{k^{1/4}}$.
-For the PF-AGP-NC algorithm, we do not need to preset specific parameters.
-In the experiments, we counted the number of gradient calls of the PF-AGP-NC algorithm.
-The results of alternating gradient projection in the $k$ th iteration are represented by nodes in the graph and the number of backtracking step computations is represented by the connecting lines.
-Results Figure 1 shows the convergence curves of the AGP and PF-AGP-NC algorithms.
-It can be seen that although the PF-AGP-NC algorithm involves a double-layer structure, requiring the determination of points in the outer layer through computations in the inner layer, even when considering the number of iterations in the inner layer, the overall convergence speed is still faster than that of the AGP algorithm.
-Figure 1: number of gradient calls $i$ vs $\log(||\nabla f(x_{i},y_{i})||)$
-4 Conclusions and discussion
+We compare the PF-AGP-NC algorithm with the SGDA-B algorithm and the AGP algorithm.
+The initial point of all algorithms is chosen as $(1,1)$.
+For the PF-AGP-NC algorithm, we set $l_{11}^{1}=0.01$ , $l_{12}^{1}=1$ and $l_{22}^{1}=0.01$.
+For the SGDA-B algorithm, we set $\mu=1$, $\gamma=0.5$, $N=1$, $p=0.5$ and $\tilde{L}$ = 1.
+For the AGP algorithm, we set $\alpha_{x}=\frac{0.8}{\sqrt{k}}$, $\beta_{y}=0.3$ and $c_{k}=\frac{0.5}{k^{1/4}}$.
+The algorithms terminate when $\|\nabla f(x_{k},y_{k})\|\leq 10^{-5}$ satisfied.
+Results.
+Figure 2 shows the stationarity gap and the sequences generated by all three test algorithms.
+We find that the SGDA-B algorithm fails to converge, both the AGP algorithm and the PF-AGP-NC algorithms successfully approach the unique stationary point.
+Moreover, the PF-AGP-NC algorithm exhibits significantly better convergence performance than that of the AGP algorithm.
+Figure 2: Numerical performance of three tested algorithms for solving the Dirac-GAN problem
+3.3 Robust learning over multiple domains
+In this subsection, we consider the following robust learning problem over multiple domains (Qian et al., 2019), formulated as a nonconvex-linear minimax problem:
+$\min_{x}\max_{y\in\Delta}y^{T}F(x),$ (84)
+where $F(x):=\left[f_{1}(x);\ldots;f_{M}(x)\right]\in\mathbb{R}^{M\times 1}$ with $f_{m}(x)=\frac{1}{\left|\mathcal{S}_{m}\right|}\sum_{i=1}^{\left|\mathcal{S}_{m}\right|}\ell\left(d_{i}^{m},l_{i}^{m},x\right)$ and $\ell$ can be any non-negative loss function, $M$ is the number of tasks and $x$ denotes the network parameters, $y\in\Delta$ describes the weights over different tasks and $\Delta$ is the simplex, i.e., $\Delta=\left\{(y_{1},\cdots,y_{M})^{T}\mid\sum_{m=1}^{M}y_{m}=1,y_{m}\geq 0\right\}$.
+We consider two image classification problems with MNIST (LeCun et al., 1998) and CIFAR10 datasets (Krizhevsky et al., 2009).
+Our goal is to train a neural network that works on these two completely unrelated problems simultaneously.
+Since cross-entropy loss is popular in multi-class classification problem, we use it as our loss function.
+We should point out that the quality of the algorithm for solving a robust learning problem over multiple domains is measured by the worst case accuracy over all domains (Qian et al., 2019).
+Experiment setup.
+We compare the proposed PF-AGP-NC algorithm with the SGDA-B algorithm and the AGP algorithm.
+For the PF-AGP-NC algorithm, we set $l_{11}^{1}=0.01$, $l_{12}^{1}=0.01$ and $l_{22}^{1}=0.01$.
+For the SGDA-B algorithm, we set $\mu=20$, $\gamma=0.9$, $N=1$, $p=0.5$ and $\tilde{L}$ = 20.
+For the AGP algorithm, we set $\frac{1}{\beta_{k}}=\frac{2}{2+\sqrt{k}}$, $\gamma_{k}=100$ and $c_{k}=\frac{1}{10+k^{\frac{1}{4}}}$.
+We set the batch size in all tasks to be $128$ and run $50$ epochs for all algorithms.
+Moreover, we sample our results every epoch and calculate the average accuracy on these two testing sets to evaluate the performance of different algorithms.
+Due to hardware (especially memory) limitations, we use a mini-batch randomized gradient instead of the exact gradient at each iteration in our numerical experiment, which is the same as in (Xu et al., 2023).
+Figure 3: Performance of two algorithms for solving robust multi-task learning problem.
+Results.
+Figure 3 and Figure 3 show the testing accuracy on the MNIST dataset and the CIFAR10 dataset respectively.
+PF-AGP-NC algorithm outperforms the SGDA-B algorithm and the AGP algorithm for both datasets.
+4 Conclusions
 In this paper, based on the framework of backtracking, we propose two completely parameter-free alternating gradient projection algorithms, i.e., the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm, for solving nonconvex-(strongly) concave minimax problems respectively, which does not require prior knowledge of parameters such as the Lipschtiz constant $L$ or the strongly concave constant $\mu$.

### v3

@@
 However, many existing algorithms require prior knowledge of the problem parameters in order to achieve optimal iteration complexity.
-In this paper, we propose two completely parameter-free alternating gradient projection algorithms, i.e., the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm, to solve the smooth nonconvex-strongly concave and nonconvex-concave minimax problems respectively using a backtracking strategy, which does not require prior knowledge of parameters such as the Lipschtiz constant $L$ or the strongly concave constant $\mu$.
-Moreover, we show that the total number of gradient calls of the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm to obtain an $\varepsilon$-stationary point is upper bounded by $\mathcal{O}\left(L\kappa^{3}\varepsilon^{-2}\right)$ and $\mathcal{O}\left(L^{4}\varepsilon^{-4}\right)$ respectively, where $\kappa$ is the condition number.
-As far as we know, the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm are the first completely parameter-free algorithms for solving nonconvex-strongly concave minimax problems and nonconvex-concave minimax problems respectively.
-Numerical results validate the efficiency of the proposed PF-AGP algorithm.
-Keywords: minimax optimization problem; backtracking strategies; parameter-free alternating gradient projection algorithm; iteration complexity; machine learning
+In this paper, three completely parameter-free single-loop algorithms, namely PF-AGP-NSC algorithm, PF-AGP-NC algorithm and PF-AGP-NL algorithm, are proposed to solve the smooth nonconvex-strongly concave, nonconvex-concave minimax problems and nonconvex-linear minimax problems respectively using line search without requiring any prior knowledge about parameters such as the Lipschtiz constant $L$ or the strongly concave modulus $\mu$.
+Furthermore, we prove that the total number of gradient calls required to obtain an $\varepsilon$-stationary point for the PF-AGP-NSC algorithm, the PF-AGP-NC algorithm, and the PF-AGP-NL algorithm are upper bounded by  $\mathcal{O}\left(L^{2}\kappa^{3}\varepsilon^{-2}\right)$, $\mathcal{O}\left(\log^{2}(L)L^{4}\varepsilon^{-4}\right)$, and $\mathcal{O}\left(L^{3}\varepsilon^{-3}\right)$, respectively, where $\kappa$ is the condition number.
+To the best of our knowledge, PF-AGP-NC and PF-AGP-NL are the first completely parameter-free algorithms for solving nonconvex-concave and nonconvex-linear minimax problems, respectively.
+PF-AGP-NSC is a completely parameter-free algorithm for solving nonconvex-strongly concave minimax problems, achieving the best known complexity with respect to $\varepsilon$.
+Numerical results demonstrate the efficiency of the three proposed algorithms.
+Keywords: minimax optimization; parameter-free alternating gradient projection algorithm; iteration complexity; machine learning
 1 Introduction
@@
 $\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}f(x,y),$ (P)
-where $\mathcal{X}\subseteq\mathbb{R}^{n}$ and $\mathcal{Y}\subseteq\mathbb{R}^{m}$ are nonempty closed and bounded convex sets, and $f:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ is a smooth function.
-Minimax problem is a critical class of optimization problems, known for its extensive research history and broad applications in areas such as machine learning, signal processing, and various other fields in recent years.
-Many practical problems can be formulated as in (P), including generative adversarial networks (GANs) (Arjovsky et al., 2017; Goodfellow et al., 2014; Sanjabi et al., 2018), reinforcement learning (Dai et al., 2018), power control and transceiver design in signal processing (Lu et al., 2020), distributed nonconvex optimization (Giannakis et al., 2017; Giordano et al., 2018; Mateos et al., 2010), robust optimization (Ben and Tal, 2009; Gao and Kleywegt, 2023), statistical learning (Giordano et al., 2018; Shafieezadeh Abadeh et al., 2015) and several other applications.
-There are many research results on convex-concave minimax problems, i.e., $f(x,y)$ is convex with respect to $x$ and concave with respect to $y$.
-Nemirovski (Nemirovski, 2004) introduced the mirror-prox method, and Nestrov (Nesterov, 2007) proposed a dual extrapolation algorithm, both of which achieve an optimal iteration complexity of $\mathcal{O}(\varepsilon^{-1})$ for solving convex-concave minimax problems.
-For more details, see (Vandenberghe and Boyd, 2004; Chen et al., 2014, 2017; Dang and Lan, 2014; Monteiro and Svaiter, 2010; Nedić and Ozdaglar, 2009; Yang et al., 2020) and references therein.
-Recently, various algorithms have been proposed for nonconvex-(strongly) concave minimax problems.
-For nonconvex-strongly concave minimax problems, several recent works (Jin et al., 2020; Lin et al., 2020a; Rafique et al., 2022; Lu et al., 2020) have proposed various algorithms that achieve a gradient complexity of $\tilde{\mathcal{O}}(\kappa_{y}^{2}\varepsilon^{-2})$ for stationary points of $\Phi(\cdot)=\max_{y\in\mathcal{Y}}f(\cdot,y)$ (when $\mathcal{X}=\mathbb{R}^{n}$), or for stationary points of $f$, where $\kappa_{y}$ is the condition number for $f(x,\cdot)$.
-(Lin et al., 2020b) proposed an accelerated algorithm that improves the gradient complexity bound to $\tilde{\mathcal{O}}(\sqrt{\kappa_{y}}\varepsilon^{-2})$.
-Furthermore, Zhang et al.
-(Zhang et al., 2021) proposed a generic acceleration framework which can improve the iteration complexity to $\mathcal{O}\left(\sqrt{\kappa_{y}}\varepsilon^{-2}\right)$.
+where $\mathcal{X}\subseteq\mathbb{R}^{n}$ and $\mathcal{Y}\subseteq\mathbb{R}^{m}$ are non-empty closed compact convex sets, and $f:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ is a smooth function.
+In recent years, the minimax problems have attracted the attention of scholars in the cross-disciplinary fields of optimization and machine learning due to its wide application.
+Many practical problems can be modeled as (P), such as generative adversarial networks (GAN) (Arjovsky et al., 2017; Goodfellow et al., 2014; Sanjabi et al., 2018), reinforcement learning (Dai et al., 2018), power control and transceiver design in signal processing Lu et al.
+(2020), distributed non-convex optimization (Giannakis et al., 2017; Giordano et al., 2018; Mateos et al., 2010), robust optimization (Ben-Tal et al., 2009; Gao and Kleywegt, 2023), statistical learning (Giordano et al., 2018; Shafieezadeh-Abadeh et al., 2015), etc.
+When $f(x,y)$ is (non)convex with respect to $x$ and (non)concave with respect to $y$, we call (P) the (non)convex-(non)concave minimax problem respectively.
+There are many research results on convex-concave minimax problems.
+The mirror neighbor method was proposed in (Nemirovski, 2004), and the dual extrapolation algorithm was proposed in (Nesterov, 2007), both of which achieve the optimal iteration complexity $\mathcal{O}(\varepsilon^{-1})$ for solving convex-concave minimax problems.
+For more related research results, please refer to (Vandenberghe and Boyd, 2004; Chen et al., 2014, 2017; Dang and Lan, 2014; Monteiro and Svaiter, 2010; Nedić and Ozdaglar, 2009; Yang et al., 2020) and their references.
+Recently, various algorithms have been proposed to solve the nonconvex-(strongly) concave minimax problems.
+For the nonconvex-strongly concave minimax problems, some recent studies (Jin et al., 2020; Lin et al., 2020a; Rafique et al., 2022; Lu et al., 2020) proposed various algorithms that achieve gradient complexity of $\tilde{\mathcal{O}}(\kappa_{y}^{2}\varepsilon^{-2})$ for stationary points of $\Phi(\cdot)=\max_{y\in\mathcal{Y}}f(\cdot,y)$ when $\mathcal{X}=\mathbb{R}^{n}$ or for stationary points of $f$ where $\kappa_{y}$ is the condition number of $f(x,\cdot)$.
+An acceleration algorithm was proposed in (Lin et al., 2020b), which improves the gradient complexity to $\tilde{\mathcal{O}}(\sqrt{\kappa_{y}}\varepsilon^{-2})$.
+In addition, Zhang et al.
+(Zhang et al., 2021) proposed a general acceleration framework, which improves the iteration complexity to $\mathcal{O}\left(\sqrt{\kappa_{y}}\varepsilon^{-2}\right)$.
 For general nonconvex-concave minimax problems, there are two types of algorithms, i.e., multi-loop algorithms and single-loop algorithms.
-Various multi-loop algorithms have been proposed in (Kong and Monteiro, 2021; Lin et al., 2020b; Nouiehed et al., 2019; Ostrovskii et al., 2021; Rafique et al., 2022; Thekumparampil et al., 2019).
+Various multi-loop algorithms have been proposed in (Kong and Monteiro, 2021; Lin et al., 2020b; Nouiehed et al., 2019b; Ostrovskii et al., 2021; Rafique et al., 2022; Thekumparampil et al., 2019).
 The best known iteration complexity of multi-loop algorithms for solving nonconvex-concave minimax problems is $\tilde{\mathcal{O}}(\varepsilon^{-2.5})$, which was achieved by (Lin et al., 2020b).
@@
 (Pan et al., 2021) proposed a new alternating gradient projection algorithm with the iteration complexity of $\mathcal{O}(\varepsilon^{-3})$.
-Various single-loop algorithms have also been proposed, e.g., the gradient descent-ascent (GDA) algorithm (Lin et al., 2020a), the hybrid block successive approximation (HiBSA) algorithm (Lu et al., 2020), the unified single-loop alternating gradient projection (AGP) algorithm (Xu et al., 2023), and the smoothed GDA algorithm (Zhang et al., 2020).
-Both the AGP algorithm and the smoothed GDA algorithm achieve the best known iteration complexity, i.e., $\mathcal{O}(\varepsilon^{-4})$, among single-loop algorithms for solving nonconvex-concave minimax problems.
-It should be noted that for most of the existing results mentioned above, achieving optimal complexity of the algorithm requires the assumption of knowing precise information about some parameters of the problem, such as the Lipschitz constant, or the strongly concave constant $\mu$, etc.
-Accurately estimating these parameters is often challenging, and conservative estimates can significantly impact algorithm performance (Lan et al., 2023).
-Therefore, designing parameter-free algorithms with complexity guarantees without relying on inputs of these parameters has attracted considerable attention recently.
-For nonconvex-strongly concave minimax problems, there are two main parameter-free algorithms: one based on AdaGrad stepsizes and the other based on backtracking.
-In the framework of AdaGrad stepsizes, Yang et al.
-(Yang et al., 2022) proposed NeAda, which can achieve convergence rates of $\tilde{\mathcal{O}}(\varepsilon^{-2})$ and $\tilde{\mathcal{O}}(\varepsilon^{-4})$ in deterministic and stochastic settings, respectively.
-Li et al.
-(Li et al., 2022) proposed TiAda, achieving convergence rates of ${\mathcal{O}}(\varepsilon^{-2})$ and ${\mathcal{O}}(\varepsilon^{-4})$ in deterministic and stochastic settings, respectively.
-In the framework of backtracking, Xu et al.
-(Xu et al., 2024) proposed a stochastic gradient descent ascent (GDA) method with backtracking (SGDA-B), achieving an upper bound on the total number of gradient calls of $\mathcal{O}(L\kappa^{2}\log(\kappa)\varepsilon^{-2})$ with $\mathcal{O}(L\kappa^{2}\varepsilon^{-2})$ gradient calls required per backtracking step.
-Meanwhile, Zhang et al.
-(Zhang et al., 2024) proposed a proximal alternating gradient descent ascent method (AGDA+), achieving an upper bound on the total number of gradient calls of $\mathcal{O}(L\kappa^{4}\varepsilon^{-2})$ with $\mathcal{O}(1)$ gradient calls required per backtracking step.
-Both methods do not require knowledge of the specific value of Lipschtiz constant $L$, but still require information about the strongly concave constant $\mu$ in the function.
-In the framework of backtracking, Xu et al.
-(Xu et al., 2024) proposed a stochastic GDA method with backtracking (SGDA-B) for nonconvex-concave minimax problems, achieving an upper bound on the total number of gradient calls of $\tilde{\mathcal{O}}(L^{3}\varepsilon^{-4})$ with $\mathcal{O}(L^{3}\varepsilon^{-4})$ gradient calls required per backtracking step.
+Various single-loop algorithms have also been proposed for solving nonconvex-concave minimax problems, e.g., the gradient descent-ascent (GDA) algorithm (Lin et al., 2020a), the hybrid block successive approximation (HiBSA) algorithm (Lu et al., 2020), the unified single-loop alternating gradient projection (AGP) algorithm (Xu et al., 2023), and the smoothed GDA algorithm (Zhang et al., 2020).
+Both the AGP algorithm and the smoothed GDA algorithm achieve the best known iteration complexity, i.e., $\mathcal{O}(\varepsilon^{-4})$, among single-loop algorithms.
+It should be noted that for most of the above existing results, achieving the optimal complexity of the algorithm requires assuming accurate information about certain parameters of the known problem, such as the Lipschitz constant or the strong concavity modulus $\mu$, etc.
+Accurately estimating these parameters is usually challenging, and conservative estimates will significantly affect the algorithm performance (Lan et al., 2023).
+Therefore, designing parameter-free algorithms that do not rely on these parameter inputs and have complexity guarantees has attracted widespread attention in recent years.
+For nonconvex-strongly concave minimax problems, there are two main types of parameter-free algorithms: one based on AdaGrad steps and the other based on line search.
+In the framework of AdaGrad steps, the NeAda algorithm was proposed in (Yang et al., 2022), which obtains an $\varepsilon$-stationary point with an iteration complexity of $\tilde{\mathcal{O}}(\varepsilon^{-2})$ and $\tilde{\mathcal{O}}(\varepsilon^{-4})$ in deterministic and stochastic settings, respectively.
+The TiAda algorithm was proposed in (Li et al., 2022), which obtains an $\epsilon$-stationary point with an iteration complexity of ${\mathcal{O}}(\varepsilon^{-2})$ and ${\mathcal{O}}(\varepsilon^{-4})$ in deterministic and stochastic settings, respectively.
+Based on the line search framework, a stochastic gradient descent with backtracking (SGDA-B) algorithm was proposed in (Xu et al., 2024), which has an upper bound of $\mathcal{O}(L\kappa^{2}\log(\kappa)\varepsilon^{-2})$ on the total number of gradient calls to obtain an $\varepsilon$-stationary point.
+A proximal alternating gradient descent (AGDA+) method was proposed in (Zhang et al., 2025) which achieves an upper bound of the total number of gradient calls of $\mathcal{O}(\bar{L}^{2}\bar{\kappa}^{4}\log(\mathcal{R})\varepsilon^{-2})$, where $\bar{\kappa}=\bar{L}/\bar{\mu}$, $\bar{L}=\mathcal{R}\tilde{l}$ and $\bar{\mu}=\tilde{\mu}/\mathcal{R}$, $\mathcal{R}=\max\{\tilde{\mu}/\mu,L/\tilde{l},1\}$, $\tilde{\mu}\geq\mu$ and $\tilde{l}\leq L$ denote the initial estimate of the concavity modulus $\mu$ and the local Lipschitz constant $L$.
+For the nonconvex-concave minimax problems, a stochastic GDA method with backtracking (SGDA-B) is proposed in (Xu et al., 2024), which obtains an upper bound of $\tilde{\mathcal{O}}(L^{3}\varepsilon^{-4})$ on the total number of gradient calls to obtain an $\varepsilon$-stationary point.
+For nonconvex-linear minimax problems (a special class of nonconvex-concave minimax problems), before the results of this paper, there were no relevant results on parameter-free optimization algorithms with an iteration complexity of $\mathcal{O}(\varepsilon^{-3})$.
 1.1 Contributions
-In this paper, based on the framework of backtracking, we propose two completely parameter-free alternating gradient projection algorithms, i.e., the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm, for solving nonconvex-(strongly) concave minimax problems respectively, which does not require prior knowledge of parameters such as the Lipschtiz constant $L$ or the strongly concave constant $\mu$.
-Moreover, we show that the total number of gradient calls of the PF-AGP-NSC algorithm to obtain an $\varepsilon$-stationary point for nonconvex-strongly concave minimax problems is upper bounded by $\mathcal{O}\left(L\kappa^{3}\varepsilon^{-2}\right)$ where $\kappa$ is the condition number, while the total number of gradient calls of the PF-AGP-NC algorithm to obtain an $\varepsilon$-stationary point for nonconvex-concave minimax problems is upper bounded by $\mathcal{O}\left(L^{4}\varepsilon^{-4}\right)$.
-As far as we know, the PF-AGP-NSC algorithm is the first completely parameter-free algorithm for solving nonconvex-strongly concave minimax problems, and the PF-AGP-NC algorithm is the first completely parameter-free algorithm which achieves the best iteration complexity in single loop algorithms for solving nonconvex-concave minimax problems.
+Based on the line search framework, this paper proposes three completely parameter-free alternating gradient projection algorithms, namely PF-AGP-NSC algorithm, PF-AGP-NC algorithm and PF-AGP-NL algorithm, which are used to solve nonconvex-strongly concave, nonconvex-concave and nonconvex-linear minimax problems respectively.
+These algorithms do not require prior knowledge of any problem-related parameters, including the Lipschtiz constant $L$ and the strongly concave modulus $\mu$.
+Furthermore, we prove that the total number of gradient calls required to obtain an $\varepsilon$-stationary point for the PF-AGP-NSC algorithm, the PF-AGP-NC algorithm, and the PF-AGP-NL algorithm are upper bounded by  $\mathcal{O}\left(L^{2}\kappa^{3}\varepsilon^{-2}\right)$, $\mathcal{O}\left(\log^{2}(L)L^{4}\varepsilon^{-4}\right)$, and $\mathcal{O}\left(L^{3}\varepsilon^{-3}\right)$, respectively.
+Table 1 provides a detailed comparison of the three proposed parameter-free optimization algorithms with existing parameter-free optimization algorithms for solving nonconvex-strongly concave (NC-SC), nonconvex-concave (NC-C), and nonconvex-linear (NC-L) minimax problems.
+Column “Completely PF” indicates whether the algorithm is completely parameter-free (yes/no), and column “Prior Information” specifies any prior knowledge or parameters required by the algorithm, where $\underline{S}=\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}f(x,y)$ and $\mathcal{D}_{y}=\max\{\|y_{1}-y_{2}\|\mid y_{1},y_{2}\in\mathcal{Y}\}$ and ”None” indicates that no prior information is required.
+In “Gradient Complexity” column, $\bar{\kappa}\geq\kappa$, $\mathcal{R}$ is a constant which depends on $L$ and $\mu$ and $\mathcal{R}\leq\kappa$.
+To the best of our knowledge, the PF-AGP-NL algorithm is the first completely parameter-free single-loop algorithm for solving nonconvex-linear minimax problems, achieving the best known gradient complexity within the single-loop algorithms.
+The PF-AGP-NSC and the PF-AGP-NC algorithm is the first completely parameter-free single-loop algorithm for solving nonconvex-strongly concave and nonconvex-concave minimax problems, achieving the best known gradient complexity with respect to $\varepsilon$ with the existing single-loop algorithms respectively.
+It should be pointed out that the complexity of the proposed PF-AGP-NSC and PF-AGP-NC algorithms is slightly worse than that of the SGDA-B algorithm in terms of the Lipschitz constant L and the condition number $\kappa$, but the SGDA-B algorithm requires some prior information about the remaining parameters and is not a completely parameter-free algorithm.
+If some other prior information is also allowed, the complexity of the proposed algorithm can also be improved.
+To better illustrate this point, if two additional parameters $\mu$ and $\underline{S}$ are known in advance, in Appendix A we propose a restarted variant of the PF-AGP-NSC algorithm, named rPF-AGP-SC, and prove that the gradient complexity of the algorithm is improved to $\mathcal{O}\left(\log(L)L\kappa^{2}\varepsilon^{-2}\right)$, which is better than that of the SGDA-B.
+Problems Algorithms Gradient Complexity Completely PF Prior Information
+NC-SC NeAda $\mathcal{O}(\log(\kappa)\kappa^{4.5}L^{4}\varepsilon^{-2})$ Yes None
+TiAda $\mathcal{O}(\kappa^{10}\varepsilon^{-2})$ Yes None
+SGDA-B $\mathcal{O}(L\kappa^{2}\log(\kappa)\varepsilon^{-2})$ No $\mu$, $\underline{S}$
+AGDA+ $\mathcal{O}(\bar{L}^{2}\bar{\kappa}^{4}\log(\mathcal{R})\varepsilon^{-2})$ Yes None
+PF-AGP-NSC $\mathcal{O}(L^{2}\kappa^{3}\varepsilon^{-2})$ Yes None
+NC-C SGDA-B $\mathcal{O}(L^{3}\log({L}{\varepsilon^{-1}})\varepsilon^{-4})$ No $\mathcal{D}_{y}$
+PF-AGP-NC $\mathcal{O}(\log^{2}(L)L^{4}\varepsilon^{-4})$ Yes None
+NC-L PF-AGP-NL $\mathcal{O}(L^{3}\varepsilon^{-3})$ Yes None
+Table 1: Comparison of Gradient Complexity of Existing Algorithms for Nonconvex-Concave Minimax Problems.
 1.2 Organization
-In Sect.
-2, we propose two parameter-free alternating gradient projection (PF-AGP) algorithms for nonconvex-strongly concave and nonconvex-concave minimax problems, and we then analyze their corresponding gradient complexity under two different settings.
-We report some numerical results in Sect.
-3 and make some conclusions in the last section.
+In Section 2, we propose two parameter-free single-loop algorithms, PF-AGP-NSC and PF-AGP-NC, for the nonconvex-strongly concave and nonconvex-concave minimax problems, and analyze their corresponding gradient complexities in two different settings.
+In Section 3, we propose a parameter-free alternating gradient projection (PF-AGP-NL) algorithm for the nonconvex-linear minimax problem and analyze its corresponding gradient complexity.
+We report some numerical results in Section 4 and draw some conclusions in the last section.
 […]
 $f(y)\geq f(x)+\langle\nabla f(x),y-x\rangle+\frac{\theta}{2}\|y-x\|^{2}.$ (1)
-If $-f$ satisfies (1), $f(\cdot)$ is called $\theta$-strongly concave. Denote the sign function as $\mbox{sgn}(x)$, i.e.,
+$f(\cdot)$ is called $\theta$-strongly concave if $-f$ satisfies (1). The symbolic function is represented as $\mbox{sgn}(x)$, that is:
 $\mbox{sgn}(x)=\begin{cases}1&\text{if }x>0,\\ -1&\text{if }x\leq 0.\end{cases}$ (2)
-2 Two completely parameter-free alternating gradient projection algorithms
-In this section, we propose two completely parameter-free alternating gradient projection (PF-AGP) algorithms for solving the nonconvex-strongly concave and the nonconvex-concave minimax problems (P) respectively.
-Both algorithms do not require prior knowledge of Lipschitz constants, and the algorithm under nonconvex-strong concave settings also does not require prior knowledge of strong concave constants.
+2  Completely parameter-free single-loop algorithms for nonconvex- (strongly) concave setting
+In this section, we propose two completely parameter-free single-loop algorithms for solving the nonconvex-strongly concave and the nonconvex-concave minimax problems (P) respectively.
+Both algorithms do not require prior knowledge of Lipschitz constants, and the algorithm under nonconvex-strong concave settings also does not require prior knowledge of strong concavity coefficient.
 The two proposed algorithms are based on the AGP algorithm (Xu et al., 2023), which uses the gradient of a regularized version of the original function, i.e.,
-$f_{k}(x,y)=f(x,y)-\frac{c_{k}}{2}\|y_{k}\|^{2},$
-where $c_{k}\geq 0$ is a regularization parameters. At the $k$-th iteration, AGP algorithm consists of the following two gradient projection steps for updating both $x$ and $y$:
+$f_{k}(x,y)=f(x,y)-\frac{c_{k}}{2}\|y\|^{2},$
+where $c_{k}\geq 0$ is a regularization parameter. At the $k$-th iteration, AGP algorithm consists of the following two gradient projection steps for updating both $x$ and $y$:
 $x_{k+1}$ $=\mathcal{P}_{\mathcal{X}}\left(x_{k}-\tfrac{1}{\beta_{k}}\nabla_{x}f(x_{k},y_{k})\right),$
-$y_{k+1}$ $=\mathcal{P}_{\mathcal{Y}}\left(y_{k}+\tfrac{1}{\gamma_{k}}\nabla_{y}{f}(x_{k},y_{k})-\frac{1}{\gamma_{k}}c_{k}y_{k}\right),$
+$y_{k+1}$ $=\mathcal{P}_{\mathcal{Y}}\left(y_{k}+\tfrac{1}{\gamma_{k}}\nabla_{y}{f}(x_{k+1},y_{k})-\frac{1}{\gamma_{k}}c_{k}y_{k}\right),$
 where $\mathcal{P}_{\mathcal{X}}$ and $\mathcal{P}_{\mathcal{Y}}$ is the projection operator onto $\mathcal{X}$ and $\mathcal{Y}$, respectively, and $\beta_{k}>0$, $\gamma_{k}>0$ are stepsize parameters.
-Although the AGP algorithm achieves the optimal iteration complexity among single loop algorithms for nonconvex-(strongly) concave minimax problems, it requires the knowledge of the Lipschtiz constant $L$ and the strongly concave coefficient $\mu$, or at least knowledge of the Lipschtiz constant $L$ to calculate the stepsizes $\beta_{k}$ and $\gamma_{k}$ under the nonconvex-strongly concave setting and the nonconvex-concave setting, respectively, which limits its applicability.
+Although the AGP algorithm achieves the optimal iteration complexity dependence on $\varepsilon$ among single loop algorithms for nonconvex-(strongly) concave minimax problems, it requires the knowledge of the Lipschtiz constant $L$ and the strongly concave coefficient $\mu$, or at least knowledge of the Lipschtiz constant $L$ to calculate the stepsizes $\beta_{k}$ and $\gamma_{k}$ under the nonconvex-strongly concave setting and the nonconvex-concave setting, respectively, which limits its applicability.
 To eliminate the need for the prior knowledge of the Lipschtiz constant $L$ or the strongly concave coefficient $\mu$, we propose two parameter-free AGP algorithms by employing backtracking strategies for solving the nonconvex-strongly concave and the nonconvex-concave minimax problems, respectively.
@@
 We first propose a completely parameter-free AGP algorithm for solving nonconvex-strongly concave minimax problems.
-More specifically, in order to estimate the Lipschitz constants for $\nabla_{x}f(\cdot,y)$, $\nabla_{y}f(\cdot,y)$, $\nabla_{y}f(x,\cdot)$ denoted as $L_{11}$, $L_{12}$ and $L_{22}$, respectively, and the strongly concave constant $\mu$, at each iteration of the proposed algorithm, we aim to find a tuple $(l_{11}^{k,i},l_{12}^{k,i},l_{22}^{k,i},\mu_{k,i})$ by backtracking such that the following conditions for $(x_{k,i},y_{k,i})$ are satisfied:
-$C_{1}^{k,i}$ $=f(x_{k,i},y_{k})-f(x_{k},y_{k})-\langle\nabla_{x}f(x_{k},y_{k}),x_{k,i}-x_{k}\rangle-\frac{l^{k,i}_{11}}{2}\|x_{k,i}-x_{k}\|^{2}\leq 0,$ (C1)
+More specifically, in order to estimate the Lipschitz constants for $\nabla_{x}f(\cdot,y)$, $\nabla_{y}f(\cdot,y)$, $\nabla_{y}f(x,\cdot)$ denoted as $L_{11}$, $L_{12}$ and $L_{22}$, respectively, and the strongly concavity modulus $\mu$, at each iteration of the proposed algorithm, we aim to find a tuple $(l_{11}^{k,i},l_{12}^{k,i},l_{22}^{k,i},\mu_{k,i})$ by backtracking such that the following conditions for $(x_{k,i},y_{k,i})$ are satisfied:
+$C_{1}^{k,i}=$ $f(x_{k,i},y_{k})-f(x_{k},y_{k})-\langle\nabla_{x}f(x_{k},y_{k}),x_{k,i}-x_{k}\rangle-\frac{l^{k,i}_{11}}{2}\|x_{k,i}-x_{k}\|^{2}\leq 0,$
 $C_{2}^{k,i}=$ $\|\nabla_{y}f(x_{k,i},y_{k})-\nabla_{y}f(x_{k},y_{k})\|-l_{12}^{k,i}\|x_{k,i}-x_{k}\|\leq 0,$ (C2)
-$C_{3}^{k,i}$ $=\|\nabla_{y}f(x_{k,i},y_{k,i})-\nabla_{y}f(x_{k,i},y_{k})\|-l_{22}^{k,i}\|y_{k,i}-y_{k}\|\leq 0,$ (C3)
+$C_{3}^{k,i}=$ $l_{22}^{k,i}\langle\nabla_{y}f(x_{k,i},y_{k,i})-\nabla_{y}f(x_{k,i},y_{k,i}),y_{k,i}-y_{k}\rangle$
+$+\|\nabla_{y}f(x_{k,i},y_{k,i})-\nabla_{y}f(x_{k,i},y_{k})\|^{2}\leq 0,$ (C3)
 $C_{4}^{k,i}=$ $\langle\nabla_{y}f(x_{k,i},y_{k,i})-\nabla_{y}f(x_{k,i},y_{k}),y_{k,i}-y_{k}\rangle+\mu_{k,i}\|y_{k,i}-y_{k}\|^{2}\leq 0.$ (C4)
@@
 Algorithm 1 A parameter-free alternating gradient projection (PF-AGP-NSC) algorithm for nonconvex-strongly concave minimax problems
-  Step 1: Input $x_{1},y_{1}$, $\beta_{1}$, $\gamma_{1}$, $l_{11}^{1}$, $l_{12}^{1}$, $l_{22}^{1}$, $\mu_{1}$; Set $k=1$.
-  Step 2: Updata $x_{k}$ and $y_{k}$:
-   (a): Set $i=1$, $l_{11}^{k,i}=l_{11}^{k}$, $l_{12}^{k,i}=l_{12}^{k}$, $l_{22}^{k,i}=l_{22}^{k}$, $\mu_{k,i}=\mu_{k}$, $\beta_{k,i}=\beta_{k}$, $\gamma_{k,i}=\gamma_{k}$.
+  Step 1: Input $x_{1},y_{1}$, $\beta_{0}$, $\gamma_{0}$, $l_{11}^{0}$, $l_{12}^{0}$, $l_{22}^{0}$, $\mu_{0}$; Set $k=1$.
+  Step 2: Update $x_{k}$ and $y_{k}$:
+   (a): Set $i=0$, $l_{11}^{k,i}={l_{11}^{k-1}}{}$, $l_{12}^{k,i}={l_{12}^{k-1}}$, $l_{22}^{k,i}=l_{22}^{k-1}$, $\mu_{k,i}=\mu_{k-1}$, $\beta_{k,i}=\beta_{k-1}$, $\gamma_{k,i}=\gamma_{k-1}$.
    (b): Update $x_{{k,i}}$ and $y_{{k,i}}$:
@@
 $y_{{k,i}}$ $=\mathcal{P}_{\mathcal{Y}}\left(y_{k}+\frac{1}{\gamma_{k,i}}\nabla_{y}f(x_{{k,i}},y_{k})\right).$ (4)
-   (c): Compute $C_{1}^{k,i}$, $C_{2}^{k,i}$, $C_{3}^{k,i}$, $C_{4}^{k,i}$ as in (C1), (C2), (C3), (C4), respectively;
+   (c): Compute $C_{1}^{k,i}$, $C_{2}^{k,i}$, $C_{3}^{k,i}$, $C_{4}^{k,i}$ as in (2.1), (C2), (C3), (C4), respectively;
    (d): Update $l^{k,i}_{11}$, $l^{k,i}_{12}$, $l^{k,i}_{22}$, $\mu_{k,i}$:
@@
      Otherwise, $i=i+1$,
-$\beta_{k,i}={l_{11}^{k,i}}+\frac{l_{12}^{k,i}\mu_{k-1,i}}{8(l_{22}^{k-1,i})^{2}}+\frac{32(l_{12}^{k,i})^{2}(l_{22}^{k-1,i})^{2}}{\mu_{k,i}\mu_{k-1,i}^{2}},\quad\gamma_{k,i}=\frac{8(l_{22}^{k,i})^{2}}{\mu_{k,i}},$
+$\beta_{k,i}={l_{11}^{k,i}}+l_{12}^{k,i}+\frac{32(l_{12}^{k,i})^{2}(l^{k-1}_{12}+l^{k-1}_{22})}{\mu_{k,i}\mu_{k-1}},\quad\gamma_{k,i}=l_{12}^{k,i}+l_{22}^{k,i},$
      go to Step 2(b).
@@
 It is worth noting that the parameter-free NeAda (Yang et al., 2022) and TiAda (Li et al., 2022) algorithms with adaptive step size are only used to solve the minimax problems when $x$ is unconstrained, while the proposed PF-AGP-NSC algorithm can solve the constrained case.
-Compared with the SGDA-B (Xu et al., 2023) algorithm and AGDA+ (Zhang et al., 2024) algorithm using the backtracking framework under the nonconvex-strongly concave setting, AGDA+ and SGDA-B still require knowledge of the strongly concave constant $\mu$, while the proposed PF-AGP-NSC algorithm is a completely parameter-free algorithm that does not require any prior knowledge of the parameters.
+Compared with the SGDA-B (Xu et al., 2024) algorithm and AGDA+ (Zhang et al., 2025) algorithm using the backtracking framework under the nonconvex-strongly concave setting, SGDA-B still requires knowledge of the strongly concavity modulus $\mu$, while the proposed PF-AGP-NSC algorithm is a completely parameter-free algorithm that does not require any prior knowledge of the parameters.
+Moreover, the complexity of the proposed PF-AGP-NSC is better than AGDA+.
 2.1.1 Complexity analysis
@@
 At each iteration of Algorithm 1, the stationarity gap for problem (P) with respect to $f_{k}(x,y)$ is defined as:
-$\nabla\tilde{G}(x_{k},y_{k})=\begin{bmatrix}\beta_{k}\left(x_{k}-P_{x}\left(x_{k}-\frac{1}{\beta_{k}}\nabla_{x}f_{k}(x_{k},y_{k})\right)\right)\\ \gamma_{k}\left(y_{k}-P_{y}\left(y_{k}+\frac{1}{\gamma_{k}}\nabla_{y}f_{k}(x_{k},y_{k})\right)\right)\end{bmatrix}.$
+$\nabla\tilde{G}(x_{k},y_{k})=\begin{bmatrix}\beta_{k}\left(x_{k}-P_{\mathcal{X}}\left(x_{k}-\frac{1}{\beta_{k}}\nabla_{x}f_{k}(x_{k},y_{k})\right)\right)\\ \gamma_{k}\left(y_{k}-P_{\mathcal{Y}}\left(y_{k}+\frac{1}{\gamma_{k}}\nabla_{y}f_{k}(x_{k},y_{k})\right)\right)\end{bmatrix}.$
 Under Assumption 1, we first establish the following lemma, which provides bounds on the changes in the function value when $x_{k}$ is updated at each iteration of Algorithm 1.
@@
 $\langle\nabla_{x}f\left(x_{k},y_{k}\right)+\beta_{k}\left(x_{k+1}-x_{k}\right),x_{k}-x_{k+1}\rangle\geq 0.$ (9)
-The backtracking strategy (C1) implies that
+The backtracking strategy (2.1) implies that
 $f\left(x_{k+1},y_{k}\right)-f\left(x_{k},y_{k}\right)\leq\langle\nabla_{x}f\left(x_{k},y_{k}\right),x_{k+1}-x_{k}\rangle+\tfrac{l^{k}_{11}}{2}\|x_{k+1}-x_{k}\|^{2}.$ (10)
@@
 which completes the proof.
-Next, we provide an estimate of upper bound of the difference between $f(x_{k+1},y_{k+1})$ and $f(x_{k},y_{k})$. First, we need to make the following assumption on the parameter $\gamma_{k}$.
+Next, we provide an upper bound of the difference between $f(x_{k+1},y_{k+1})$ and $f(x_{k},y_{k})$. First, we need to make the following assumption on the parameter $\gamma_{k}$.
 Assumption 2
@@
 Lemma 4
-Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 1, then we have
+Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 1,  if $\gamma_{k}\geq l_{22}^{k}$, then we have
 $f(x_{k+1},y_{k+1})-f(x_{k},y_{k})$
-$\quad\leq-\left(\beta_{k}-\frac{l_{11}^{k}}{2}-\frac{l_{12}^{k}}{2\gamma_{k-1}}\right)\|x_{k+1}-x_{k}\|^{2}+{\gamma_{k-1}}\|y_{k+1}-y_{k}\|^{2}$
-$\quad\quad-\left(\mu_{k-1}-\frac{\gamma_{k-1}}{2}-\frac{(l_{22}^{k-1})^{2}}{2\gamma_{k-1}}\right)\|y_{k}-y_{k-1}\|^{2}.$ (11)
+$\leq$ $-\left(\beta_{k}-\frac{l_{11}^{k}}{2}-\frac{l_{12}^{k}}{2}\right)\|x_{k+1}-x_{k}\|^{2}+\left(\frac{l^{k}_{12}}{2}+\frac{\gamma_{k-1}}{2}\right)\|y_{k+1}-y_{k}\|^{2}$
+$-\left(\frac{\mu_{k-1}}{2}-\frac{\gamma_{k-1}}{2}\right)\|y_{k}-y_{k-1}\|^{2}.$ (11)
 Proof  The optimality condition for $y_{k}$ in (4) implies that
@@
 $\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle$
-$=$ $\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle+\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),v_{k+1}\rangle$
+$=$ $\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle$
+$+\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),v_{k+1}\rangle$
 $+\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle.$ (14)
-Next, we estimate the three terms on the right-hand side of (14) respectively. By the backtracking strategy (C2), (C3) and the Cauchy-Schwarz inequality, we have
-$\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle\leq\frac{l_{12}^{k}}{2\gamma_{k}}\|x_{k+1}-x_{k}\|^{2}+\frac{\gamma_{k}}{2}\|y_{k+1}-y_{k}\|^{2}$ (15)
+Next, we estimate the three terms on the right-hand side of (14) respectively. By the backtracking strategy (C2) and the Cauchy-Schwarz inequality, we have
+$\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle\leq\frac{l_{12}^{k}}{2}\|x_{k+1}-x_{k}\|^{2}+\frac{l^{k}_{12}}{2}\|y_{k+1}-y_{k}\|^{2}$ (15)
 and
-$\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),v_{k+1}\rangle\leq\frac{(l_{22}^{k-1})^{2}}{2\gamma_{k-1}}\|y_{k}-y_{k-1}\|^{2}+\frac{\gamma_{k-1}}{2}\|v_{k+1}\|^{2}.$ (16)
-By the backtracking strategy (C4) of $\mu_{k-1}$, we obtain
-$\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle\leq-\mu_{k-1}\|y_{k}-y_{k-1}\|^{2}.$ (17)
+$\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),v_{k+1}\rangle\leq\frac{1}{2\gamma_{k-1}}\|f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1})\|+\frac{\gamma_{k-1}}{2}\|v_{k+1}\|.$ (16)
+By the backtracking strategy (C3) and (C4), we obtain
+$\begin{split}&\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle\\ \leq&-\frac{1}{2l_{22}^{k-1}}\|f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1})\|-\frac{\mu_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}.\end{split}$ (17)
 Moreover, it can be easily checked that
@@
 Plugging (14)-(18) into (13) and rearranging the terms, we conclude that
-$f(x_{k+1},y_{k+1})-f(x_{k+1},y_{k})$ $\leq\frac{l_{12}^{k}}{2\gamma_{k}}\|x_{k+1}-x_{k}\|^{2}-\left(\mu_{k-1}-\frac{\gamma_{k-1}}{2}-\frac{(l_{22}^{k-1})^{2}}{2\gamma_{k-1}}\right)\|y_{k}-y_{k-1}\|^{2}$
-$\quad+\left(\frac{\gamma_{k}}{2}+\frac{\gamma_{k-1}}{2}\right)\|y_{k+1}-y_{k}\|^{2}.$ (19)
+$f(x_{k+1},y_{k+1})-f(x_{k+1},y_{k})$
+$\leq$ $\frac{l^{k}_{12}}{2}\|x_{k+1}-x_{k}\|^{2}+\left(\frac{l^{k}_{12}}{2}+\frac{\gamma_{k-1}}{2}\right)\|y_{k}-y_{k-1}\|^{2}-\left(\frac{\mu_{k-1}}{2}-\frac{\gamma_{k-1}}{2}\right)\|y_{k+1}-y_{k}\|^{2}.$ (19)
 By Lemma 8, we have
@@
 Lemma 5
-Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 1, and denoting
-$f_{k+1}$ $=f(x_{k+1},y_{k+1}),\quad S_{k+1}=\frac{2\gamma^{2}_{k}}{\mu_{k}}\|y_{k+1}-y_{k}\|^{2},$
-$F_{k+1}$ $=f_{k+1}+S_{k+1}-\left(\mu_{k}+\frac{7\gamma_{k}}{2}-\frac{(l^{k}_{22})^{2}}{2\gamma_{k}}-\frac{2(l^{k}_{22})^{2}}{\mu_{k}}\right)\|y_{k+1}-y_{k}\|^{2}-\frac{2\gamma_{k}^{2}}{\mu_{k}}\mathcal{D}^{2}_{y},$ (21)
-then we have
-$F_{k+1}-F_{k}\leq$ $-\left(\frac{2\mu_{k}\gamma_{k}-(l_{22}^{k})^{2}}{2\gamma_{k}}+\frac{\mu_{k}\gamma_{k}-4(l_{22}^{k})^{2}}{2\mu_{k}}\right)\|y_{k+1}-y_{k}\|^{2}$
-$\quad-\left(\beta_{k}-\frac{l_{11}^{k}}{2}-\frac{l_{12}^{k}}{2\gamma_{k-1}}-\frac{2(l_{12}^{k})^{2}\gamma_{k-1}}{\mu_{k-1}\mu_{k}}\right)\|x_{k+1}-x_{k}\|^{2},$ (22)
+Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 1, and denote
+$f_{k+1}$ $=f(x_{k+1},y_{k+1}),\quad S_{k+1}=\frac{4\gamma^{2}_{k}}{\mu_{k}}\|y_{k+1}-y_{k}\|^{2},$
+$F_{k+1}$ $=f_{k+1}+S_{k+1}-\frac{7\gamma_{k}}{2}\|y_{k+1}-y_{k}\|^{2}-\frac{4\gamma_{k}^{2}}{\mu_{k}}\mathcal{D}^{2}_{y}.$ (21)
+If $\gamma_{k}\geq l_{22}^{k}$, then we have
+$F_{k+1}-F_{k}\leq$ $-\left(2\gamma_{k}-\frac{l_{12}^{k}}{2}\right)\|y_{k+1}-y_{k}\|^{2}$
+$-\left(\beta_{k}-\frac{l_{11}^{k}}{2}-\frac{l_{12}^{k}}{2}-\frac{16(l_{12}^{k})^{2}\gamma_{k-1}}{\mu_{k-1}\mu_{k}}\right)\|x_{k+1}-x_{k}\|^{2},$ (22)
 where $\mathcal{D}_{y}=\max\{\|y_{1}-y_{2}\|\mid y_{1},y_{2}\in\mathcal{Y}\}$.
@@
 Combining (23) and (12), we have
-$\langle\nabla_{y}f(x_{k},y_{k-1})-\nabla_{y}f(x_{k+1},y_{k})+\gamma_{k}(y_{k+1}-y_{k})-\gamma_{k-1}(y_{k}-y_{k-1}),y_{k+1}-y_{k}\rangle\leq 0,$ (24)
+$\langle\nabla_{y}f(x_{k},y_{k-1})-\nabla_{y}f(x_{k+1},y_{k})+\gamma_{k}(y_{k+1}-y_{k})-\gamma_{k-1}(y_{k}-y_{k-1}),y_{k+1}-y_{k}\rangle$
+$\leq 0,$ (24)
 which implies that
-$\gamma_{k}\|y_{k+1}-y_{k}\|^{2}-\gamma_{k-1}\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle\leq\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle.$ (25)
+$\gamma_{k}\|y_{k+1}-y_{k}\|^{2}-\gamma_{k-1}\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle$
+$\leq\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle.$ (25)
 Denoting $v_{k+1}=(y_{k+1}-y_{k})-(y_{k}-y_{k-1})$, and using the identity $\frac{1}{2}\langle v_{k+1},y_{k+1}-y_{k}\rangle=\frac{1}{2}\|y_{k+1}-y_{k}\|^{2}+\frac{1}{2}\|v_{k+1}\|^{2}-\frac{1}{2}\|y_{k}-y_{k-1}\|^{2}$, we conclude that
@@
 $\leq$ $\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle$
-$=$ $\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle+\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),v_{k+1}\rangle$
+$=$ $\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle$
+$+\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),v_{k+1}\rangle$
 $+\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle.$ (27)
 By backtracking strategy (C2) and the Cauchy-Schwarz inequality, we can easily get that
-$\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle\leq\frac{(l_{12}^{k})^{2}}{2\mu_{k}}\|x_{k+1}-x_{k}\|^{2}+\frac{\mu_{k}}{2}\|y_{k+1}-y_{k}\|^{2}.$ (28)
+$\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle$
+$\leq$ $\frac{2(l_{12}^{k})^{2}}{\mu_{k}}\|x_{k+1}-x_{k}\|^{2}+\frac{\mu_{k}}{8}\|y_{k+1}-y_{k}\|^{2}.$ (28)
 Combining (15), (16), (27) and (28), we have
-$\frac{\gamma_{k}}{2}\|y_{k+1}-y_{k}\|^{2}-\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}$ $\leq\frac{(l_{12}^{k})^{2}}{2\mu_{k}}\|x_{k+1}-x_{k}\|^{2}+\left(\frac{\mu_{k}}{2}+\frac{\gamma_{k-1}}{2}-\frac{\gamma_{k}}{2}\right)\|y_{k+1}-y_{k}\|^{2}$
-$\quad-\left(\mu_{k-1}-\frac{(l_{22}^{k-1})^{2}}{2\gamma_{k-1}}\right)\|y_{k}-y_{k-1}\|^{2}.$ (29)
-By multiplying $\frac{4\gamma_{k-1}}{\mu_{k-1}}$ on both sides of (29), using the definition of $S_{k+1}$ and Assumption 2, we obtain
-$\begin{split}S_{k+1}-S_{k}&\leq\frac{2(l_{12}^{k})^{2}\gamma_{k-1}}{\mu_{k-1}\mu_{k}}\|x_{k+1}-x_{k}\|^{2}+2\gamma_{k-1}\|y_{k+1}-y_{k}\|^{2}\\ &\quad-\left({4\gamma_{k-1}}-\frac{2(l_{22}^{k-1})^{2}}{\mu_{k-1}}\right)\|y_{k}-y_{k-1}\|^{2}+\left(\frac{2\gamma_{k}^{2}}{\mu_{k}}-\frac{2\gamma^{2}_{k-1}}{\mu_{k-1}}\right)\mathcal{D}_{y}^{2}.\end{split}$
+$\frac{\gamma_{k}}{2}\|y_{k+1}-y_{k}\|^{2}-\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}$
+$\leq$ $\frac{2(l_{12}^{k})^{2}}{\mu_{k}}\|x_{k+1}-x_{k}\|^{2}+\left(\frac{\mu_{k}}{8}+\frac{\gamma_{k-1}}{2}-\frac{\gamma_{k}}{2}\right)\|y_{k+1}-y_{k}\|^{2}$
+$-\frac{\mu_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}.$ (29)
+By multiplying  $\frac{8\gamma_{k-1}}{\mu_{k-1}}$ on both sides of (29), using the definition of $S_{k+1}$ and Assumption 2, we obtain
+$\begin{split}S_{k+1}-S_{k}&\leq\frac{16(l_{12}^{k})^{2}\gamma_{k-1}}{\mu_{k-1}\mu_{k}}\|x_{k+1}-x_{k}\|^{2}+\gamma_{k-1}\|y_{k+1}-y_{k}\|^{2}\\ &\quad-{4\gamma_{k-1}}\|y_{k}-y_{k-1}\|^{2}+\left(\frac{4\gamma_{k}^{2}}{\mu_{k}}-\frac{4\gamma^{2}_{k-1}}{\mu_{k-1}}\right)\mathcal{D}_{y}^{2}.\end{split}$
 The proof is then completed by Lemma 4 and the definition.
@@
 Suppose that Assumptions 1 and 2 hold.
-Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 1, $\beta_{k}={l_{11}^{k}}+\frac{l_{12}^{k}\mu_{k-1}}{8(l_{22}^{k-1})^{2}}+\frac{32(l_{12}^{k})^{2}(l_{22}^{k-1})^{2}}{\mu_{k}\mu_{k-1}^{2}}$, $\gamma_{k}=\frac{8(l_{22}^{k})^{2}}{\mu_{k}}$, then it holds that
+Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 1, $\beta_{k}={l_{11}^{k}}+l_{12}^{k}+\frac{32(l_{12}^{k})^{2}(l^{k-1}_{12}+l_{22}^{k-1})}{\mu_{k}\mu_{k-1}},\gamma_{k}=l_{12}^{k}+l_{22}^{k}$, then it holds that
 ${T}(\varepsilon)\leq\frac{F_{2}-\underline{F}}{d_{1}\varepsilon^{2}}+1,$ (31)
-where $F_{2}$ is defined in (21), $\underline{F}=\underline{f}-\left(\frac{224L_{22}^{2}}{\mu}+\mu_{1}-\frac{2(l_{22}^{1})^{2}}{\mu_{1}}-\frac{\left(l_{22}^{1}\mu\right)^{2}}{128L_{22}^{2}}+\frac{2^{13}L_{22}^{4}}{\mu^{3}}\right)\mathcal{D}^{2}_{y}$, with $\underline{f}=\min_{(x,y)\in X\times Y}f(x,y)$, $\mathcal{D}_{y}=\max\{\|y_{1}-y_{2}\|\mid y_{1},y_{2}\in\mathcal{Y}\}$, $d_{1}=\min\{\frac{\beta_{\text{min}}}{2\beta_{\text{max}}^{2}+4L^{2}_{12}},\frac{15\mu^{3}+2^{7}l_{22}^{1}\mu}{2^{15}L_{22}^{4}}\}$, $\beta_{\text{min}}={l^{1}_{11}}+\frac{l^{1}_{12}\mu}{64(L_{22})^{2}}+\frac{32(l_{12}^{1})^{2}(l^{1}_{22})^{2}}{\mu_{1}^{3}}$ and $\beta_{\text{max}}=2L_{11}+\frac{\mu_{1}L_{12}}{4(l^{1}_{22})^{2}}+\frac{4096(L_{12})^{2}(L_{22})^{2}}{\mu^{3}}$.
+where $F_{2}$ is defined in (21), $\underline{F}=\underline{f}-\left(14L_{12}+\frac{128L_{12}}{\mu}\right)\mathcal{D}^{2}_{y}$, with $\mathcal{D}_{y}=\max\{\|y_{1}-y_{2}\|\mid y_{1},y_{2}\in\mathcal{Y}\}$, $d_{1}=\min\{\frac{3}{16l^{1}_{12}},\frac{\beta_{\text{min}}}{8L^{2}_{12}},\frac{1}{4\beta_{\max}}\}$, $\underline{f}=\min_{(x,y)\in X\times Y}f(x,y)$, $\beta_{\text{min}}={l^{1}_{11}}+{l^{1}_{12}}+\frac{32(l_{12}^{1})^{2}(l^{1}_{11}+l^{1}_{22})}{\mu_{1}^{2}}$ and $\beta_{\text{max}}=2L_{11}+2{L_{12}}+\frac{2^{9}L_{12}^{2}(L_{12}+L_{22})}{\mu^{2}}$.
 Proof  It follows immediately from Definition 1 and (3) that
@@
 $\|\nabla G_{k}\|^{2}$ $\leq(\beta_{k}^{2}+2L_{12}^{2})\|x_{k+1}-x_{k}\|^{2}+2\gamma_{k}^{2}\|y_{k+1}-y_{k}\|^{2}.$ (34)
-Let $d^{(k)}_{1}={\min\left\{\frac{\frac{2\mu_{k}\gamma_{k}-\left(l^{k}_{22}\right)^{2}}{2\gamma_{k}}+\frac{\mu_{k}\gamma_{k}-4\left(l_{22}^{k}\right)^{2}}{2\mu_{k}}}{2\gamma_{k}^{2}},\frac{\beta_{k}-\frac{l^{k}_{11}}{2}-\frac{l^{k}_{12}}{2\gamma_{k-1}}-\frac{2(l^{k}_{12})^{2}\gamma_{k-1}}{\mu_{k-1}\mu_{k}}}{\beta_{k}^{2}+2L_{12}^{2}}\right\}}>0$.
+Let $d^{(k)}_{1}={\min\left\{\frac{2\gamma_{k}-\frac{l^{k}_{12}}{2}}{2\gamma_{k}^{2}},\frac{\beta_{k}-\frac{l^{k}_{11}}{2}-\frac{l^{k}_{12}}{2}-\frac{16(l^{k}_{12})^{2}\gamma_{k-1}}{\mu_{k-1}\mu_{k}}}{\beta_{k}^{2}+2L_{12}^{2}}\right\}}>0$.
 Multiplying both sides of (34) by $d^{(k)}_{1}$, and using (22) in Lemma 5, we have
@@
 Note that by the definition of $F_{k+1}$ in Lemma 5, we have
-$F_{T(\varepsilon)+1}$ $=f_{T(\varepsilon)+1}+S_{T(\varepsilon)+1}-\frac{2\gamma_{T(\varepsilon)}^{2}}{\mu_{T(\varepsilon)}}\mathcal{D}^{2}_{y}$
-$\quad-\left(\frac{7}{2}\gamma_{T(\varepsilon)+1}+\mu_{T(\varepsilon)+1}-\frac{2(l_{22}^{T(\varepsilon)+1})^{2}}{\mu_{T(\varepsilon)+1}}-\frac{(l_{22}^{T(\varepsilon)+1})^{2}}{2\gamma_{T(\varepsilon)+1}}\right)||y_{T(\varepsilon)+1}-y_{T(\varepsilon)}||^{2}$
-$\geq\underline{f}-\left(\frac{224L_{22}^{2}}{\mu}+\mu_{1}-\frac{2(l_{22}^{1})^{2}}{\mu_{1}}-\frac{\left(l_{22}^{1}\mu\right)^{2}}{128L_{22}^{2}}+\frac{2^{13}L_{22}^{4}}{\mu^{3}}\right)\mathcal{D}^{2}_{y}$
-$=\underline{F},$ (38)
+$F_{T(\varepsilon)+1}$ $=f_{T(\varepsilon)+1}+S_{T(\varepsilon)+1}-\frac{2\gamma_{T(\varepsilon)}^{2}}{\mu_{T(\varepsilon)}}\mathcal{D}^{2}_{y}-\frac{7}{2}\gamma_{T(\varepsilon)+1}||y_{T(\varepsilon)+1}-y_{T(\varepsilon)}||^{2}$
+$\geq\underline{f}-\left(14L_{12}+\frac{128L_{12}}{\mu}\right)\mathcal{D}^{2}_{y}=\underline{F},$ (38)
 where the inequality follows from the definitions of $\underline{f}$ and $\mathcal{D}_{y}$, and the fact that $S_{k}\geq 0$ ($\forall k\geq 1$). We then conclude from (37) that
@@
 Remark 7
-At the $k$-th iteration of Algorithm 1, the number of backtracking step is upper bounded by $\log_{2}\left(\frac{2L_{11}}{l_{11}^{1}}\right)+\log_{2}\left(\frac{2L_{12}}{l_{12}^{1}}\right)+\log_{2}\left(\frac{2L_{22}}{l_{22}^{1}}\right)+\log_{2}\left(\frac{2\mu_{1}}{\mu}\right)=\mathcal{O}(1)$.
-By Theorem 6, the total number of gradient calls of Algorithm 1 to obtain an $\varepsilon$-stationary point that satisfies $\|\nabla G(x_{k},y_{k})\|\leq\varepsilon$ is upper bounded by $\mathcal{O}(L\kappa^{3}\varepsilon^{-2})$ under the nonconvex-strongly concave setting, where $\kappa$ is the condition number.
+In PF-AGP-NSC algorithm, the  total number of backtracking step is upper bounded by  $\sum\limits_{k=2}^{T(\varepsilon)}(\log_{2}\left(\frac{l^{k}_{11}}{l_{11}^{k-1}}\right)+\log_{2}\left(\frac{l^{k}_{12}}{l_{12}^{k-1}}\right)+\log_{2}\left(\frac{l^{k}_{22}}{l_{22}^{k-1}}\right)+\log_{2}\left(\frac{\mu_{k-1}}{\mu_{k}}\right))=\log_{2}\left(\frac{2L_{11}}{l_{11}^{1}}\right)+\log_{2}\left(\frac{2L_{12}}{l_{12}^{1}}\right)+\log_{2}\left(\frac{2L_{22}}{l_{22}^{1}}\right)+\log_{2}\left(\frac{2\mu_{1}}{\mu}\right)=O({\log(\kappa}))$.
+According to Theorem 6,   $d_{1}=O(\frac{1}{L\kappa^{2}})$, $\underline{F}=O(L\kappa)$ , the total number of gradient calls of PF-AGP-NSC to obtain an $\varepsilon$-stationary point that satisfies $\|\nabla G(x_{k},y_{k})\|\leq\varepsilon$ is upper bounded by  $\mathcal{O}(L^{2}\kappa^{3}\varepsilon^{-2})$ under the nonconvex-strongly concave setting, where $\kappa$ is the condition number.
+Remark 8
+If additional information is available, the iteration complexity of the algorithm can be further improved.
+To illustrate this point, if $\mu$ and $\underline{S}$ are known in advance, in Appendix A we propose a restarted variant of the PF-AGP-NSC algorithm, named rPF-AGP-NSC, and show that the gradient complexity of the algorithm can be improved to $\mathcal{O}\left(\log(L)L\kappa^{2}\varepsilon^{-2}\right)$, which is better than SGDA-B.
 2.2 Nonconvex-concave Setting
-In this subsection, we propose a completely parameter-free AGP algorithm for solving nonconvex-concave minimax problems.
-Similar to the idea of the PF-AGP-NSC Algorithm, we estimate the Lipschitz constants $L_{11}$, $L_{12}$ and $L_{22}$ iteratively, however we do not need to estimate the strongly concave constant $\mu$.
-At each iteration of the proposed algorithm, we aim to find a tuple $(l^{k,i}_{11},l^{k,i}_{12},l^{k,i}_{22})$ by backtracking such that conditions (C1), (C2), (C3) and the following condition (C5) are satisfied at $(x_{k,i},y_{k,i})$:
-$C_{5}^{k,i}$ $=\langle\nabla_{y}f_{k,i}(x_{k,i},y_{k,i})-\nabla_{y}f_{k,i}(x_{k,i},y_{k}),y_{k,i}-y_{k}\rangle+\frac{c^{2}_{k,i}+c_{k,i}{l}^{k,i}_{22}}{{l}^{k,i}_{22}+2c_{k,i}}\|y_{k,i}-y_{k}\|^{2}$
-$\qquad+\frac{1}{{l}^{k,i}_{22}+2c_{k,i}}\|\nabla_{y}f_{k,i}(x_{k,i},y_{k,i})-\nabla_{y}f_{k,i}(x_{k,i},y_{k})\|^{2}\leq 0.$ (C5)
-The proposed algorithm, denoted as PF-AGP-NC, is formally presented in Algorithm 2.
-Please note that the existing SGDA-B algorithm (Xu et al., 2023) for solving nonconvex-concave minimax problems requires the realization of the diameter of the feasible set $\mathcal{Y}$ in order to construct the key auxiliary problem, while the PF-AGP-NC algorithm we propose is a completely parameter-free algorithm that does not require any prior parameters.
 Algorithm 2 A parameter-free alternating gradient projection (PF-AGP-NC) algorithm for nonconvex-concave minimax problems
-  Step 1: Input $x_{1},y_{1}$, $\beta_{1}$, $\gamma_{1}$, $c_{1}$, $l_{11}^{1}$, $l_{12}^{1}$, $l_{22}^{1}$; Set $k=1$.
+  Step 1: Input $x_{1},y_{1}$, $\beta_{0}$, $\gamma_{0}$, $c_{0}$, $l_{11}^{0}$, $l_{12}^{0}$, $l_{22}^{0}$; Set $k=1$.
   Step 2: Updata $x_{k}$ and $y_{k}$:
-   (a): Set $i=1$, $l_{11}^{k,i}=l_{11}^{k}$, $l_{12}^{k,i}=l_{12}^{k}$, $l_{22}^{k,i}=l_{22}^{k}$, $\beta_{k,i}=\beta_{k}$, $\gamma_{k,i}=\gamma_{k}$, $c_{k,i}=c_{k}$.
+   (a): Set $i=0$$l_{11}^{k,i}={l_{11}^{k-1}}$, $l_{12}^{k,i}={l_{12}^{k-1}}$, $l_{22}^{k,i}=l_{22}^{k-1}$, $\beta_{k,i}=\beta_{k-1}$, $\gamma_{k,i}=\gamma_{k-1}$, $c_{k,i}=c_{k-1}$.
    (b): Update $x_{{k,i}}$ and $y_{{k,i}}$:
@@
 $y_{{k,i}}$ $=\mathcal{P}_{\mathcal{Y}}\left(y_{k}+\frac{1}{\gamma_{k,i}}\nabla_{y}f(x_{{k,i}},y_{k})-\frac{1}{\gamma_{k,i}}{c_{k,i}}y_{k}\right).$ (43)
-   (c): Compute $C_{1}^{k,i}$, $C_{2}^{k,i}$, $C_{3}^{k,i}$, $C_{5}^{k,i}$ as in (C1), (C2), (C3), (C5);
+   (c):  Compute $C_{1}^{k,i}$, $C_{2}^{k,i}$, $C_{5}^{k,i}$, as in (2.1), (C2), (C5);
    (d): Update $l^{k,i}_{11}$, $l^{k,i}_{12}$, $l^{k,i}_{22}$:
 $l^{k,i+1}_{11}$ $=\frac{\mbox{sgn}(C_{1}^{k,i})+3}{2}l^{k,i}_{11},\quad l^{k,i+1}_{12}=\frac{\mbox{sgn}(C_{2}^{k,i})+3}{2}l^{k,i}_{12},$ (44)
-$l^{k,i+1}_{22}$ $=\frac{\max\{\mbox{sgn}(C_{3}^{k,i}),\mbox{sgn}(C_{5}^{k,i})\}+3}{2}l^{k,i}_{22}.$ (45)
-   (e): If $C_{1}^{k,i}\leq 0$, $C_{2}^{k,i}\leq 0$, $C_{3}^{k,i}\leq 0$ and $C_{5}^{k,i}\leq 0$, then
-     $x_{k+1}=x_{k,i}$, $y_{k+1}=y_{k,i}$, $l_{11}^{k}=l_{11}^{k,i+1}$, $l_{12}^{k}=l_{12}^{k,i+1}$, $l_{22}^{k}=l_{22}^{k,i+1}$, $\beta_{k}=\beta_{k,i}$, $\gamma_{k}=\gamma_{k,i}$,
-     $c_{k}=c_{k,i}$, go to Step 3;
+$l^{k,i+1}_{22}$ $=\frac{\mbox{sgn}(C_{5}^{k,i})+3}{2}l^{k,i}_{22}.$ (45)
+   (e):  If $C_{1}^{k,i}\leq 0$, $C_{2}^{k,i}\leq 0$ and $C_{5}^{k,i}\leq 0$, then
+     $x_{k+1}=x_{k,i}$, $y_{k+1}=y_{k,i}$, $l_{11}^{k}=l_{11}^{k,i+1}$, $l_{12}^{k}=l_{12}^{k,i+1}$, $l_{22}^{k}=l_{22}^{k,i+1}$, $\beta_{k}=\beta_{k,i}$,
+     $\gamma_{k}=\gamma_{k,i}$, $c_{k}=c_{k,i}$, go to Step 3;
      Otherwise, $i=i+1$,
-$\beta_{k,i}=\frac{l_{12}^{k,i}}{20l^{k-1,i}_{22}}+\frac{1280(l_{12}^{k,i})^{2}\sqrt{k}}{19^{2}l_{12}^{k-1,i}},\quad\gamma_{k,i}=20l^{k,i}_{22},\quad c_{k,i}=\frac{19l^{k,i}_{22}}{k^{1/4}},$ (46)
+$\beta_{k,i}=\frac{l_{12}^{k,i}}{20l^{k-1,i}_{22}}+\frac{2(l_{12}^{k,i})^{2}\sqrt{k}}{l_{12}^{k-1,i}},\quad\gamma_{k,i}=20l^{k,i}_{22},\quad c_{k,i}=\frac{19l^{k,i}_{22}}{k^{1/4}},$ (46)
      go to Step 2(b).
   Step 3: If some stationary condition is satisfied, stop; Otherwise, set $k=k+1$, go to Step 2.
+In this subsection, we propose a completely parameter-free AGP algorithm for solving nonconvex-concave minimax problems.
+Similar to the idea of the PF-AGP-NSC Algorithm, we estimate the Lipschitz constants $L_{11}$, $L_{12}$ and $L_{22}$ iteratively, however we do not need to estimate the strongly concavity modulus $\mu$.
+At each iteration of the proposed algorithm, we aim to find a tuple $(l^{k,i}_{11},l^{k,i}_{12})$ by backtracking such that conditions (2.1), (C2).
+The proposed algorithm, denoted as PF-AGP-NC, is formally presented in Algorithm 2.
+Please note that the existing SGDA-B algorithm (Xu et al., 2024) for solving nonconvex-concave minimax problems requires the realization of the diameter of the feasible set $\mathcal{Y}$ in order to construct the key auxiliary problem, while the proposed PF-AGP-NC algorithm is a completely parameter-free algorithm that does not require any prior parameters.
 2.2.1 Complexity Analysis
 In this subsection, we analyze the iteration complexity of Algorithm 2 for solving (P) under the general nonconvex-concave setting.
-The backtracking strategy (C5) is a key inequality that we will use to establish some important recursions for the PF-AGP-NC algorithm under the nonconvex-concave setting in the following two lemmas.
+Similar to the idea of the PF-AGP-NSC Algorithm, we estimate the Lipschitz constants $L_{11}$, $L_{12}$ and $L_{22}$ iteratively, however we do not need to estimate the strongly concavity modulus $\mu$.
+At each iteration of the proposed algorithm, we aim to find a tuple $(l^{k,i}_{11},l^{k,i}_{12},l^{k,i}_{22})$ by backtracking such that conditions (2.1), (C2) and the following condition (C5) are satisfied:
+${C}_{5}^{k,i}=$ $(l_{22}^{k,i}+c_{k,i})\langle\nabla_{y}f_{k,i}(x_{k,i},y_{k,i})-\nabla_{y}f_{k,i}(x_{k,i},y_{k}),y_{k,i}-y_{k}\rangle$
+$+\|\nabla_{y}f_{k,i}(x_{k,i},y_{k,i})-\nabla_{y}f_{k,i}(x_{k,i},y_{k})\|^{2}\leq 0$ (C5)
+The backtracking strategy (C5) is a key inequality that we will use in the following two lemmas to establish some important recursions of the PF-AGP-NC algorithm in the nonconvex-concave setting.
 This is also one of the main differences between proofs in the nonconvex-strongly concave setting and proofs in the nonconvex-concave setting.
 Lemma 9
-Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 2. If $\forall k$, $\beta_{k}>l^{k}_{11}$ and $\gamma_{k}\geq\frac{{l}_{22}^{k}+2c_{k}}{2}$, then
+Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 2. If $\forall k$, $\beta_{k}>l^{k}_{11}$ and  $\gamma_{k}\geq{{l}_{22}^{k}+c_{k}}$, then
 $f({x_{k+1}},y_{k+1})-f({x_{k}},y_{k})$
@@
 The concavity of $f_{k}(x_{k+1},y)$ with respect to $y$ together with (48) then imply that
-$f_{k}(x_{k+1},y_{k+1})-f_{k}(x_{k+1},y_{k})$ $\leq\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k}),y_{k+1}-y_{k}\rangle$
-$\quad+\langle\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle$
+$f_{k}(x_{k+1},y_{k+1})-f_{k}(x_{k+1},y_{k})$
+$\leq$ $\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k}),y_{k+1}-y_{k}\rangle$
+$+\langle\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle$
 $=\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k}),y_{k+1}-y_{k}\rangle$
-$\quad+\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),v_{k+1}\rangle$
-$\quad+\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle$
-$\quad+\gamma_{k-1}\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle,$ (49)
+$+\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),v_{k+1}\rangle$
+$+\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle$
+$+\gamma_{k-1}\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle,$ (49)
 where $v_{k+1}=y_{k+1}-y_{k}-(y_{k}-y_{k-1})$.
@@
 $\leq$ $\frac{1}{2\gamma_{k-1}}\|\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1})\|^{2}+\frac{\gamma_{k-1}}{2}\|v_{k+1}\|^{2}.$ (51)
-Thirdly, by (C4) we have that
+Thirdly, by (C5) , we have that
 $\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle$
-$\leq$ $-\frac{1}{{l}_{22}^{k-1}+2c_{k-1}}\|\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1})\|^{2}$
-$\quad-\frac{c_{k-1}{l}_{22}^{k-1}+c^{2}_{k-1}}{{l}_{22}^{k-1}+2c_{k-1}}\|y_{k}-y_{k-1}\|^{2}$
-$\leq$ $-\frac{1}{{l}_{22}^{k-1}+2c_{k-1}}\|\nabla_{y}f_{k}(x_{k},y_{k})-\nabla_{y}f_{k}(x_{k},y_{k-1})\|^{2}.$ (52)
+$\leq$ $-\frac{1}{{l}_{22}^{k-1}+c_{k-1}}\|\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1})\|^{2}.$ (52)
 We also have the following equation that
-$\begin{split}\gamma_{k-1}\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle=\frac{\gamma_{k-1}}{2}\|y_{k+1}-y_{k}\|^{2}+\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}-\frac{\gamma_{k-1}}{2}\|v_{k+1}\|^{2}.\end{split}$ (53)
-By $\gamma_{k}\geq\frac{{l}_{22}^{k}+2c_{k}}{2}$, plugging (50)-(53) into (49), and using the definition of $f_{k}(x_{k+1},y_{k+1})$ and $f_{k}(x_{k+1},y_{k})$, we obtain
+$\gamma_{k-1}\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle$
+$=$ $\frac{\gamma_{k-1}}{2}\|y_{k+1}-y_{k}\|^{2}+\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}-\frac{\gamma_{k-1}}{2}\|v_{k+1}\|^{2}.$ (53)
+By $\gamma_{k}\geq{{l}_{22}^{k}+c_{k}}$, plugging (50)-(53) into (49), and using the definition of $f_{k}(x_{k+1},y_{k+1})$ and $f_{k}(x_{k+1},y_{k})$, we obtain
 $f(x_{k+1},y_{k+1})-f(x_{k+1},y_{k})$
@@
 If
-$\beta_{k}>l^{k}_{11},\quad\frac{\gamma_{k+1}}{c_{k+1}}-\frac{\gamma_{k}}{c_{k}}\leq\frac{1}{5},\quad\gamma_{k}\geq\frac{{l}^{k}_{22}+2c_{k}}{2},$ (58)
+$\beta_{k}>l^{k}_{11},\quad\frac{\gamma_{k+1}}{c_{k+1}}-\frac{\gamma_{k}}{c_{k}}\leq\frac{1}{5},\quad{c_{k+1}}-{c_{k}}\leq{\gamma_{k+1}-\gamma_{k}},\quad\gamma_{k}\geq{{l}^{k}_{22}+c_{k}},$ (58)
 then $\forall k\geq 2$,
-$F_{k+1}-F_{k}=$ $-\left(\frac{9\gamma_{k}}{10}-\frac{c_{k}}{2}\right)\|y_{k+1}-y_{k}\|^{2}$
-$-\left(+\frac{\beta_{k}}{2}-\frac{(l^{k}_{12})^{2}}{2\gamma_{k-1}}-\frac{16(l^{k}_{12})^{2}\gamma_{k-1}}{c_{k-1}^{2}}\right)\|x_{k+1}-x_{k}\|^{2}$
+$F_{k+1}-F_{k}=$ $-\frac{19\gamma_{k}}{10}\|y_{k+1}-y_{k}\|^{2}-\left(\frac{\beta_{k}}{2}-\frac{(l^{k}_{12})^{2}}{2\gamma_{k-1}}-\frac{16(l^{k}_{12})^{2}\gamma_{k-1}}{c_{k-1}^{2}}\right)\|x_{k+1}-x_{k}\|^{2}$
 $+8\left({\gamma_{k}}\frac{c_{k+1}}{c_{k}}-{\gamma_{k-1}}\frac{c_{k}}{c_{k-1}}\right)\|y_{k+1}\|^{2}+\frac{c_{k-1}-c_{k}}{2}\|y_{k+1}\|^{2}.$ (59)
@@
 By (48) and (60), we have
-$\langle\nabla_{y}f_{k-1}(x_{k},y_{k-1})-\nabla_{y}f_{k}(x_{k+1},y_{k})+$ $\gamma_{k}(y_{k+1}-y_{k})-\gamma_{k-1}(y_{k}-y_{k-1}),y_{k+1}-y_{k}\rangle\leq 0,$
+$\langle\nabla_{y}f_{k-1}(x_{k},y_{k-1})-\nabla_{y}f_{k}(x_{k+1},y_{k})+\gamma_{k}(y_{k+1}-y_{k})-\gamma_{k-1}(y_{k}-y_{k-1}),y_{k+1}-y_{k}\rangle$
+$\leq 0,$
 and
@@
 $\leq$ $\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle$
-$=$ $\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k}),y_{k+1}-y_{k}\rangle+\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),v_{k+1}\rangle$
+$=$ $\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k}),y_{k+1}-y_{k}\rangle$
+$+\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),v_{k+1}\rangle$
 $+\langle\nabla_{y}f_{k}(x_{k-1},y_{k-1})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle.$ (62)
-Using an argument similar to the proof of (50)-(53), by (C5) and the Cauchy-Schwarz inequality, we conclude from the above inequality that
+Using an argument similar to the proof of (50)-(53), by (C5), $f_{k}(x,y)$ is concave with respect to $y$ and the Cauchy-Schwarz inequality, we conclude from the above inequality that
 $\left(\gamma_{k}-\frac{\gamma_{k-1}}{2}\right)\|y_{k+1}-y_{k}\|^{2}+\frac{\gamma_{k-1}}{2}\|v_{k+1}\|^{2}-\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}$
@@
 $+\frac{\gamma_{k-1}}{2}\|v_{k+1}\|-\frac{1}{{l}_{22}^{k-1}+2c_{k-1}}\|\nabla_{y}f_{k}(x_{k},y_{k})-\nabla_{y}f_{k}(x_{k},y_{k-1})\|^{2}$
-$-\frac{c_{k-1}{l}_{22}^{k-1}+c^{2}_{k-1}}{{l}_{22}^{k-1}+2c_{k-1}}\|y_{k}-y_{k-1}\|^{2},$ (63)
-for any $a_{k}>0$.
-Obviously, we have $-\frac{c_{k-1}{l}^{k-1}_{22}+c^{2}_{k-1}}{2c_{k-1}+{l}^{k-1}_{22}}\leq-\frac{c_{k-1}{l}^{k-1}_{22}+c^{2}_{k-1}}{2({l}^{k-1}_{22}+c_{k-1})}=-\frac{c_{k-1}}{2}.$ Combining $\gamma_{k}\geq\frac{{l}^{k}_{22}+2c_{k}}{2}$ and rearranging the terms in (63), we obtain
+$-\frac{c_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2},$ (63)
+for any $a_{k}>0$. Combining $\gamma_{k}\geq{{l}^{k}_{22}+c_{k}}$ and rearranging the terms in (63), we obtain
 $\frac{\gamma_{k}}{2}\|y_{k+1}-y_{k}\|^{2}+\frac{c_{k}-c_{k-1}}{2}\|y_{k+1}\|^{2}$
-$\quad\leq\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}+\left(\frac{\alpha_{k}}{2}+\frac{\gamma_{k-1}}{2}-\frac{\gamma_{k}}{2}\right)\|y_{k+1}-y_{k}\|^{2}+\frac{c_{k}-c_{k-1}}{2}\|y_{k}\|^{2}$
-$\quad\quad+\frac{(l^{k}_{12})^{2}}{2a_{k}}\|x_{k+1}-x_{k}\|^{2}-\frac{c_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}.$ (64)
+$\quad\leq\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}+\left(\frac{a_{k}}{2}+\frac{\gamma_{k-1}}{2}-\frac{\gamma_{k}}{2}+\frac{c_{k}}{2}-\frac{c_{k-1}}{2}\right)\|y_{k+1}-y_{k}\|^{2}+\frac{c_{k}-c_{k-1}}{2}\|y_{k}\|^{2}$
+$\quad\quad+\frac{(l^{k}_{12})^{2}}{2a_{k}}\|x_{k+1}-x_{k}\|^{2}-\frac{c_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}$
+$\quad\leq\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}+\frac{a_{k}}{2}\|y_{k+1}-y_{k}\|^{2}+\frac{c_{k}-c_{k-1}}{2}\|y_{k}\|^{2}+\frac{(l^{k}_{12})^{2}}{2a_{k}}\|x_{k+1}-x_{k}\|^{2}$ (64)
+$\quad\quad-\frac{c_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}.$ (65)
 By multiplying $\frac{16\gamma_{k-1}}{c_{k-1}}$ on both sides of the above inequality, setting $a_{k}=\frac{c_{k}}{2}$, and using the definition of $S_{k+1}$, (58) and $\{\gamma_{k}\}$ is nodecreasing, we then obtain
@@
 Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 2.
-If $\beta_{k}=\frac{l_{12}^{k}}{20l^{k-1}_{22}}+\frac{1280(l_{12}^{k})^{2}\sqrt{k}}{19^{2}l_{12}^{k-1}}$, $\gamma_{k}=20l^{k}_{22}$ , $c_{k}=\frac{19l^{k}_{22}}{k^{1/4}}$, then for any given $\varepsilon>0$, $k\geq 2$, we have
-$T(\varepsilon)\leq\max\left(\left(\frac{5120\left(L_{12}\right)^{2}d_{2}d_{3}}{19^{2}l_{22}^{1}\varepsilon^{2}}+1\right)^{2},\frac{76^{4}L_{22}^{4}\hat{\mathcal{D}}_{y}^{4}}{\varepsilon^{4}}\right),$
-where $d_{2}=F_{2}-\underline{F}+339L_{22}\hat{\mathcal{D}}_{y}^{2}$, $d_{3}=\max\{\bar{d}_{1},\frac{20\cdot 19^{2}L_{22}^{2}}{17(l_{12}^{1})^{2}}\}$, $\hat{\mathcal{D}}_{y}=\max\{\|y\|\,|y\in Y\}$ with $\underline{F}={\underline{f}}-639L_{22}\hat{\mathcal{D}}_{y}^{2}$, $\underline{f}=\min_{(x,y)\in X\times Y}f(x,y)$, $\bar{d}_{1}=32+\frac{1+1600L^{2}_{22}}{128({l}_{12}^{1})^{2}}$.
+If $\beta_{k}=\frac{l_{12}^{k}}{20l^{k-1}_{22}}+\frac{2(l_{12}^{k})^{2}\sqrt{k}}{l_{22}^{k-1}}$, $\gamma_{k}=20l^{k}_{22}$ , $c_{k}=\frac{19l^{k}_{22}}{k^{1/4}}$, then for any given $\varepsilon>0$, $k\geq 2$, we have
+$T(\varepsilon)\leq\max\left(\left(\frac{\hat{\alpha}d_{2}d_{3}}{\varepsilon^{2}}+1\right)^{2},\frac{76^{4}L_{22}^{4}\hat{\mathcal{D}}_{y}^{4}}{\varepsilon^{4}}\right),$
+where $d_{2}=F_{2}-\underline{F}+339L_{22}\hat{\mathcal{D}}_{y}^{2}$, $d_{3}=\max\{\bar{d}_{1},\max\limits_{{k=2,\cdots,T(\varepsilon)}}\{\frac{19^{2}\cdot 20^{2}l_{22}^{k}l_{22}^{k-1}}{16(l_{12}^{k})^{2}\}}\}\}$, $\hat{\mathcal{D}}_{y}=\max\{\|y\|\,|y\in Y\}$ with $\underline{F}={\underline{f}}-639L_{22}\hat{\mathcal{D}}_{y}^{2}$, $\underline{f}=\min_{(x,y)\in X\times Y}f(x,y)$, $\hat{\alpha}=\max\limits_{{k=2,\cdots,T(\varepsilon)}}\{\frac{2^{8}\cdot 5^{2}(l^{k}_{12})^{2}}{19^{2}\cdot 20l_{22}^{k-1}}\}$ and $\bar{d}_{1}=\max\limits_{{k=2,\cdots,T(\varepsilon)}}\left\{4+\frac{19^{4}}{2^{15}\cdot 5^{4}}+\frac{19^{4}\cdot 20^{2}(l_{22}^{k})^{2}}{2^{15}\cdot 5^{4}\cdot(l_{12}^{k})^{2}}\right\}$.
 Proof  We can easily see that the relations in (58) are satisfied by the settings of $c_{k}$, $\gamma_{k}$ and $\beta_{k}$.
@@
 $\frac{\beta_{k}}{2}-\frac{(l^{k}_{12})^{2}}{2\gamma_{k-1}}-\frac{16(l^{k}_{12})^{2}\gamma_{k-1}}{c_{k-1}^{2}}=\alpha_{k}.$ (67)
-This observation, in view of Lemma 8, $\frac{c_{k}}{2}\leq\frac{19\gamma_{k}}{40k^{1/4}}\leq\frac{19\gamma_{k}}{40}$ implies that
-$\begin{split}&{\alpha_{k}}\left\|x_{k+1}-x_{k}\right\|^{2}+\frac{17\gamma_{k}}{40}\left\|y_{k+1}-y_{k}\right\|^{2}\\ \leq&F_{k}-F_{k+1}+8\left({\gamma_{k}}\frac{c_{k+1}}{c_{k}}-{\gamma_{k-1}}\frac{c_{k}}{c_{k-1}}\right)\|y_{k+1}\|^{2}+\frac{c_{k-1}-c_{k}}{2}\|y_{k+1}\|^{2}.\end{split}$ (67)
+This observation, in view of Lemma 9, implies that
+${\alpha_{k}}\left\|x_{k+1}-x_{k}\right\|^{2}+\frac{19\gamma_{k}}{10}\left\|y_{k+1}-y_{k}\right\|^{2}$
+$\leq$ $F_{k}-F_{k+1}+8\left({\gamma_{k}}\frac{c_{k+1}}{c_{k}}-{\gamma_{k-1}}\frac{c_{k}}{c_{k-1}}\right)\|y_{k+1}\|^{2}+\frac{c_{k-1}-c_{k}}{2}\|y_{k+1}\|^{2}.$ (68)
 We can easily check from the definition of $f_{k}(x,y)$ that
@@
 $\frac{(\beta_{k})^{2}+2({l^{k}_{12}})^{2}}{\alpha_{k}^{2}}$ $=\frac{\left(\frac{{l^{k}_{12}}}{\gamma_{k-1}}+\frac{64({l}_{12}^{k})^{2}\gamma_{k-1}}{c_{k-1}^{2}}\right)^{2}+2({l}^{k}_{12})^{2}}{\alpha_{k}^{2}}$
-$\leq 32+\frac{1+1600L^{2}_{22}}{128({l}_{12}^{1})^{2}}=\bar{d}_{1}.$ (72)
+$\leq\max\limits_{k=2,3\cdots T(\varepsilon)}\{4+\frac{19^{4}}{2^{15}\cdot 5^{4}}+\frac{19^{4}\gamma^{2}_{k}}{2^{15}\cdot 5^{4}({l}_{12}^{k})^{2}}\}=\bar{d}_{1}.$ (73)
 Combining the previous two inequalities in (72) and (73), we obtain
 $\|\nabla{\tilde{G}_{k}}\|^{2}\leq{\bar{d}_{1}(\alpha_{k})^{2}}\|x_{k+1}-x_{k}\|^{2}+2\gamma_{k}^{2}\|y_{k+1}-y_{k}\|^{2}.$ (74)
-Denote $d_{k}^{(2)}=\frac{1}{\max\{\bar{d}_{1}\alpha_{k},{\frac{17}{80}\gamma_{k}}\}}$, $\tilde{c}_{k}=\frac{19}{20k^{1/4}}$, by multiplying $d^{(2)}_{k}$ on the both sides of (73), and using (67), we have
+Denote  $d_{k}^{(2)}=\frac{1}{\max\{\bar{d}_{1}\alpha_{k},{\frac{19}{20}\gamma_{k}}\}}$, $\tilde{c}_{k}=\frac{19}{20k^{1/4}}$, by multiplying $d^{(2)}_{k}$ on the both sides of (74), and using (68), we have
 $d_{k}^{(2)}\|\nabla\tilde{G}_{k}\|^{2}$
@@
 summing both sides of (75) from $k=2$ to $k=\tilde{T}(\varepsilon)$, we then obtain
-$\sum_{k=2}^{\tilde{T}{(\varepsilon)}}d^{(2)}_{k}\left\|\nabla\tilde{G}_{k}\right\|^{2}$ $\leq F_{2}-F_{\tilde{T}(\varepsilon)}+8\left(\frac{\gamma_{\tilde{T}(\varepsilon)-1}c_{\tilde{T}(\varepsilon)}}{c_{\tilde{T}(\varepsilon)-1}}-\frac{\gamma_{1}c_{2}}{c_{1}}\right)\hat{\mathcal{D}}_{y}^{2}+40L_{22}\left(\frac{\tilde{c}_{1}}{2}-\frac{\tilde{c}_{\tilde{T}(\varepsilon)}}{2}\right)\hat{\mathcal{D}}^{2}_{y}$
+$\sum_{k=2}^{\tilde{T}{(\varepsilon)}}d^{(2)}_{k}\left\|\nabla\tilde{G}_{k}\right\|^{2}$
+$\leq$ $F_{2}-F_{\tilde{T}(\varepsilon)}+8\left(\frac{\gamma_{\tilde{T}(\varepsilon)-1}c_{\tilde{T}(\varepsilon)}}{c_{\tilde{T}(\varepsilon)-1}}-\frac{\gamma_{1}c_{2}}{c_{1}}\right)\hat{\mathcal{D}}_{y}^{2}+40L_{22}\left(\frac{\tilde{c}_{1}}{2}-\frac{\tilde{c}_{\tilde{T}(\varepsilon)}}{2}\right)\hat{\mathcal{D}}^{2}_{y}$
 $\leq$ $F_{2}-F_{\tilde{T}(\varepsilon)}+339L_{22}\hat{\mathcal{D}}_{y}^{2}.$ (77)
@@
 $\sum_{k=2}^{\infty}d^{(2)}_{k}\left\|\nabla\tilde{G}_{k}\right\|^{2}\leq F_{2}-\underline{F}+339L_{22}\hat{\mathcal{D}}_{y}^{2}=d_{2}.$ (79)
-We can see from the selection of $d_{3}$ that $d_{3}\geq\max\{\bar{d}_{1},\frac{80\gamma_{k}}{17\alpha_{k}}\}$, which implies $d^{(2)}_{k}\geq\frac{1}{d_{3}\alpha_{k}}$, by multiplying $d_{3}$ on the both sides of (78), and combining the definition of $d_{2}$, we have $\sum_{k=2}^{\infty}\frac{1}{\alpha_{k}}\left\|\nabla\tilde{G}_{k}\right\|^{2}\leq d_{2}d_{3}$, which, by the definition of $\tilde{T}(\varepsilon)$, implies that
+We can see from the selection of $d_{3}$ that  $d_{3}\geq\max\{\bar{d}_{1},\frac{19\gamma_{k}}{20\alpha_{k}}\}$, which implies $d^{(2)}_{k}\geq\frac{1}{d_{3}\alpha_{k}}$, by multiplying $d_{3}$ on the both sides of (79), and combining the definition of $d_{2}$, we have $\sum_{k=2}^{\infty}\frac{1}{\alpha_{k}}\left\|\nabla\tilde{G}_{k}\right\|^{2}\leq d_{2}d_{3}$, which, by the definition of $\tilde{T}(\varepsilon)$, implies that
 $\frac{\varepsilon^{2}}{4}\leq\frac{d_{2}d_{3}}{\sum_{k=2}^{\tilde{T}{(\varepsilon)}}\frac{1}{\alpha_{k}}}\leq\frac{d_{2}d_{3}}{\sum_{k=2}^{\tilde{T}{(\varepsilon)}}\frac{1}{\bar{\alpha}}}.$ (80)
-By using the fact $\sum_{k=2}^{\tilde{T}{(\varepsilon)}}\frac{1}{\sqrt{k}}\geq\sqrt{\tilde{T}{(\varepsilon)}}-1$ and (79), we conclude $\frac{\varepsilon^{2}}{4}\leq\frac{1280\cdot\left(L_{12}\right)^{2}d_{2}d_{3}}{19^{2}l_{22}^{1}\left(\sqrt{\tilde{T}(\varepsilon)}-1\right)}$ or equivalently,
-$\tilde{T}(\varepsilon)\leq\left(\frac{5120\left(L_{12}\right)^{2}d_{2}d_{3}}{19^{2}l_{22}^{1}\varepsilon^{2}}+1\right)^{2}.$ (80)
+Where $\bar{\alpha}={(k-1)^{1/2}}\hat{\alpha}$, by using the fact $\sum_{k=2}^{\tilde{T}{(\varepsilon)}}\frac{1}{\sqrt{k}}\geq\sqrt{\tilde{T}{(\varepsilon)}}-1$ and (80), we conclude  $\frac{\varepsilon^{2}}{4}\leq\frac{\hat{\alpha}d_{2}d_{3}}{\sqrt{\tilde{T}(\varepsilon)}-1}$ or equivalently,
+$\tilde{T}(\varepsilon)\leq\left(\frac{\hat{\alpha}d_{2}d_{3}}{\varepsilon^{2}}+1\right)^{2}.$ (81)
 On the other hand, if $k\geq\frac{76^{4}L_{22}^{4}\hat{\mathcal{D}}_{y}^{4}}{\varepsilon^{4}}$, then $c_{k}=\frac{19\gamma_{k}}{20k^{1/4}}\leq\frac{\varepsilon}{2\hat{{\mathcal{D}_{y}}}}$.
@@
 $T(\varepsilon)$ $\leq\max\left(\tilde{T}{(\varepsilon)},\frac{76^{4}L_{22}^{4}\hat{\mathcal{D}}_{y}^{4}}{\varepsilon^{4}}\right)$
-$\leq\max\left(\left(\frac{5120\left(L_{12}\right)^{2}d_{2}d_{3}}{19^{2}l_{22}^{1}\varepsilon^{2}}+1\right)^{2},\frac{76^{4}L_{22}^{4}\hat{\mathcal{D}}_{y}^{4}}{\varepsilon^{4}}\right),$ (81)
+$\leq\max\left(\left(\left(\frac{\hat{\alpha}d_{2}d_{3}}{\varepsilon^{2}}+1\right)^{2}+1\right)^{2},\frac{76^{4}L_{22}^{4}\hat{\mathcal{D}}_{y}^{4}}{\varepsilon^{4}}\right),$ (82)
 such that $\left\|\nabla G_{k}\right\|\leq\left\|\nabla\tilde{G}_{k}\right\|+c_{k}\left\|y_{k}\right\|\leq\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon$.
 Remark 12
-At the $k$-th iteration of Algorithm 2, the number of backtracking step is upper bounded by $\log_{2}\left(\frac{2L_{11}}{l_{11}^{1}}\right)+\log_{2}\left(\frac{2L_{12}}{l_{12}^{1}}\right)+\log_{2}\left(\frac{2L_{22}}{l_{22}^{1}}\right)=\mathcal{O}(1)$.
-According to Theorem 10, the total number of gradient calls of Algorithm 2 to obtain an $\varepsilon$-stationary point that satisfies $\|\nabla G(x_{k},y_{k})\|\leq\varepsilon$ is $\mathcal{O}(L^{4}\varepsilon^{-4})$ under the nonconvex-concave setting.
+In PF-AGP-NC algorithm, the total number of backtracking step is upper bounded by $\sum\limits_{k=2}^{T(\varepsilon)}\left(\log_{2}\left(\frac{l^{k}_{11}}{l_{11}^{k-1}}\right)+\log_{2}\left(\frac{l^{k}_{12}}{l_{12}^{k-1}}\right)+\log_{2}\left(\frac{l^{k}_{22}}{l_{22}^{k-1}}\right)\right)=\log_{2}\left(\frac{2L_{11}}{l_{11}^{1}}\right)+\log_{2}\left(\frac{2L_{12}}{l_{12}^{1}}\right)+\log_{2}\left(\frac{2L_{22}}{l_{22}^{1}}\right)=\mathcal{O}(\log(L))$.
+According to Theorem 11, $d_{2}=O(L)$, $\bar{d}_{1}=O(1)$, $d_{3}=O(1)$, $\hat{\alpha}=O(\log(L)L)$, hence the total number of gradient calls of PF-AGP-NC to obtain an $\varepsilon$-stationary point that satisfies $\|\nabla G(x_{k},y_{k})\|\leq\varepsilon$ is $\mathcal{O}(\log^{2}(L)L^{4}\varepsilon^{-4})$ under the nonconvex-concave setting.
+Remark 13
+In the nonconvex-concave setting, when $\varepsilon<\frac{L}{{2}^{\log^{2}(L)L}}$ holds, $\log^{2}(L)L^{4}\varepsilon^{-4}<L^{3}\log(L\varepsilon^{-1})\varepsilon^{-4}$, which means that the total number of gradient calls for PF-AGP-NC algorithm, i.e., $\mathcal{O}(\log^{2}(L)L^{4}\varepsilon^{-4})$, which is asymptotically smaller than the total number of gradient calls for SGDA-B algorithm, i.e., $\mathcal{O}(L^{3}\log(L\varepsilon^{-1})\varepsilon^{-4})$, especially when $\varepsilon\to 0^{+}$.
+On the other hand, SGDA-B algorithm requires additional prior information, notably $\underline{S}$ and $\mathcal{D}_{y}$.
+3 A completely parameter-free single-loop algorithm for nonconvex-linear minimax problems
+In this section, we propose a completely parameter-free AGP algorithm, denoted as PF-AGP-NL, for solving nonconvex-linear minimax problems.
+The proposed algorithm is similar to PF-AGP-NC, while utilizes the gradient of a regularized variant of the original function, i.e.,
+$\tilde{f}_{k}(x,y)=f(x,y)-\frac{c_{k}}{2}\|y\|^{2}-\frac{d_{k}}{2}\|y-y_{k}\|^{2},$
+where $c_{k}\geq 0$ and $d_{k}\geq 0$ are regularization parameters.
+At the $k$th iteration, the variable $x$ is updated by minimizing a linear approximation of $f(x,y_{k})$ plus a regularization term, i.e.,
+$x_{k+1}$ $=\arg\min_{x\in\mathcal{X}}\left\langle\nabla_{x}f(x_{k},y_{k}),x-x_{k}\right\rangle+\frac{\beta_{k}}{2}\|x-x_{k}\|^{2}$ (83)
+$=\mathcal{P}_{\mathcal{X}}\left(x_{k}-\frac{1}{\beta_{k}}\nabla_{x}f(x_{k},y_{k})\right),$
+where $\mathcal{P}_{\mathcal{X}}$ denotes the projection operator and $\beta_{k}>0$ is the step-size parameter.
+Meanwhile, for the variable $y$ , the proposed algorithm solves the following inner maximization subproblem with respect to $y$ using a regularized function instead of the original $f(x,y)$,
+$y_{k+1}=\arg\max_{y\in\mathcal{Y}}\tilde{f}_{k}(x_{k+1},y).$ (84)
+Similar to the idea of PF-AGP-NSC and PF-AGP-NC algorithms, we iteratively estimate the Lipschitz constants $L_{11}$ and $L_{12}$.
+In each iteration of the algorithm, our goal is to find a tuple $(l^{k,i}_{11},l^{k,i}_{12})$ that satisfies conditions (2.1) and (C2) by backtracking.
+The proposed algorithm, denoted as PF-AGP-NL, is formally presented in Algorithm 3.
+Algorithm 3 A parameter-free alternating gradient projection (PF-AGP-NL) algorithm for nonconvex-linear minimax problems
+  Step 1: Input $x_{1},y_{1}$, $\beta_{0}$, $\gamma_{0}$, $c_{0}$, $d_{0}$, $l_{11}^{0}$, $l_{12}^{0}$; Set $k=1$.
+  Step 2: Updata $x_{k}$ and $y_{k}$:
+   (a): Set $i=0$$l_{11}^{k,i}={l_{11}^{k-1}}$, $l_{12}^{k,i}={l_{12}^{k-1}}$, $\beta_{k,i}=\beta_{k-1}$, $c_{k,i}=c_{k-1}$, $d_{k,i}=d_{k-1}$.
+   (b): Update $x_{{k,i}}$ and $y_{{k,i}}$:
+$x_{{k,i}}$ $=\mathcal{P}_{\mathcal{X}}\left(x_{k}-\frac{1}{\beta_{k,i}}\nabla_{x}f(x_{k},y_{k})\right),$ (85)
+$y_{{k,i}}$ $=\arg\max_{y\in\mathcal{Y}}f(x_{k,i},y)-\frac{c_{k,i}}{2}\|y\|^{2}-\frac{d_{k,i}}{2}\|y-y_{k}\|^{2},.$ (86)
+   (c):  Compute $C_{1}^{k,i}$ and $C_{2}^{k,i}$, as in (2.1) and (C2);
+   (d): Update $l^{k,i}_{11}$ and $l^{k,i}_{12}$:
+$l^{k,i+1}_{11}=\frac{\mbox{sgn}(C_{1}^{k,i})+3}{2}l^{k,i}_{11},\quad l^{k,i+1}_{12}=\frac{\mbox{sgn}(C_{2}^{k,i})+3}{2}l^{k,i}_{12},$ (87)
+   (e):  If $C_{1}^{k,i}\leq 0$ and $C_{2}^{k,i}\leq 0$, then
+     $x_{k+1}=x_{k,i}$, $y_{k+1}=y_{k,i}$, $l_{11}^{k}=l_{11}^{k,i+1}$, $l_{12}^{k}=l_{12}^{k,i+1}$, $\beta_{k}=\beta_{k,i}$, $c_{k}=c_{k,i}$, $d_{k}=d_{k,i}$,
+     go to Step 3;
+     Otherwise, $i=i+1$,
+$\rho_{k,i}=2\max\{l_{11}^{k,i},l^{k,i}_{12}\},\quad\beta_{k,i}=2\rho_{k,i}k^{1/3}+\rho_{k,i},\quad c_{k,i}=\frac{\rho_{k,i}}{k^{1/3}},\quad d_{k,i}=\frac{\rho_{k,i}}{16k^{1/3}}$ (88)
+     go to Step 2(b).
+  Step 3: If some stationary condition is satisfied, stop; Otherwise, set $k=k+1$, go to Step 2.
+3.1 Complexity Analysis
+In this subsection, we analyze the iteration complexity of Algorithm PF-AGP-NL for solving (P) under the general nonconvex-linear setting.
+Lemma 14
+Under Assumption 1, let $\{(x_{k},y_{k})\}$ be the sequence generated by Algorithm 3. Then we have
+$\begin{split}&\tilde{f}_{k+1}(x_{k+1},y_{k+1})-\tilde{f}_{k}(x_{k},y_{k})\\ \leq&-\left(\beta_{k}-\frac{l_{11}^{k}}{2}-\frac{(l_{12}^{k})2}{c_{k}}\right)\|x_{k+1}-x_{k}\|^{2}\\ &-(-\frac{c_{k}}{4}+\frac{c_{k-1}}{2}+\frac{d_{k}-d_{k-1}}{2})\|y_{k+1}-y_{k}\|^{2}+\frac{c_{k-1}-c_{k+1}}{2}\|y_{k+1}\|^{2}\\ &-\frac{c_{k-1}-c_{k}}{2}\|y_{k}\|^{2}+\frac{d_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}.\end{split}$ (89)
+Proof  First, by the strong concavity of ${f}_{k}(x,y)$ with respect to $y$, we have
+$\tilde{f}_{k}(x_{k+1},y_{k+1})-\tilde{f}_{k}(x_{k+1},y_{k})\leq\langle\nabla_{y}\tilde{f}_{k}(x_{k+1},y_{k}),y_{k+1}-y_{k}\rangle-\frac{c_{k}+d_{k}}{2}\|y_{k+1}-y_{k}\|^{2}.$ (90)
+From the optimality condition of (86) in Algorithm 1, we obtain
+$\langle\nabla_{y}\tilde{f}_{k-1}(x_{k},y_{k}),y-y_{k}\rangle\leq 0.$ (91)
+Let $y=y_{k+1}$ and we then get
+$\langle\nabla_{y}\tilde{f}_{k-1}(x_{k},y_{k}),y_{k+1}-y_{k}\rangle\leq 0.$ (92)
+Substituting (92) into (90), we get
+$\begin{split}\tilde{f}_{k}(x_{k+1},y_{k+1})-\tilde{f}_{k}(x_{k+1},y_{k})\leq&\langle\nabla_{y}\tilde{f}_{k}(x_{k+1},y_{k})-\nabla_{y}\tilde{f}_{k-1}(x_{k},y_{k}),y_{k+1}-y_{k}\rangle\\ &-\frac{c_{k}+d_{k}}{2}\|y_{k+1}-y_{k}\|^{2}.\end{split}$ (93)
+By the Cauchy-Schwarz inequality and the backtracking strategy (C2), we obtain
+$\langle\nabla_{y}\tilde{f}(x_{k+1},y_{k})-\nabla_{y}\tilde{f}(x_{k},y_{k}),y_{k+1}-y_{k}\rangle\leq\frac{(l_{12}^{k})^{2}}{c_{k}}\|x_{k+1}-x_{k}\|^{2}+\frac{c_{k}}{4}\|y_{k+1}-y_{k}\|^{2}.$ (94)
+Moreover, we have
+$\begin{split}\langle y_{k},y_{k+1}-y_{k}\rangle=\frac{1}{2}\|y_{k+1}\|^{2}-\frac{1}{2}\|y_{k}\|^{2}-\frac{1}{2}\|y_{k+1}-y_{k}\|^{2}.\end{split}$ (95)
+Combining (94)-(95) with (93) and applying the Cauchy-Schwarz inequality, we obtain
+$\begin{split}&\tilde{f}_{k}(x_{k+1},y_{k+1})-\tilde{f}_{k}(x_{k+1},y_{k})\\ \leq&\frac{(l_{12}^{k})^{2}}{c_{k}}\|x_{k+1}-x_{k}\|^{2}-(-\frac{c_{k}}{4}+\frac{c_{k-1}}{2}+\frac{d_{k}-d_{k-1}}{2})\|y_{k+1}-y_{k}\|^{2}\\ &+\frac{d_{k-1}}{2}\|y_{k}-y_{k-1}\|+\frac{c_{k-1}-c_{k}}{2}(\|y_{k+1}\|^{2}-\|y_{k}\|^{2}).\end{split}$ (96)
+From Lemma 3 and (96), we obtain
+$\tilde{f}_{k}(x_{k+1},y_{k+1})-\tilde{f}_{k}(x_{k},y_{k})\leq$ $-\left(\beta_{k}-\frac{l_{11}^{k}}{2}-\frac{(l_{12}^{k})^{2}}{c_{k}}\right)\|x_{k+1}-x_{k}\|^{2}+\frac{d_{k-1}}{2}\|y_{k}-y_{k-1}\|$
+$-(-\frac{c_{k}}{4}+\frac{c_{k-1}}{2}+\frac{d_{k}-d_{k-1}}{2})\|y_{k+1}-y_{k}\|^{2}$
+$+\frac{c_{k-1}-c_{k}}{2}\left(\|y_{k+1}\|^{2}-\|y_{k}\|^{2}\right).$ (97)
+Finally, combining the definition of $f_{k}(x,y)$ with (97), we conclude that
+$\tilde{f}_{k+1}(x_{k+1},y_{k+1})-\tilde{f}_{k}(x_{k},y_{k})\leq$ $-\left(\beta_{k}-\frac{l_{11}^{k}}{2}-\frac{(l_{12}^{k})^{2}}{c_{k}}\right)\|x_{k+1}-x_{k}\|^{2}+\frac{d_{k-1}}{2}\|y_{k}-y_{k-1}\|$
+$+\frac{c_{k-1}-c_{k+1}}{2}\|y_{k+1}\|^{2}-\frac{c_{k-1}-c_{k}}{2}\|y_{k}\|^{2}$
+$-(-\frac{c_{k}}{4}+\frac{c_{k-1}}{2}-\frac{d_{k-1}}{2})\|y_{k+1}-y_{k}\|^{2},$ (98)
+which completes the proof.
+Theorem 15
+Suppose Assumption 1 holds. Set
+$H_{k+1}:=\tilde{f}_{k+1}(x_{k+1},y_{k+1})+\frac{d_{k}}{2}\|y_{k+1}-y_{k}\|^{2}-\frac{c_{k}-c_{k+1}}{2}\|y_{k+1}\|^{2}.$ (99)
+For any given $\varepsilon>0$, choose
+$c_{k}=\frac{\rho_{k}}{k^{1/3}},\quad d_{k}=\frac{c_{k}}{8},\quad\beta_{k}=2\rho_{k}k^{1/3}+\rho_{k}.$ (100)
+Let $\rho_{k}=2\max\{l_{11}^{k},l^{k}_{12}\}$, then
+$T(\varepsilon)\leq\max\left\{\left(\frac{H_{2}-\underline{H}+11\hat{\mathcal{D}}_{y}^{2}}{\tau\varepsilon^{2}}+1\right)^{3/2},\left(\frac{8L\hat{\mathcal{D}}_{y}}{\varepsilon}\right)^{3}\right\},$ (101)
+where $\tau=\frac{1}{2^{6}3^{2}L}$, $\underline{H}=\underline{F}-4{L}\hat{\mathcal{D}}_{y}^{2}$, $\underline{F}=\min_{(x,y)\in\mathcal{X}\times\mathcal{Y}}\tilde{F}(x,y)$ and $\hat{\mathcal{D}}_{y}=\max\{\|y\|\,|y\in Y\}$.
+Proof  From Lemma 14 and the definition of $H_{k+1}$, we get
+$\begin{split}&\left(\beta_{k}-\frac{l_{11}^{k}}{2}-\frac{(l_{12}^{k})^{2}}{c_{k}}\right)\|x_{k+1}-x_{k}\|^{2}+(\frac{c_{k}}{4}-d_{k})\|y_{k+1}-y_{k}\|^{2}\\ \leq&H_{k}-H_{k+1}+\frac{c_{k-1}-c_{k}}{2}\|y_{k+1}\|^{2}+\frac{c_{k}-c_{k-1}}{2}\|y_{k+1}-y_{k}\|^{2}+\frac{d_{k-1}-d_{k}}{2}\|y_{k+1}-y_{k}\|^{2}.\end{split}$ (102)
+By setting the parameters of $\beta_{k}$, we can easily obtain
+$\beta_{k}-\frac{l_{11}^{k}}{2}-\frac{(l_{12}^{k})^{2}}{c_{k}}\geq\frac{\rho_{k}^{2}}{c_{k}}\quad\text{and}\quad\frac{c_{k}}{4}-d_{k}=\frac{c_{k}}{8}$ (103)
+Combined with (102)-(103), we get
+$\begin{split}&\frac{\rho_{k}^{2}}{c_{k}}\|x_{k+1}-x_{k}\|^{2}+\frac{c_{k}}{8}\|y_{k+1}-y_{k}\|^{2}\\ \leq&H_{k}-H_{k+1}+\frac{c_{k-1}-c_{k}}{2}\|y_{k+1}\|^{2}+\frac{c_{k}-c_{k-1}}{2}\|y_{k+1}-y_{k}\|^{2}+\frac{d_{k-1}-d_{k}}{2}\|y_{k+1}-y_{k}\|^{2}.\end{split}$ (104)
+According to (3), we immediately get
+$\|\left(\nabla\tilde{G}_{k}\right)_{x}\|=\beta_{k}\|x_{k+1}-x_{k}\|.$ (105)
+On the other hand, according to (4) and the triangle inequality, we can prove that
+$\begin{split}\|\left(\nabla\tilde{G}_{k}\right)_{y}\|&\leq{\rho_{k}}\|y_{k}-A+A-B\|\\ &\leq l_{12}^{k}\|x_{k+1}-x_{k}\|+\left(2\rho_{k}+c_{k}\right)\|y_{k+1}-y_{k}\|,\end{split}$ (106)
+where $A=\mathcal{P}_{y}\left(y_{k+1}+\frac{1}{\rho_{k}}\nabla_{y}f_{k}(x_{k+1},y_{k+1})\right)$ and $B=\mathcal{P}_{y}\left(y_{k}+\frac{1}{\rho_{k}}\nabla_{y}f_{k}(x_{k},y_{k})\right)$.
+Combining (105), (106) and (C2), and applying the Cauchy-Schwarz inequality and $c_{k}\leq\rho_{k}$, we get
+$\begin{split}\|\nabla\tilde{G}_{k}\|^{2}&\leq\left(\beta_{k}^{2}+2(l_{12}^{k})^{2}\right)\|x_{k+1}-x_{k}\|^{2}+2\left(2\rho_{k}+c_{k}\right)^{2}\|y_{k+1}-y_{k}\|^{2}\\ &\leq\left(\beta_{k}^{2}+2(l_{12}^{k})^{2}\right)\|x_{k+1}-x_{k}\|^{2}+18\rho_{k}^{2}\|y_{k+1}-y_{k}\|^{2}.\end{split}$ (107)
+Multiplying both sides of (107) by $m_{k}c_{k}$, combined with (104) and using the definition of $\hat{\mathcal{D}}_{y}$, we get
+$\begin{split}m_{k}c_{k}\|\nabla\tilde{G}_{k+1}\|^{2}&\leq\frac{\rho_{k}^{2}}{c_{k}}\|x_{k+1}-x_{k}\|^{2}+\frac{c_{k}}{8}\|y_{k+1}-y_{k}\|^{2}\\ &\leq H_{k}-H_{k+1}+\frac{c_{k-1}-c_{k}}{2}\|y_{k+1}\|^{2}+\frac{c_{k}-c_{k-1}}{2}\|y_{k+1}-y_{k}\|^{2}\\ &\quad+\frac{d_{k-1}-d_{k}}{2}\|y_{k+1}-y_{k}\|^{2}\\ &\leq H_{k}-H_{k+1}+4L({\bar{c}_{k-1}-\bar{c}_{k}})\hat{\mathcal{D}}_{y}^{2}+\frac{7}{4}(\rho_{k}-\rho_{k-1})\hat{\mathcal{D}}_{y}^{2},\end{split}$ (108)
+where $m_{k}=\min\left\{\frac{\rho^{2}_{k}}{c^{2}_{k}(2(l_{12}^{k})^{2}+\beta^{2}_{k})},\frac{1}{144\rho^{2}_{k}}\right\}=\frac{1}{144\rho^{2}_{k}}$ and $\bar{c}_{k}=\frac{1}{k^{1/3}}$.
+Define $\tilde{T}(\varepsilon):=\min\{k\mid\|\nabla\tilde{G}(x_{k},y_{k})\|\leq\frac{\varepsilon}{2},k\geq 2\}$.
+Summing (108) from $k=2$ to $\tilde{T}(\varepsilon)$, and using $c_{k}\leq\rho_{k}$, we obtain
+$\sum_{k=2}^{\tilde{T}(\varepsilon)}m_{k}c_{k}\|\nabla\tilde{G}_{k+1}\|^{2}\leq\sum_{k=2}^{\tilde{T}(\varepsilon)}\left(H_{k}-H_{k+1}+4L({\bar{c}_{k-1}-\bar{c}_{k}})\hat{\mathcal{D}}_{y}^{2}+11(\rho_{k}-\rho_{k-1})\hat{\mathcal{D}}_{y}^{2},\right).$ (109)
+Further simplification yields
+$\sum_{k=2}^{\tilde{T}(\varepsilon)}m_{k}c_{k}\|\nabla\tilde{G}_{k+1}\|^{2}\leq H_{2}-H_{\tilde{T}(\varepsilon)+1}+11L\hat{\mathcal{D}}_{y}^{2}.$ (110)
+From the definition of $H_{k+1}$, we have
+$\begin{split}H_{\tilde{T}(\varepsilon)+1}&=f_{\tilde{T}(\varepsilon)+1}(x_{\tilde{T}(\varepsilon)+1},y_{\tilde{T}(\varepsilon)+1})+\frac{d_{\tilde{T}(\varepsilon)+1}}{2}\|y_{\tilde{T}(\varepsilon)+1}-y_{\tilde{T}(\varepsilon)}\|^{2}-\frac{c_{\tilde{T}(\varepsilon)}-c_{\tilde{T}(\varepsilon)+1}}{2}\|y_{\tilde{T}(\varepsilon)+1}\|^{2}\\ &\geq\underline{F}-{\rho_{\tilde{T}(\varepsilon)}}\hat{\mathcal{D}}_{y}^{2}\geq\underline{F}-4{L}\hat{\mathcal{D}}_{y}^{2}=\underline{H}.\end{split}$ (111)
+Combining (111) with (110), we then obtain
+$\sum_{k=2}^{\tilde{T}(\varepsilon)}\frac{{\tau}}{{k}^{1/3}}\|\nabla\tilde{G}_{k+1}\|^{2}\leq\frac{1}{144\rho_{k}{k}^{1/3}}\|\nabla\tilde{G}_{k+1}\|^{2}=m_{k}c_{k}\|\nabla\tilde{G}_{k+1}\|^{2}\leq H_{2}-\underline{H}+11L\hat{\mathcal{D}}_{y}^{2}.$ (112)
+According to the setting of $c_{k}$ and (112), when $k\geq 2$, we can get
+$\begin{split}\frac{\varepsilon^{2}}{4}&\leq\frac{H_{2}-\underline{H}+11L\hat{\mathcal{D}}_{y}^{2}}{\sum^{\tilde{T(\varepsilon)}}_{k=2}\frac{\tau}{k^{1/3}}}\\ &\leq\frac{H_{2}-\underline{H}+11L\hat{\mathcal{D}}_{y}^{2}}{\tau(\tilde{T}(\varepsilon)^{2/3}-1)},\end{split}$ (113)
+and then
+$\tilde{T}(\varepsilon)\leq\left(\frac{H_{2}-\underline{H}+11L\hat{\mathcal{D}}_{y}^{2}}{\tau\varepsilon^{2}}+1\right)^{3/2}.$ (114)
+On the other hand, when $k\geq\left(\frac{8L\hat{\mathcal{D}}_{y}}{\varepsilon}\right)^{3}\geq\left(\frac{2\rho_{k}\hat{\mathcal{D}}_{y}}{\varepsilon}\right)^{3}$, the parameter setting of $c_{k}$ ensures
+$c_{k}\hat{\mathcal{D}}_{y}\leq\frac{\varepsilon}{2}.$ (115)
+Therefore, there exists $T(\varepsilon)\leq\max\left(\tilde{T}(\varepsilon),\left(\frac{8L\hat{\mathcal{D}}_{y}}{\varepsilon}\right)^{3}\right)$ such that
+$\|\nabla G_{k+1}\|\leq\|\nabla\tilde{G}_{k+1}\|+c_{k}\hat{\mathcal{D}}_{y}\leq\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon,$ (116)
+which completes the proof.
+Remark 16
+In the PF-AGP-NL algorithm, the upper bound of the total number of backtracking steps is  $\sum\limits_{k=2}^{T(\varepsilon)}\left(\log_{2}\left(\frac{l^{k}_{11}}{l_{11}^{k-1}}\right)+\log_{2}\left(\frac{l^{k}_{12}}{l_{12}^{k-1}}\right)\right)=\log_{2}\left(\frac{2L_{11}}{l_{11}^{1}}\right)+\log_{2}\left(\frac{2L_{12}}{l_{12}^{1}}\right)=\mathcal{O}(\log(L))$, and the upper bound of the total number of line searches is $\mathcal{O}(L)$.
+According to Theorem 15, in the nonconvex-linear setting, the number of gradient evaluations required by the PF-AGP-NL algorithm to obtain an $\varepsilon$stationary point is $\mathcal{O}(L^{3}\varepsilon^{-3})$.
 4 Numerical results
@@
 4.1 Synthetic Minimax Problem
-Synthetic minimax problem can be formulated as the following nonconvex-strongly-concave minimax problem (Luo et al., 2022):
+Synthetic minimax problem can be formulated as the following nonconvex-strongly concave minimax problem (Luo et al., 2022):
 $\min_{\mathbf{x}\in\mathbb{R}^{3}}\max_{\mathbf{y}\in\mathbb{R}^{2}}f(\mathbf{x},\mathbf{y})=w(x_{3})-\frac{y_{1}^{2}}{40}+x_{1}y_{1}-\frac{5y_{2}^{2}}{2}+x_{2}y_{2},$ (117)
@@
 $w(x)=\begin{cases}\sqrt{\epsilon}\left(x+(\lambda+1)\sqrt{\epsilon}\right)^{2}-\frac{1}{3}\left(x+(\lambda+1)\sqrt{\epsilon}\right)^{3}-\frac{1}{3}(3\lambda+1)\epsilon^{3/2},&x\leq-\lambda\sqrt{\epsilon};\\ \epsilon x+\frac{\epsilon^{3/2}}{3},&-\lambda\sqrt{\epsilon}<x\leq-\sqrt{\epsilon};\\ -\sqrt{\epsilon}x^{2}-\frac{x^{3}}{3},&-\sqrt{\epsilon}<x\leq 0;\\ -\sqrt{\epsilon}x^{2}+\frac{x^{3}}{3},&0<x\leq\sqrt{\epsilon};\\ -\epsilon x+\frac{\epsilon^{3/2}}{3},&\sqrt{\epsilon}<x\leq\lambda\sqrt{\epsilon};\\ \sqrt{\epsilon}\left(x-(\lambda+1)\sqrt{\epsilon}\right)^{2}+\frac{1}{3}\left(x-(\lambda+1)\sqrt{\epsilon}\right)^{3}-\frac{1}{3}(3\lambda+1)\epsilon^{3/2},&\lambda\sqrt{\epsilon}\leq x.\end{cases}$
-It is easy to verify that this problem has a strict saddle point at $(\mathbf{x}^{*},\mathbf{y}^{*})=([0,0,0]^{\top},[0,0]^{\top})$.
+It is easy to verify that (117) has a strict saddle point at $(\mathbf{x}^{*},\mathbf{y}^{*})=([0,0,0]^{\top},[0,0]^{\top})$.
 Experimental setup.
-We compare the proposed PF-AGP-NSC algorithm with the SGDA-B algorithm, the TiAda algorithm and the AGP algorithm for solving (82) with $\epsilon=0.01$, $\lambda=5$.
+We compare the proposed PF-AGP-NSC algorithm with the SGDA-B algorithm, the NeAda algorithm, the TiAda algorithm and the AGP algorithm for solving (117) with $\epsilon=0.01$ and $\lambda=5$.
 The initial point is chosen as $(x_{0},y_{0})=([0,0,2]^{\top},[0,0]^{\top})$ for all four tested algorithms.
-For the PF-AGP-NSC algorithm, we set $l_{11}^{1}=0.01$ , $l_{12}^{1}=0.01$, $l_{22}^{1}=0.01$ and $\mu_{1}=1$.
+For the PF-AGP-NSC algorithm, we set $l_{11}^{1}=0.01$ , $l_{12}^{1}=0.01$, $l_{22}^{1}=0.01$ and $\mu_{1}=0.01$.
 For the SGDA-B algorithm, we set $\mu=1$, $\gamma=0.9$, $N=1$, $p=0.5$ and $\tilde{L}=\frac{10}{9}$.
+For the NeAda algorithm, we set $\eta_{x}=0.5$ and $\eta_{y}=0.1$.
 For the TiAda algorithm, we set $\alpha=0.6$, $\beta=0.4$, $\eta_{x}=0.8$, $\eta_{y}=0.4$, $v^{x}_{0}=v^{y}_{0}=1$.
@@
 The algorithms terminate when $\|\nabla f(x_{k},y_{k})\|\leq 10^{-5}$ satisfied.
-Figure 1: Numerical Performance of four tested algorithms for solving the Synthetic Minimax problem
+Figure 1: Numerical Performance of five tested algorithms for solving the Synthetic Minimax problem
 Results.
@@
 We find that the SGDA-B algorithm fails to converge, both the AGP algorithm and the PF-AGP-NC algorithms successfully approach the unique stationary point.
-Moreover, the PF-AGP-NC algorithm exhibits significantly better convergence performance than that of the AGP algorithm.
+Moreover, the PF-AGP-NC algorithm exhibits significantly better convergence performance than that of AGP algorithm.
 Figure 2: Numerical performance of three tested algorithms for solving the Dirac-GAN problem
@@
 where $F(x):=\left[f_{1}(x);\ldots;f_{M}(x)\right]\in\mathbb{R}^{M\times 1}$ with $f_{m}(x)=\frac{1}{\left|\mathcal{S}_{m}\right|}\sum_{i=1}^{\left|\mathcal{S}_{m}\right|}\ell\left(d_{i}^{m},l_{i}^{m},x\right)$ and $\ell$ can be any non-negative loss function, $M$ is the number of tasks and $x$ denotes the network parameters, $y\in\Delta$ describes the weights over different tasks and $\Delta$ is the simplex, i.e., $\Delta=\left\{(y_{1},\cdots,y_{M})^{T}\mid\sum_{m=1}^{M}y_{m}=1,y_{m}\geq 0\right\}$.
-We consider two image classification problems with MNIST (LeCun et al., 1998) and CIFAR10 datasets (Krizhevsky et al., 2009).
+We consider two image classification problems with MNIST (LeCun et al., 1998) and CIFAR10 datasets (Krizhevsky and Hinton, 2009).
+The dimension of $x$ is 57,044,810.
+The sample sizes of the CIFAR-10 and MNIST datasets are both 60,000.
 Our goal is to train a neural network that works on these two completely unrelated problems simultaneously.
@@
 Experiment setup.
-We compare the proposed PF-AGP-NC algorithm with the SGDA-B algorithm and the AGP algorithm.
-For the PF-AGP-NC algorithm, we set $l_{11}^{1}=0.01$, $l_{12}^{1}=0.01$ and $l_{22}^{1}=0.01$.
+We compare the proposed PF-AGP-NC algorithm, PF-AGP-NL algorithm with the SGDA-B algorithm and the AGP algorithm.
+For the PF-AGP-NC algorithm, we set $l_{11}^{1}=0.01$ , $l_{12}^{1}=0.01$ and $l_{22}^{1}=5$.
+For the PF-AGP-NL algorithm, we set $l_{11}^{1}=0.1$ and $l_{12}^{1}=0.1$.
 For the SGDA-B algorithm, we set $\mu=20$, $\gamma=0.9$, $N=1$, $p=0.5$ and $\tilde{L}$ = 20.
@@
 Due to hardware (especially memory) limitations, we use a mini-batch randomized gradient instead of the exact gradient at each iteration in our numerical experiment, which is the same as in (Xu et al., 2023).
-Figure 3: Performance of two algorithms for solving robust multi-task learning problem.
+Figure 3: Numerical Performance of four tested algorithms for solving robust multi-task learning problem.
 Results.
 Figure 3 and Figure 3 show the testing accuracy on the MNIST dataset and the CIFAR10 dataset respectively.
-PF-AGP-NC algorithm outperforms the SGDA-B algorithm and the AGP algorithm for both datasets.
+The PF-AGP-NL algorithm and PF-AGP-NC algorithm outperform the SGDA-B algorithm and the AGP algorithm for both datasets.
 5 Conclusions
-In this paper, based on the framework of backtracking, we propose two completely parameter-free alternating gradient projection algorithms, i.e., the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm, for solving nonconvex-(strongly) concave minimax problems respectively, which does not require prior knowledge of parameters such as the Lipschtiz constant $L$ or the strongly concave constant $\mu$.
-Moreover, we show that the total number of gradient calls of the PF-AGP-NSC algorithm to obtain an $\varepsilon$-stationary point for nonconvex-strongly concave minimax problems is upper bounded by $\mathcal{O}\left(L\kappa^{3}\varepsilon^{-2}\right)$ where $\kappa$ is the condition number, while the total number of gradient calls of the PF-AGP-NC algorithm to obtain an $\varepsilon$-stationary point for nonconvex-concave minimax problems is upper bounded by $\mathcal{O}\left(L^{4}\varepsilon^{-4}\right)$.
-As far as we know, the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm are the first completely parameter-free algorithms for solving nonconvex-strongly concave minimax problems and nonconvex-concave minimax problems respectively.
-Numerical results demonstrate the efficiency of the two proposed algorithms.
+In this paper, based on the framework of backtracking, we propose two completely parameter-free alternating gradient projection algorithms, i.e., the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm, for solving nonconvex-(strongly) concave minimax problems respectively, which does not require prior knowledge of parameters such as the Lipschtiz constant $L$ or the strongly concavity modulus $\mu$.
+Moreover, we show that the total number of gradient calls of the PF-AGP-NSC algorithm to obtain an $\varepsilon$-stationary point for nonconvex-strongly concave minimax problems is upper bounded by  $\mathcal{O}\left(L^{2}\kappa^{3}\varepsilon^{-2}\right)$ where $\kappa$ is the condition number, while the total number of gradient calls of the PF-AGP-NC algorithm to obtain an $\varepsilon$-stationary point for nonconvex-concave minimax problems is upper bounded by  $\mathcal{O}\left(\log^{2}(L)L^{4}\varepsilon^{-4}\right)$.
+Moreover, we also propose a completely parameter-free alternating gradient projection algorithm, i.e., the PF-AGP-NL algorithm, for solving nonconvex-linear minimax problems, which does not require any prior knowledge of parameters.
+As far as we know, the PF-AGP-NSC algorithm, the PF-AGP-NC algorithm and the PF-AGP-NL algorithm are the first completely parameter-free algorithms for solving nonconvex-strongly concave minimax problems, nonconvex-concave minimax problems and nonconvex-linear minimax problems respectively.
+Numerical results demonstrate the efficiency of the proposed algorithms.
+Appendix A A restarted parameter-free single-loop algorithm for nonconvex-strongly concave minimax problems
+In this section, we propose a restarting parameter-free alternating gradient projection algorithm (rPF-AGP-NSC) for the nonconvex-strongly concave minimax problem.
+The algorithm is a restarting variant of PF-AGP-NSC that does not require the prior knowledge of the Lipschtiz constant $L$, but requires the prior knowledge of two additional parameters $\mu$ and $\underline{S}$.
+The algorithm is formally presented in Algorithm 4.
+Next, we prove the iteration complexity of Algorithm 4 under the nonconvex-strongly concave setting.
+Denote $\Phi(x):=\max_{y\in\mathcal{Y}}f(x,y)$ and $y^{*}(x):=\arg\max_{y\in\mathcal{Y}}f(x,y)$.
+By Lemma B.1 in (Nouiehed et al., 2019a), $\Phi(x)$ is $L_{\Phi}$-Lipschitz smooth with $L_{\Phi}=L+\frac{L^{2}}{\mu}$.
+Moreover, we have
+$\nabla_{x}\Phi(x)=\nabla_{x}f(x,y^{*}(x)).$ (120)
+Algorithm 4 A restarted parameter-free alternating gradient projection (rPF-AGP-NSC) algorithm for nonconvex-strongly concave minimax problems
+  Step 1: Input $x_{1,1},y_{1,1}$, $l_{1}$, $\mu$, $\varepsilon$, $\underline{S}$; Set $i=1$. Compute: $S_{1}=2\max_{y\in\mathcal{Y}}f(x_{1,1},y)-f(x_{1,1},y_{1,1})$.
+  Step 2: Comppute $\beta_{i}=3l_{i}$, $\alpha_{i}=\frac{158l_{i}^{3}}{\mu^{2}}$, $\eta_{i}=\frac{(2\beta_{i}+\mu)(\beta_{i}+l_{i})}{\mu\beta_{i}}$, $\tilde{d}^{i}_{1}=\min\{\frac{\alpha_{i}-\frac{l_{i}(l_{i}+\beta_{i})^{2}\eta_{i}^{2}}{\beta_{i}^{2}}-\frac{l_{i}^{2}}{\mu}-\frac{3l_{i}}{2}}{2\alpha_{i}^{2}},\frac{\beta_{i}-\frac{3l_{i}}{2}}{\beta_{i}^{2}+2l_{i}^{2}}\}$, $\tilde{T}_{i}=\frac{S_{1}-\underline{S}}{\varepsilon^{2}\tilde{d}^{i}_{1}}$; Set $k=1$;
+  Step 3: Update $x_{i,k+1}$ and $y_{i,k+1}$:
+   (a): If $k\geq\tilde{T}_{i}$ then, go to Step 4;
+   (b): Update $x_{i,{k+1}}$ and $y_{i,{k+1}}$
+$y_{i,{k+1}}$ $=\mathcal{P}_{\mathcal{Y}}\left(y_{i,k}+\frac{1}{\beta_{i}}\nabla_{y}f(x_{i,{k}},y_{i,k})\right),$ (121)
+$x_{i,{k+1}}$ $=\mathcal{P}_{\mathcal{X}}\left(x_{i,k}-\frac{1}{\alpha_{i}}\nabla_{x}f(x_{i,k},y_{i,k+1})\right).$ (122)
+   (c): Set $k=k+1$ go to Step 3.
+  Step 4: If $\|\nabla G(x_{i,k},y_{i,k})\|\leq\varepsilon$, stop; Otherwise $x_{i+1,1}=x_{1,1}$, $y_{i+1,1}=y_{1,1}$, $l_{i+1}=2l_{i}$, $i=i+1$, go to Step 2.
+Lemma 17
+Suppose that Assumption 1 holds.
+Let $\{\left(x_{i,k},y_{i,k}\right)\}$ be a sequence generated by Algorithm 4.
+Let $\eta_{i}=\frac{(2\beta_{i}+\mu)(\beta_{i}+L)}{\mu\beta_{i}}$, then $\forall k,i\geq 1$, and for any $a>0$,
+$\Phi(x_{i,k+1})-\Phi(x_{i,k})$
+$\leq$ $\langle\nabla_{x}f(x_{i,k},y_{i,k+1}),x_{i,k+1}-x_{i,k}\rangle+\frac{L^{2}(L+\beta_{i})^{2}\eta_{i}^{2}}{2a\beta_{i}^{2}}\|y_{i,k+1}-y_{i,k}\|^{2}+\frac{a+L_{\Phi}}{2}\|x_{i,k+1}-x_{i,k}\|^{2}.$ (123)
+Proof  Since $\Phi(x)$ is $L_{\Phi}$ smooth with respect to $x$, and by (120), we have
+$\Phi(x_{i,k+1})-\Phi(x_{i,k})\leq$ $\langle\nabla_{x}f(x_{i,k},y^{*}(x_{i,k}))-\nabla_{x}f(x_{i,k},y_{i,k+1}),x_{i,k+1}-x_{i,k}\rangle$
+$+\frac{L_{\Phi}}{2}\|x_{i,k+1}-x_{i,k}\|^{2}+\langle\nabla_{x}f(x_{i,k},y_{i,k+1}),x_{i,k+1}-x_{i,k}\rangle.$ (124)
+By the Cauchy-Schwarz inequality, for any $a>0$, we obtain
+$\langle\nabla_{x}f(x_{i,k},y^{*}(x_{i,k}))-\nabla_{x}f(x_{i,k},y_{i,k+1}),x_{i,k+1}-x_{i,k}\rangle$
+$\leq$ $\frac{L^{2}}{2a}\|y_{i,k+1}-y^{*}(x_{i,k})\|^{2}+\frac{a}{2}\|x_{i,k+1}-x_{i,k}\|^{2}.$ (125)
+By Lemma 1 in (Zhang et al., 2022), we have
+$\|y_{i,k+1}-y^{*}(x_{i,k})\|\leq\eta_{i}\left\|y_{i,k+1}-\mathcal{P}_{\mathcal{Y}}\left(y_{i,k+1}+\frac{1}{\beta_{i}}\nabla_{y}f\left(x_{i,k},y_{i,k+1}\right)\right)\right\|.$ (126)
+By (121), the non-expansive property of the projection operator $\mathcal{P}_{\mathcal{Y}}(\cdot)$ and Assumption 1, we obtain
+$\left\|y_{i,k+1}-\mathcal{P}_{\mathcal{Y}}\left(y_{i,k+1}+\frac{1}{\beta_{i}}\nabla_{y}f\left(x_{i,k},y_{i,k+1}\right)\right)\right\|\leq\frac{L+\beta_{i}}{\beta_{i}}\|y_{i,k+1}-y_{i,k}\|.$ (127)
+Combing (A), (125), (126) and (127), we get
+$\Phi(x_{i,k+1})-\Phi(x_{i,k})\leq$ $\frac{L^{2}\eta_{i}^{2}(L+\beta_{i})^{2}}{2a\beta_{i}^{2}}\|y_{i,k+1}-y_{i,k}\|^{2}+\frac{a+L_{\Phi}}{2}\|x_{i,k+1}-x_{i,k}\|^{2}$
+$+\langle\nabla_{x}f(x_{i,k},y_{i,k+1}),x_{i,k+1}-x_{i,k}\rangle,$ (128)
+which completes the proof.
+Next, we provide an lower bound for the difference between $f(x_{i,k+1},y_{i,k+1})$ and $f(x_{i,k},y_{i,k})$.
+Lemma 1
+Suppose that Assumption 1 holds. Let $\{\left(x_{i,k},y_{i,k}\right)\}$ be a sequence generated by Algorithm 4. Then $\forall i,k\geq 1$,
+$f(x_{i,k+1},y_{i,k+1})-f(x_{i,k},y_{i,k})$
+$\geq$ $\left(\beta_{i}-\frac{L}{2}\right)\|y_{i,k+1}-y_{i,k}\|^{2}-\frac{L}{2}\|x_{i,k+1}-x_{i,k}\|^{2}+\langle\nabla_{x}f(x_{i,k},y_{i,k+1}),x_{i,k+1}-x_{i,k}\rangle,$ (129)
+Proof  The optimality condition for $y_{i,k}$ in (121) implies that $\forall y\in\mathcal{Y}$ and $\forall i,k\geq 1$,
+$\langle\nabla_{y}f(x_{i,k},y_{i,k}),y_{i,k+1}-y_{i,k}\rangle\geq\beta_{i}\|y_{i,k+1}-y_{i,k}\|^{2}.$ (130)
+By Assumption 1, the gradient of $f(x,y)$ is Lipschitz continuous with respect to $y$ and (130), we get
+$f(x_{i,k},y_{i,k+1})-f(x_{i,k},y_{i,k})\geq$ $\left(\beta_{i}-\frac{L}{2}\right)\|y_{i,k+1}-y_{i,k}\|^{2}.$ (131)
+Since that the gradient of $f(x,y)$ is Lipschitz continuous with respect to $x$, we obtain
+$f(x_{i,k+1},y_{i,k+1})-f(x_{i,k},y_{i,k+1})$
+$\geq$ $\langle\nabla_{x}f(x_{i,k},y_{i,k+1}),x_{i,k+1}-x_{i,k}\rangle-\frac{L}{2}\|x_{i,k+1}-x_{i,k}\|^{2}.$ (132)
+The proof is then completed by adding (131) and (A).
+We now establish an important recursion for the Algorithm 4.
+Lemma 2
+Suppose that Assumption 1 holds. Denote $S(x,y)=2\Phi(x)-f(x,y).$ Let $\{\left(x_{i,k},y_{i,k}\right)\}$ be a sequence generated by Algorithm 4. Then $\forall i,k\geq 1$,
+$S(x_{i,k+1},y_{i,k+1})-S(x_{i,k},y_{i,k})\leq$ $-\left(\alpha_{i}-\frac{L(L+\beta_{i})^{2}\eta_{i}^{2}}{\beta_{i}^{2}}-\frac{L^{2}}{\mu}-\frac{3L}{2}\right)\|x_{i,k+1}-x_{i,k}\|^{2}$
+$-\left(\beta_{i}-\frac{3L}{2}\right)\|y_{i,k+1}-y_{i,k}\|^{2}.$ (133)
+Proof  Denote $S(x,y)=2\Phi(x)-f(x,y)$. Combining (17) and (1), we have
+$S(x_{i,k+1},y_{i,k+1})-S(x_{i,k},y_{i,k})$
+$\leq$ $-\left(\beta_{i}-\frac{L}{2}-\frac{(L+\beta_{i})^{2}\eta_{i}^{2}L^{2}}{a\beta_{i}^{2}}\right)\|y_{i,k+1}-y_{i,k}\|^{2}+\left(a+L_{\Phi}+\frac{L}{2}\right)\|x_{i,k+1}-x_{i,k}\|^{2}$
+$+\langle\nabla_{x}f(x_{i,k},y_{i,k+1}),x_{i,k+1}-x_{i,k}\rangle.$ (134)
+The optimality condition for $x_{i,k}$ in (122) implies that $\forall x\in\mathcal{X}$ and $\forall k\geq 1$,
+$\langle\nabla_{x}f(x_{i,k},y_{i,k+1}),x_{i,k+1}-x_{i,k}\rangle\leq-\alpha_{i}\|x_{i,k+1}-x_{i,k}\|^{2}.$ (135)
+Plugging (135) into (A), then we obtain
+$S(x_{i,k+1},y_{i,k+1})-S(x_{i,k},y_{i,k})$
+$\leq$ $-\left(\beta_{i}-\frac{L}{2}-\frac{L^{2}(L+\beta_{i})^{2}\eta^{2}}{a\beta_{i}^{2}}\right)\|y_{i,k+1}-y_{i,k}\|^{2}-\left(\alpha_{i}-a-L_{\Phi}-\frac{L}{2}\right)\|x_{i,k+1}-x_{i,k}\|^{2}.$ (136)
+The proof is then completed by the definiton of $L_{\Phi}=L+\frac{L^{2}}{\mu}$ and choosing $a=\frac{L(L+\beta_{i})^{2}\eta_{i}^{2}}{\beta_{i}^{2}}$ in (A).
+We are now ready to establish the iteration complexity for the rPF-AGP-NSC algorithm.
+In particular, letting $\nabla G(x_{i,k},y_{i,k})$ be defined as in Definition 1 and $\varepsilon>0$ be a given target accuracy, we provide a bound on $T_{i}(\varepsilon)$, the inner iteration index to achieve an $\varepsilon$-stationary point, i.e., $\|\nabla G(x_{i,k},y_{i,k})\|\leq\varepsilon$, which is equivalent to
+$T_{i}(\varepsilon)=\min\{k\mid\|\nabla G(x_{i,k},y_{i,k})\|\leq\varepsilon\}.$ (137)
+The following theorem gives an upper bound on $T_{i}(\varepsilon)$.
+Theorem 1
+Suppose that Assumptions 1 hold. Let $\{\left(x_{i,k},y_{i,k}\right)\}$ be a sequence generated by Algorithm 4. Let $\eta_{i}=\frac{(2\beta_{i}+\mu)(\beta_{i}+L)}{\mu\beta_{i}}$. If
+$\beta_{i}>\frac{3L}{2},\alpha_{i}>\frac{L(L+\beta_{i})^{2}\eta_{i}^{2}}{\beta_{i}^{2}}+\frac{L^{2}}{\mu}+\frac{3L}{2},$ (138)
+then $\forall\varepsilon>0$, we have
+$T_{i}\left(\varepsilon\right)\leq\frac{d_{2}^{i}}{\varepsilon^{2}d^{i}_{1}},$
+where $d^{i}_{1}:=\min\{{\frac{\alpha_{i}-\frac{L(L+\beta_{i})^{2}\eta_{i}^{2}}{\beta_{i}^{2}}-\frac{L^{2}}{\mu}-\frac{3L}{2}}{2\alpha_{i}^{2}},\frac{\beta_{i}-\frac{3L}{2}}{\beta_{i}^{2}+2L^{2}}\}}$, $d_{2}^{i}:=S(x_{i,1},y_{i,1})-\underline{S}$ with $\underline{S}:=\min\limits_{x\in\mathcal{X}}\max\limits_{y\in\mathcal{Y}}f(x,y)$.
+Proof  By (121), we immediately obtain
+$\left\|\beta_{i}\left(y_{i,k}-\mathcal{P}_{\mathcal{Y}}\left(y_{i,k}+\frac{1}{\beta_{i}}\nabla_{y}f\left(x_{i,k},y_{i,k},\right)\right)\right)\right\|=\beta_{i}\|y_{i,k+1}-y_{i,k}\|.$ (139)
+On the other hand, by (122) and the Cauchy-Schwartz inequality, we have
+$\|\alpha_{i}(x_{i,k}-\mathcal{P}_{\mathcal{Y}}(x_{i,k}-\frac{1}{\alpha_{i}}\nabla_{x}f\left(x_{i,k},y_{i,k}\right)))\|\leq\alpha_{i}\|x_{i,k+1}-x_{i,k}\|+L\|y_{i,k+1}-y_{i,k}\|,$ (140)
+where the last inequality is by the nonexpansive property of the projection operator and Assumption 1. Combing (139), (140), and using Cauchy-Schwarz inequality, we have
+$\|\nabla G(x_{i,k},y_{i,k})\|^{2}\leq(\beta_{i}^{2}+2L^{2})\|y_{i,k+1}-y_{i,k}\|^{2}+2\alpha_{i}^{2}\|x_{i,k+1}-x_{i,k}\|^{2}.$ (141)
+By (138), it can be easily checked that $d^{i}_{1}>0$. By multiplying $d^{i}_{1}$ on the both sides of (141) and using (2) in Lemma 2, we get
+$d^{i}_{1}\|\nabla G(x_{i,k},y_{i,k})\|^{2}\leq S(x_{i,k},y_{i,k})-S(x_{i,k+1},y_{i,k+1}).$ (142)
+Summing both sides of (142) from $k=1$ to $T_{i}(\varepsilon)$, we then obtain
+$\sum_{k=1}^{T_{i}\left(\varepsilon\right)}{d^{i}_{1}\|\nabla G(x_{i,k},y_{i,k})\|^{2}}\leq S(x_{i,1},y_{i,1})-S(x_{i,T_{i}(\varepsilon)+1},y_{i,T_{i}(\varepsilon)+1}).$ (143)
+Then, by (143) we obtain $\sum_{k=1}^{T_{i}\left(\varepsilon\right)}{d^{i}_{1}\|\nabla G(x_{i,k},y_{i,k})\|^{2}}\leq S(x_{i,1},y_{i,1})-\underline{S}=d_{2}^{i}.$ In view of the definition of $T_{i}(\varepsilon)$, the above inequality implies that $\varepsilon^{2}\leq\frac{d_{2}^{i}}{T_{i}(\varepsilon)d^{i}_{1}}$ or equivalently, $T_{i}\left(\varepsilon\right)\leq\frac{d_{2}^{i}}{\varepsilon^{2}d^{i}_{1}}$.
+This completes the proof.
+Remark 18
+Let $\beta_{i}=3l_{i},\alpha_{i}=\frac{158l_{i}^{3}}{\mu^{2}}$, then the outer iteration of the rPF-AGP-NSC algorithm requires at most $\mathcal{O}(\log(L))$ steps, so that $\beta_{i}$ and $\alpha_{i}$ satisfy (138).
+According to Theorem 1, the inner iteration of $\|\nabla G(x_{i,k},y_{i,k})\|\leq\varepsilon$ requires at most $\frac{d_{2}^{i}}{\varepsilon^{2}d^{i}_{1}}$ steps.
+Remark 19
+Denote $\kappa=L/\mu$.
+If we set $\beta_{i}=3l_{i}$, by Theorem 1 we have that $\eta_{i}=\mathcal{O}(\kappa)$, $d_{1}^{i}=\mathcal{O}(\frac{1}{L\kappa^{2}})$ and $d_{2}^{i}=O(1)$, which implies the total number of gradient calls is upper bounded by $\sum\limits_{i=1}^{\log_{2}(L)+1}T_{i}(\varepsilon)=\mathcal{O}(\log(L)L\kappa^{2}\varepsilon^{-2})$ under the nonconvex-strongly concave setting.
+It means that if two additional parameters $\mu$ and $\underline{S}$ are known in advance, the gradient complexity of the rPF-AGP-NSC algorithm is improved to $\mathcal{O}\left(\log(L)L\kappa^{2}\varepsilon^{-2}\right)$, which is better than that of the SGDA-B and the PF-AGP-NSC algorithm.